A Complete Bibliography of Publications in Isis, 1950–1959

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 0.07

Title word cross-reference

-1662 [300].

1 [1123]. 11.-18 [925]. 1238 [389]. 1267 [791]. 12th [804]. 1362 [349]. 1439 [294]. 15 [145]. 1500 [927]. 1524 [1045]. 1553 [595]. 1576 [471]. 1577 [297]. 1584 [864]. 1605 [820]. 1610 [614]. 1650 [513]. 1655 [1116]. 1666/7 [986]. 1679 [502]. 1683 [156]. 1700 [643]. 1714 [1105]. 1745 [506]. 1750 [1042]. 1793 [1054]. 1794 [731]. 1799 [1052]. 1818 [729]. 1823 [356]. 1829 [823]. 1833 [936]. 1850 [242, 780]. 1854 [1057]. 1859 [705]. 18th [492]. 1900 [693, 440, 966, 575, 1015]. 1905 [1119]. 1910 [782]. 1911 [596]. 1914 [993, 1084]. 1919 [824]. 1926 [781]. 1939 [782]. 1940 [886, 1110]. 1944 [121]. 1946 [24]. 1947 [597]. 1950 [622, 622, 480, 932]. 1950. [411]. 1951 [403, 403]. 1953 [594]. 1954 [648]. 1955 [999, 977]. 1956 [876, 832]. 1957 [1040]. 1959 [1058, 1123]. 19de [38]. 19th [340, 267].

200th [430].

31st [268]. 37th [151].

4 [442].

6 [617].

8512 [656].

9 [367].

A.D [513, 305, 791, 927]. A.D. [207, 389, 214]. A.H [349]. Abbildungen [694]. Abbot [1121]. Abbott [672]. Abdu’r [859]. Abdu’r-Rahman [859]. Abel [1003]. Aberdeen [1052]. Abetti [483, 483, 170]. abgehaltenen [119]. Abich [211]. Abraham [19, 365, 344, 503]. absolus [169]. Absorption [655]. Abu [899]. Abu’l [859]. Abu’l-Husayn [859]. Abundant [1046]. Abuse [409]. ac [646]. Academies [920]. Academy [494, 181]. Acare [788]. Accademia [23]. Acceptance [888]. accidents [850]. Account [320, 649, 944, 80]. Accused [448]. Acids [267]. Ackerknecht [810, 1073, 549]. Aconcio [758]. Acosta [1090]. Acta [931]. Action [685, 749]. Activities [803, 886]. Acute [206]. ad [646, 976]. Adam [1116]. Adams [4, 1085]. Adamson [695]. Address [385]. Adelard [420]. Adiabatic [943]. Administrationibus [858]. Administrative [488, 412, 550, 680, 937, 952, 1059, 454, 798, 98]. Admirable [975]. Adolf [211]. Adriano [1000]. Advances [267]. Adventures [257, 1121]. Aesthetic [709]. Aether [781, 371]. Aetii [207]. Aëtios [207]. aevi [291]. Affective [318]. Africa [917]. against [786]. Agassiz [580]. Age [432, 912, 1046, 1108, 305, 1108]. Ages [927, 434, 435, 593]. Agnes [131, 484]. Agnus [612]. Agricola [1046, 976, 977, 976, 154, 1046]. Agricole [784]. Agricultura [725]. Aim [721]. Aimant [559]. Air [829, 884]. Al [1018, 17, 161, 818, 197, 54, 285, 1064, 899, 469, 349, 899, 285]. al- [17]. al-Biruni [899, 349]. al-Bitrûjî’s [197]. Al-Bitruji [469]. al-Kāshī [285]. al-Kāshī’s [54]. Al-Kashi [1018]. al-Manāteq [54, 285]. Al-Qanunu’ [899]. al-Shātir [1064]. al-Tahafut [818]. Alan [1119, 405, 475]. Alard [145]. Albert [534, 95, 473, 930, 173]. Albertus [816, 941]. Alberuni [1093]. Albumasar [524]. Alchemical [828]. Alchemie [923]. Alchemist [104, 547]. Alchemists [83]. Alchemy [34, 602, 994]. Alchimistes [994]. Alessandro [1045]. Alex [893]. Alexander [705, 361, 650, 448, 823, 207, 364, 653, 670, 51, 1052]. Alexandre [973, 1116, 821]. Alexandrian [364, 774]. Alexis [36]. Alfred [670, 1082, 583]. Algorismus [755]. alicuius [1032]. aliorumque [291]. Alison [993]. Alkalis [267]. Allan [477]. Allen [401, 376]. Almagest [468]. Almirante [486]. along [102]. Alphabet [512]. Alsos [73]. Alten [77]. Altimetria [804]. altindischer [931]. Alum [946, 32]. Alwin [308, 476, 578]. Amateurs [96, 81]. Amber [257]. America [864, 605, 676, 539, 584, 650, 630, 125, 27, 31, 1100, 1025, 1055, 258, 641, 182]. Américaines [608]. American [574, 589, 991, 896, 376, 779, 80, 1026, 379, 534, 782, 514, 363, 569, 244, 340, 419, 222, 603, 102, 1090, 141, 496, 408, 782]. Americans [1113]. ami [729]. Amiable [71]. Amida [207]. Amideni [207]. Ammoniak [578]. Ammoniaksynthesc [308]. Amos [425]. Anaesthesia [671]. Analogies [6]. Analysis [315]. Analytic [1103]. Anastasia [36]. Anatomical [858]. Anatomicis [858]. anatomische [792]. Anatomy [311, 1115]. anciens [796, 813]. Ancient [160, 89, 432, 661, 34, 972, 246, 247, 865, 248, 673, 1107, 1094, 811, 185, 946, 959, 586, 642, 172]. Anderson [437]. Andre [132]. Andreae [38]. Andreas [751, 853, 152, 64, 555]. Andrew [87, 244]. Anecdotes [879]. Aneurysmatibus [506]. Aneurysms [506]. Anfängen [219, 822]. Angel [1074, 1111]. Angels [735]. Aniline [1024]. Anima [423, 544]. animae [816]. animalibus [816]. Animalium [1112]. Animals [640, 640]. Anlass [119]. Anneliese [154, 63, 906, 351]. Anniversary [430]. ano [297]. Anonyma [381]. Anonyme [1032]. anonymous [1032]. Answers [57, 148, 347, 948, 201, 274, 13, 988, 111, 288, 393, 452, 561, 633, 715, 606]. Antecedents [1008, 795]. Anthony [187, 256]. Anthropologist [549]. Anthropology [7, 539, 312]. Antiquary [157]. Antique [990, 59]. Antiquities [364]. Antiquity [773, 362, 342, 996, 1043, 301, 302, 856]. Antoine [1116, 906, 731, 52, 227]. Antoinette [123]. Antoni [1079, 25]. Ape [1111, 434]. Apes [735, 434]. Apocrypha [78]. Appearance [314]. Applications [783]. Applied [1055, 743]. Appraisals [1063]. Appreciation [865]. Approach [756, 983]. April [385]. arabe [930]. Arabic [420, 8, 767, 846, 998]. Arabs [30]. Arabus [291]. Aram [504]. Aratus [8]. Arbeit [122, 853]. Arbeiter [935]. Arbeitsmedizin [935]. Arber [131, 484]. Archaeologists [362]. Archaeology [1122]. Archer [381]. Archibald [404, 768]. Archimedes [283, 339, 1089, 499, 929]. Architect [143]. Architecture [408]. Archives [630]. Area [327]. Arifmetiki [1018]. Aristeas [365, 365]. Aristotelem [544, 115]. Aristoteles [377]. Aristotelis [900, 544]. Aristotle [723]. Arithmetic [1018]. Armando [676]. Armstrong [757]. Arnaldez [990]. Arnold [690]. Arranged [373]. Arrhidaei [1096]. Arrived [1046]. Art [33, 366, 964, 43, 756, 373, 535, 435, 790, 401, 92, 790]. arte [862]. Artelt [540]. Arthur [173, 896, 907, 37, 158]. artigos [486]. Arts [709, 727]. ärztlichen [219, 410, 822]. As-Sufi [859]. Asclepiades [997]. Ascribed [389]. ash [883]. Ashbel [320]. Ashley [186]. Ashworth [565]. Asia [304]. Asiatic [215]. Aspect [1034]. Aspects [603, 1038, 1055, 536]. Assaying [351]. Astrolab [106]. astrolabe [106]. Astrolabes [901]. Astrological [998, 593, 743, 775, 914, 1011]. Astrologorum [609]. Astrology [615, 724, 666, 724]. Astronomer [325, 1054]. Astronomers [340]. Astronomiae [21]. Astronomical [662, 803, 998, 1075, 55, 743, 775, 914, 1011, 967]. astronómicas [19]. Astronomie. [369]. astronomische [925]. Astronomy [197, 603, 713, 511, 974, 483, 992, 171, 533]. at-Tūsī [200]. Athens [363]. Atmosphere [406]. Atom [255]. Atomic [255, 741]. Atoms [255]. Attainments [179]. Attempt [684]. Attitude [709, 678]. attraverso [70]. Attributed [200]. Aubrey [431]. Aubry [729]. Audience [281]. Audubon [376]. Auerbach [702]. Aufenthaltes [118]. Aufklärung [410]. Aufsätze [1022]. Augé [784]. Augé-Laribé [784]. August [453]. Augustin [59]. Augustine [513]. Augusto [928]. Aurele [975]. Aurelianus [206]. Aurorae [199]. Ausbau [962]. Auseinandersetzungen [962]. Ausone [271]. Austryn [969]. Auteur [1032]. author [1032]. autobiographiques [357]. Autobiography [442, 26, 301, 397, 158]. Autocrat [573, 71]. Autolycus [205]. Autopsy [108]. autour [60, 594]. aux [505]. avant [60]. Averroès [18, 818, 818]. Averrois [900, 115, 544]. Avian [36]. Avicenne [1076]. Avstryn [115]. Awakening [690]. Award [940]. Azarquiel [208]. Azevedo [125].

B [1081, 760, 732, 510, 1018, 748, 514, 152, 486, 172, 215, 424, 690, 594, 820, 1103, 788, 253, 662]. B.C [927]. B.C.E [250]. Bá [162]. Babson [116]. Babylonian [967, 1039]. Back [46, 99, 138, 190, 220, 263, 277, 321, 333, 383, 414, 455, 489, 519, 551, 599, 626, 651, 659, 681, 707, 738, 761, 770, 800, 826, 835, 877, 909, 938, 953, 979, 1005, 1027, 1060, 1087, 1125]. Background [926, 663, 782, 805]. Bacon [67, 418]. badatel [760]. Bagdadien [462]. Bailey [87, 703]. Bailly [428, 1054]. Bainton [595, 594]. Baker [503]. Bakos [1076]. Ball [151]. Ballistics [472]. Balz [473]. Band [219, 822]. Baneth [115]. Bành [162]. Banks [573]. Baptista [1020]. Baranowski [999]. Barber [481]. Barbour [720]. Barcelona [105]. Barcroft [597]. Barnett [1082]. Barock [1114]. Barraud [813]. Bartolomé [354]. Bartomeu [1044]. Basis [166, 889]. Bassal [725]. Bastholm [259]. Bates [786, 217]. Bath [420]. Battles [62]. Baumer [532]. Bayard [210]. Bayley [380]. Bd [410]. Beaglehole [871]. Beall [697]. Beaujeu [990]. Beaujouan [990]. because [1046]. Beccaria [928]. Beck [699]. Becker [1053, 368, 594]. Beddoes [671]. Beeckman [645]. Beer [573, 992]. Beers [864]. before [376, 1008, 1093]. Beginn [822]. Beginning [915, 789]. Beginnings [782, 33]. Begründung [410]. Behavior [318]. Beierlein [861]. Being [186]. Beiträge [515, 923]. Belief [780, 780]. Beliefs [250]. Bell [779]. Ben [485, 545]. Benediktinerklosters [755]. Benefactor [239]. Beneficiaries [814]. Bengt [21]. Benjamin [649, 574, 430, 26, 397, 427, 67, 50, 807, 548]. Bennett [500]. Benson [478]. Benzi [294]. Beobachtungen [145]. Bequest [815]. Bergbau [976]. Bergbüchlein [387, 863]. Bergh [818]. Bergson [852]. Bergwerk [154]. Bergwerk- [154]. Beringer [855]. Berkeley [399, 382]. Bermannus [976]. Bern [1046, 905]. Bernal [619, 887]. Bernard [1116, 34, 548, 796, 1100, 904, 1051, 80, 378, 400, 975, 401, 256, 481, 244]. Bernardo [125]. Bernhard [816]. Bernheimer [435]. Bernier [7]. Bernoulli [916]. Berriman [543]. Berry [691]. Bett [752]. Better [1046]. Betts [618]. Betty [483]. between [214, 1115, 267, 388, 684]. Beyer [255].

Beyond [704]. Bézout [882]. Bibliografia [999]. bibliografisk [22]. bibliografisk-boghistorisk [22]. Bibliographers [319]. Bibliographical [985, 52, 381]. Bibliographie [373, 693, 566]. Bibliographies [109]. Bibliography [459, 356, 658, 834, 951, 1086, 507, 458, 731, 16, 136, 229, 276, 332, 453, 847, 769, 356, 998, 1106]. Bibliotheca [851]. Bibliothèque [295]. Bicentenary [829]. Bidney [539]. Bio [356]. Bio-Bibliography [356]. Biographers [527]. Biographical [320, 196]. Biography [71, 135, 991]. Biological [80, 889]. Biologie [375, 211]. Biologist [484]. Biology [893, 485, 176]. Bird [693]. Birds [894, 601]. Birmingham [881, 698]. Biruni [899, 349]. bis [962, 377, 694, 219, 410, 822]. Bishop [399, 336, 300, 1099]. Bitrûjî’s [197]. Bitruji [469]. Björn [91]. Black [829, 884]. Blair [475]. Blake [27]. Bloch [990]. Blood [522, 629, 199, 350, 1118]. Blüh [783]. Blumenthal [33]. Blunt [373]. Boas [538]. Bober [593]. Bocheński [586]. Bock [888]. Bodenheimer [446, 545]. Bodies [1062]. Body [424]. Boerhaave [655, 632]. Boethius [586]. boghistorisk [22]. Bogtrykkeri [22]. Bolgar [814]. Bologna [235, 1011]. Boltzmann [873]. Bonaparte [729]. Bondo [366]. Boner [786]. Bonet [62]. Bonnet [357]. Book [815, 967, 978, 1000, 369, 485, 705, 320, 1041, 1118, 35, 177, 187, 749, 1101, 299, 590, 733, 443, 649, 1049, 1070, 541, 1119, 160, 502, 567, 678, 973, 1020, 214, 637, 81, 159, 499, 607, 638, 751, 853, 890, 961, 962, 1071, 1081, 596, 781, 358, 1053, 153, 156, 240, 574, 864, 933, 121, 183, 260, 442, 589, 737, 760, 787, 872, 875, 1080, 903, 19, 238, 294, 430, 447, 469, 513, 643, 755, 814, 900, 929, 969, 970, 971, 42, 478, 586, 1116, 26, 27, 31, 79, 154, 173, 184, 254, 255, 356, 367, 397, 399, 403]. Book [474, 479, 516, 573, 595, 622, 793, 794, 851, 854, 991, 993, 995, 1079, 408, 896, 376, 377, 693, 894, 591, 857, 998, 1075, 63, 996, 89, 642, 432, 1109, 568, 1002, 1085, 34, 730, 162, 311, 874, 120, 122, 165, 308, 361, 476, 515, 578, 620, 1024, 972, 409, 477, 507, 548, 619, 697, 719, 747, 242, 780, 38, 166, 704, 306, 597, 722, 353, 650, 1073, 30, 78, 246, 247, 250, 251, 292, 317, 365, 598, 259, 1045, 1019, 720, 504, 796, 906, 968, 1021, 1042, 732, 735, 97, 677, 41, 127, 150, 186]. Book [236, 304, 441, 444, 445, 448, 510, 674, 779, 1100, 893, 73, 65, 429, 731, 850, 904, 1051, 210, 211, 249, 696, 679, 823, 21, 22, 68, 241, 297, 483, 615, 866, 1046, 1121, 778, 869, 891, 926, 1022, 1023, 1057, 80, 359, 427, 886, 1026, 888, 1120, 783, 873, 1038, 1039, 71, 44, 425, 378, 379, 404, 439, 440, 534, 621, 691, 736, 782, 924, 963, 1025, 1055, 170, 95, 128, 134, 237, 1052, 1111, 446, 819, 676, 1043, 117, 501, 570, 572, 754, 865, 867, 1115, 623, 1018, 546, 539, 33, 24, 69, 472, 473]. Book [532, 699, 821, 295, 319, 354, 400, 517, 734, 750, 753, 907, 169, 934, 135, 576, 752, 895, 82, 94, 163, 167, 168, 315, 316, 370, 371, 426, 670, 692, 748, 852, 931, 1114, 23, 508, 514, 423, 928, 675, 817, 209, 728, 123, 407, 83, 360, 1044, 37, 130, 152, 158, 188, 189, 305, 312, 318, 366, 538, 695, 703, 816, 898, 486, 964, 975, 994, 1076, 90, 644, 1009, 1040, 125, 64, 351, 646, 792, 797, 72, 1110, 213, 133, 20, 39, 40, 43, 93, 234, 296, 300, 350, 352, 401, 471, 506, 537, 540, 756]. Book [791, 923, 997, 1048, 1078, 1112, 96, 512, 613, 966, 1105, 901, 925, 88, 500, 543, 131, 124, 1084, 639, 487, 1047, 930, 727, 935, 999, 1001, 1050, 239, 119, 17, 18, 25, 28, 32, 59, 60, 61, 62, 66, 70, 74, 76, 91, 115, 116, 151, 155, 157, 161, 172, 205, 206, 207, 208, 215, 248, 252, 291, 293, 298, 301, 302, 307, 313, 348, 349, 357, 363, 364, 424, 431, 433, 480, 505, 544, 545, 565, 566, 608, 609, 640, 645, 673, 690, 725, 729, 789, 45, 235, 262, 470, 701, 702, 810, 813]. Book [858, 862, 965, 1017, 1077, 84, 85, 990, 1074, 855, 892, 902, 355, 547, 647, 575, 824, 671, 698, 721, 759, 905, 1099, 132, 757, 758, 1003, 1083, 700, 549, 569, 594, 861, 863, 976, 977, 174, 362, 428, 820, 927, 1102, 612, 1106, 897, 468, 511, 592, 593, 611, 724, 992, 1107, 564, 67, 870, 1117, 171, 118, 164, 245, 368, 398, 1054, 1103, 726, 402, 610, 723, 788, 694, 542, 795, 1016, 92, 212, 219, 410, 411, 577, 822, 75, 859, 503, 616, 77, 899, 533, 936, 1122, 1108, 29, 614, 860]. Book [932, 261, 314, 672, 811, 1072, 812, 617, 887, 216, 818, 303, 373, 380, 381, 382, 535, 790, 1113, 648, 871, 974, 36, 86, 87, 126, 129, 175, 176, 178, 179, 180, 181, 182, 217, 218, 243, 244, 253, 256, 257, 258, 309, 310, 372, 374, 375, 405, 406, 434, 435, 436, 437, 438, 475, 481, 482, 484, 509, 536, 587, 588, 618, 641, 784, 785, 786, 809, 825, 889, 1015, 1056, 1082, 1104, 185, 868, 1067, 571, 856, 618]. Bookbinding [767]. Booke [151]. Bookman [354]. Books [693, 48, 373]. Boorstin [1113]. Borgerliga [571]. Boring [442, 183]. Boris [398]. Borodin [825]. Botanical [221, 373]. botanische [373]. Botany [744]. Bouligand [169]. Bourgey [990]. Boussole [529, 559]. Bovier [796]. Boyd [359]. Boyer [1103]. Boyle [101, 655, 1010, 463, 282]. Boyle’s [557]. Bradwardine [817, 971]. Brahe [21]. Brahes [22]. Brain [470]. Brasil [125]. Bread [438]. Breaking [240]. Breeding [309]. Brière [932]. Briggs [248]. Brinton [184]. Britain [1101, 242]. British [573, 223, 43, 75]. Britten [1104]. Broda [873]. Brodin [612]. Broglie [596, 254]. Brooks [509]. Brown [385, 88]. Browne [238, 534, 744, 237]. Brucker [428]. Brunhes [228]. Bruno [236, 444, 702]. Brussels [152]. Brutus [150]. Bryan [359]. Bryant [163, 692]. Bubonic [1041]. Buch [28]. Buchanan [693]. buditel [760]. Builder [27]. Building [408]. Bulfinch [3]. Bunsen [120]. Burckhardt [1019]. Burden [246]. Burgess [361]. Burial [554]. Burns [508]. Burr [896, 483]. Burrows [1054]. Bütschli [143]. Butterfield [359, 79, 427, 134].

C [358, 153, 513, 643, 793, 854, 780, 650, 317, 1019, 546, 517, 337, 305, 512, 424, 1083, 897, 871, 126, 179, 372, 374, 809, 1042, 927]. C.E. [250]. ca [791, 410]. Cabanis [1021, 1021]. Caelius [206]. Caerleon [325]. Cairns [44]. Calculator [495]. Calculus [962]. Calendar [774]. Calendars [247]. Caley [897]. Caliphate [306]. Callahan [185]. Caloric [943]. Calvert [243]. Calw [863]. Camera [993]. Cameron [573]. Campbell [360, 791, 239]. Canonistes [60]. Cape [374]. Capillary [5]. Captain [964, 871]. caractéristiques [850]. Cardano [501]. Carl [363, 855, 1103, 1053, 72, 1083, 475]. Carlo [417]. Carmody [469, 998]. Caro [353]. carteggio [23]. Cartesianism [1012]. Cartographical [676]. Cartography [676]. Cartwright [671]. Carugo [1000]. Caryl [313]. Casas [354]. Case [163, 786]. Caspar [68]. Cassirer [358]. Casson [1107].

Castellion [594]. Castus [612]. Catalogue [116, 373, 790, 593]. Catalogus [609]. Catesby [468]. Catholics [787]. Cattell [121]. causa [971]. Causalités [850]. causis [976]. Cave [506]. Ce [654]. Cecil [587]. Celebration [838]. Celebrations [24]. Cell [266]. Cell-Division [266]. Cellier [33]. celorum [469]. Census [807]. Centennial [1024, 1040]. Centuries [1106, 806, 509]. Century [733, 403, 993, 250, 1055, 472, 300, 284, 157, 207, 647, 492, 614, 648, 786, 1082, 804, 184, 340, 54, 285, 282, 767, 496, 267, 582, 710, 686, 105, 270, 556, 1082, 995, 619, 728, 500, 789, 235]. Certain [380]. Cesi [23]. Ch [990]. Cha [55]. Cha-ma-lu-ting [55]. Challenge [786]. Chalmers [165]. Chandra [972]. Changing [765]. Chapin [398]. Chapters [165, 1009]. Character [802]. Charge [692]. Charles [1041, 35, 567, 573, 595, 361, 242, 65, 823, 1121, 534, 152, 644, 32, 357, 424, 565, 673, 470, 858, 758, 927, 176, 1082, 193, 3, 703]. Chart [676]. Charts [871]. Chaucer [894]. Chauncey [320]. Chauntecleer [1065]. Check [497]. Chemical [474, 584, 605, 6, 165, 404, 534, 782, 1055, 767, 959, 741, 32]. Chemiker [1023]. Chemist [104, 398]. Chemistry [622, 34, 972, 440, 691, 1055, 632, 282, 604, 915, 972, 926, 439, 963, 83, 360]. Chesney [1119]. Chester [309]. chez [813]. Chhattisgarh [366]. Chicago [45, 408]. Chief [572]. Chimie [52]. China [677, 77, 860, 614]. Chinese [932, 142, 284]. Ching [249]. Chirurgien [341]. Chrétiennes [228]. Christiaan [155]. Christian [780, 611]. Christianity [134]. Christman [623]. Christoph [855]. Christopher [298]. Chromatography [736, 736]. Chronic [206]. Chronicle [1119]. Chronology [251]. Church [969]. Chymico [1010]. Chymist [557]. Cicero [150, 554]. Ciencia [843, 76]. científico [753]. Cina [70]. circa [1045]. Circle [324, 240]. Circulation [1118, 683, 5, 144]. Circumference [1018]. City [1053]. Civil [955]. Civilisation [677, 860, 132]. Civilization [479, 1019, 16, 136, 276, 332, 453, 640, 642, 95, 539]. Civilizations [927]. Civilized [917]. cività [161]. Cl. [106]. Clagett [675, 856]. Claire [768]. Clark [154, 637]. Classic [249, 1046]. Classical [814, 685, 691, 1043, 363]. Classification [459]. classique [795]. Claude [803, 930, 400]. Claus [693, 1106, 373]. Clio [813]. Close [385, 1013]. Close-Hauled [1013]. Closed [973]. Clow [404, 927, 404]. Clowes [20]. Cloyne [399]. Club [1047]. Coal [942]. Coast [102]. codici [928]. Codicum [609]. Codification [133]. Cohen [1051, 378, 548, 127, 904, 80, 401]. Coherence [117]. Coiter [471, 862]. Cole [1117]. Colectanea [486]. Collected [1045, 25, 1079]. Collection [116, 790]. Collections [80]. Collegio [70]. Collinder [91]. Collision [90]. Colonial [663, 1113]. Colony [728]. Color [655, 522]. Colorado [1002]. Colton [791]. Comenius [425]. Comet [389]. cometa [297]. Comets [214]. Comfort [893]. Commemorating [1024]. Commemoration [349]. commensurabilis [1032]. commensurable [1032]. Commentaries [556]. Commentariorum [544]. commentariorvm [115]. Commentarium [544]. commentarius [797]. Commentarivm [900]. Commentary [348, 679, 797]. Commentatio [872]. Comments [188].

Commission [955]. Committee [1079, 514]. Common [752, 253]. Company [298, 820, 374]. Comparative [1101, 695]. Comparison [226, 267]. Compass [487]. compendium [291]. complètes [155]. Complexionibus [984]. Compression [943]. Computer [54, 285]. Computus [571]. con [19]. Conant [163, 692, 253]. Concept [692]. Concepts [670]. concerning [665, 572, 464]. Concise [171]. Conducted [363]. conferencias [486]. Conflict [126, 189]. Confucius [29]. congrès [596]. Congress [876, 832, 528]. Connaissance [269]. Conquered [320]. Conquest [568, 752, 752]. Conrad [1057]. Consequence [586]. Conservation [342]. Consideration [179]. Construction [408]. contemporaine [443]. contemporaines [123]. Contemporary [1026, 128, 408]. Contribution [159, 142, 404, 269, 1106]. Contributions [229, 45, 315]. Controversy [86]. Conventional [637]. Conway [84]. Cook [871, 372, 964]. Coolidge [81]. Coopland [62]. Copernican [883, 974, 572, 883]. coperniciennes [1050]. Copernicus [644, 465, 956]. Copies [226]. Copper [789, 820]. Cordvbensis [900]. Corinth. [363]. Corn [641]. Corner [26]. Corpus [500, 291, 544]. Corpvs [115]. Correspondence [989, 1036, 113, 885, 777, 1097, 833, 667, 689, 918, 331, 498, 560, 230, 808, 146, 56, 960, 1068, 12, 585, 714, 746, 451, 290, 392, 273, 467, 1014, 634, 531, 203, 950]. corrvptione [900]. Cortes [216]. Cortesao [676]. Cosmic [983, 592]. Cosmimetria [804]. Cosmografía [819]. Cosmography [224]. Cosmology [1038]. costae [1032]. Cotton [697]. couchers [205]. Coulomb [692]. Coulson [780, 213]. Coulston [242]. Count [574, 1034]. Counter [409]. Counter-Revolution [409]. Countries [1101]. Courses [232]. Courtier [398]. Coutinho [486]. Cow [187]. Cowan [1108]. Cox [75]. Craft [457]. Cramer [724]. Crane [184]. Crawford [1045, 544]. Creation [1038]. Creative [360]. Creel [29]. Crew [558]. Creyente [517]. Crime [726]. Crisis [700, 379]. Critic [709, 727, 700]. crítica [19]. Critical [459, 469, 658, 834, 951, 1086, 998, 458, 1026, 465, 109, 16, 136, 276, 332, 453, 616, 769]. Criticism [495]. critique [205]. Crombie [513, 643]. Crop [182]. Crosby [817]. Crucible [759]. Crucibles [34]. Cultivated [309]. Cultural [658, 834, 951, 1086, 235, 769]. Culture [305, 59, 419, 607, 246, 640]. Cultures [215]. cum [646]. Cuneiform [967, 656]. cura [1000]. Currents [532, 564]. Currier [507]. Curtis [898]. Curtius [93]. Curwen [587]. Cusanus [696]. Cylindro [283]. Cyril [154, 306, 351]. Czechoslovakia [533].

D [320, 1081, 1080, 619, 400, 692, 152, 350, 424, 547, 887, 126, 179, 809, 325]. Dīn [200]. D’Abro [370, 168]. Daguerre [993]. dall’alba [161]. Dalton [474, 741]. Dana [500]. Dandy [1104]. Daniel [382, 1113]. dans [930]. Danske [21]. Dantzig [815, 700]. d’après [1076, 930]. Darby [546]. D’Arcy [2]. d’Aristote [123]. d’Arles [545]. Darling [896]. Darlington [126, 809]. Darwin [735, 703, 1082, 193, 1093, 1082, 1082]. Darwinism [1093]. d’Astrologia [1044]. datae [646]. Date [389]. Dati [417]. Dating [743].

Daumas [730, 505]. David [1071, 864, 819, 539, 115, 1033, 508]. Davy [671, 281]. Dawes [485]. Dawn [317]. Dean [294]. Death [595, 33, 554, 84]. DeC [152]. Decade [782]. December [1058, 332]. Dechend [621]. Decimal [686]. déclin [169]. découverte [850]. Degradation [268]. Deity [670]. del [819, 753, 928]. Delanglez [42]. Deledalle [443]. D’Elia [70]. Delivered [385]. della [1049, 23, 161]. dell’Astronomia [170]. Demonology [435]. Denis [731, 260]. Denison [783]. Denkweise [751]. Denny [87, 244]. Department [955]. Departure [344]. Deposits [941]. depuis [596]. Derek [1020, 754]. Derivation [421]. Desargues [307]. Descarte [699]. Descartes [962, 504, 473, 699, 323, 473, 699]. Descobrimentos [486]. Description [21]. Descriptive [116]. Desgranges [169]. despotisme [1001]. Destructions [364]. Deutsche [1023, 925]. deutschen [120]. Deux [1001]. Devaux [852]. Development [408, 632, 692, 817, 1110, 602, 140, 974]. DeWitt [591]. d’Humanisme [295]. Diabetes [542]. Dialog [976]. dialogorum [291]. Dialogue [572, 572]. diameter [1032]. d’Ibn [1076]. Dibner [1046, 905]. Dichtung [210]. Dictionary [991]. Did [391, 418]. Diderot [665, 108, 664, 429, 504]. Dieckmann [429]. Diepgen [219, 410, 822]. Diest [145, 145]. Dietetics [142]. Differences [684]. Digby [870]. Digges [196]. Digressions [796]. Dijksterhuis [793, 929, 164]. Dikshoorn [793]. Dilemma [372]. Diller [431]. Dílo [872]. Dimension [1074]. Dimostrazioni [1000]. Dingle [403]. Direction [514, 990]. Discipulorum [875]. Discorsi [1000]. Discourse [700]. Discourses [975]. Discovered [1092]. Discoverer [401]. Discoveries [6, 575, 868]. Discovery [266, 1024, 830, 676, 766, 490, 805, 871, 1070, 165, 892, 1099, 924]. Disease [752]. Diseases [206, 752, 206]. Dissertationes [875]. Dissertations [233, 328, 396, 562, 636, 716, 848]. Distance [765, 685, 388]. Distances [285]. Distillation [33]. dit [654]. Divination [150]. Divis [391, 390]. Division [266, 834]. do [297, 125]. Doba [872]. Dobzhansky [129, 889]. Doctor [238, 325, 300, 549]. Doctoral [328]. Doctorates [563, 635, 668, 849]. Documentary [821]. Documented [235]. Documents [488, 412, 550, 680, 937, 952, 864, 1059, 454, 798, 98, 60, 1051]. Dodds [402]. Doig [171]. Domesday [546]. Domicile [227]. Dominant [1030]. Donald [595, 701, 758, 243, 820]. Donne [1115, 828]. d’optique [794]. Dordick [129]. Dorfman [934]. Dorothea [236]. Dorothy [96]. Dorwin [438]. Douglas [698]. Dow [360]. Dr [71, 846, 239, 229]. Dr. [872, 1010, 837]. Drabkin [206]. Drachmann [160]. Dragons [257]. Drake [572, 868]. Drama [615]. Draper [243]. Drawn [871]. Drayton [894]. Dream [348]. Dresden [690]. Dreyer [511]. Duane [163, 692]. Dubos [576, 576]. Due [1000]. Dugas [254, 795]. Duhem [970, 721]. Dujardin [788]. Dulieu [965]. Dumbleton [1091]. Dunn [237]. Dunnington [1083]. Dunsheath [403]. Du’O’Ng [162]. Dupont [990]. Dupont-Sommer [990]. Dupree [886]. Durand [500]. during [227, 1055, 865, 983]. Dürr [586]. Dutch [1079, 1110]. Dutch/Shell [1110]. Duties [150]. Duveen [731]. dyameter [1032]. Dye [942]. dyes [419, 1024]. Dynamic [1034]. Dynamics [695]. Dzhemshid [1018].

E. [537]. E.-H [411]. Earle [897]. Earliest [993, 103, 32, 533, 1015]. Early [557, 1109, 80, 676, 1063, 673, 611, 587, 1030, 779, 905]. Earth [892, 388, 874]. Earthquakes [50]. East [85]. Eastern [306, 546, 172]. Eastman [997]. Eaton [936]. eaux [411]. éclairé [1001]. Eclipses [270, 343, 806]. Écoles [228]. Economic [1055]. Economics [32]. Edelstein [997, 28]. Edgar [1078, 437]. Edgerton [28]. edición [19]. Editing [416]. Edition [469, 664, 226, 916, 205]. Editions [65]. Edmond [228]. Edmund [781, 515, 371]. Edna [703]. Eduard [622, 439, 244]. Education [763, 378, 491, 772, 379]. Edward [1101, 356, 991, 69, 364, 398, 535, 1025]. Edwards [50, 209]. Edwin [183, 442, 251, 359, 1054, 618]. eeuw [38]. Effect [797]. Effects [679]. effluunt [976]. Egg [36]. Egon [574]. Egypt [246, 247]. Egypte [1017]. Egyptian [246, 774]. Ehrenfried [638]. Ehrlich [299]. Eight [1093]. Eighteenth [733, 157, 648, 710]. Eighteenth-Century [157, 648, 710]. Eighth [876, 832, 332]. Eightieth [658]. Eighty [834, 951, 1086, 769]. Eighty-First [769]. Eighty-Fourth [1086]. Eighty-Second [834]. Eighty-Third [951]. einem [656]. Einheit [407]. Einige [419]. Einstein [173, 168]. Einzelwissenschaften [407]. Eiseley [1082]. ejusdem [1032]. Elder [783]. Eldridge [791]. Eleanor [507, 71]. Election [421]. Electric [1095, 692]. Electrical [525, 905]. electricitatis [797]. Electricity [781, 904, 679, 631, 692, 797, 371]. Electromagnetism [490, 805]. Elefanten [28]. Élémentaire [52]. Elements [736, 924, 756, 420, 583]. Elephant [28]. Elephant-Lore [28]. Elephant-Sport [28]. Eleventh [993, 767]. Elèves [882]. Elgood [306]. Eli [477]. Eliade [994]. Elis [21]. Elizabeth [757]. Elizabethan [615, 570, 523, 820]. Ellenberger [1074]. Elmer [179]. Eloi [965]. Elsa [376]. Elvira [534, 924]. Elwin [366, 366]. Emanuel [608]. Emerson [703]. émire [930]. Emmeram [755]. Emperor [249]. Emphasis [420]. Empire [248]. Empires [673]. Encyclopedia [344]. End [724]. Endeavour [65, 871]. Endorsed [834]. Energie [39]. Energy [255]. Engelbert [873]. Engine [280]. Engineering [174, 896]. England [728, 870, 936, 267, 492, 523, 1041, 570, 867, 546, 472]. English [590, 864, 184, 151, 671, 612, 983]. Enigma [216]. Enlightenment [358, 504]. Entretiens [796]. Entwicklung [377, 219, 410, 822]. Entwicklungsgeschichte [118]. Entwicklungsphysiologie [577]. Environmental [436]. eorum [976]. Ephemerides [967]. Ephrem [816]. Epicurus [591]. Epicycloidal [981]. Epidaure [813]. Epidemic [320]. Epidemiology [541]. épistémologie [123]. epistolae [646, 646]. Epistula [976]. Epitome [64]. Equations [285]. Equatorie [754]. Era [1077]. Eranos [40]. Eranos-Jahrbuch [40]. Ercker [351, 861]. Erect [391]. Eric [252]. Erik [851]. Erna [935]. Ernest [369, 872, 586, 891, 675, 1074, 575, 824]. Ernst [358, 93, 925]. Errata [1124]. Erratum [876]. errlee [29]. Errors [226, 490]. ersten [1023]. Erwin [377, 1073, 727, 810, 549, 610]. Escala [61]. Esche [792]. española [76]. especialmente [125]. ’Espinasse [902]. Essai [1017, 74, 965]. Essay [802, 831, 32]. Essays [615, 565, 535, 1120, 719].

Esther [39]. Estienne [757]. Estimate [9]. Estudios [76, 208]. étapes [850]. Ethan [401]. Ethics [750]. Ethiopian [1045]. Ethnocentrism [917]. Ethnology [653]. étude [74]. Etudes [60, 933, 608, 1050]. etwa [822]. Euclid [420, 583]. Euclide [159]. Eugene [1071, 930]. eum [646]. Europaeische [93]. Europe [1101, 574, 1055, 617]. European [1101, 215, 616, 391, 390, 733]. européenne [852]. Evangelista [737, 760, 875]. Evans [312]. Evans-Pritchard [312]. Evelyn [1010]. Eventful [534]. Evolution [1124, 461, 439, 168, 565, 178, 436, 850, 735, 37, 130, 129, 536, 1082]. Evolutionary [258]. Evolutionstheorie [375]. ex [976]. Exact [996, 302]. Examinateur [882]. Example [904]. Excavations [363]. Exhibition [373]. Existence [1074]. Experience [957, 1113]. Experiment [430, 805]. Experimental [183, 643, 995, 904, 400, 163, 1120]. expérimentateur [730]. Experiments [525, 103]. Exploration [1002, 221]. Exposition [1034]. extant [291]. Extraordinary [1003]. Eye [484]. Eyeglasses [417]. Eyre [398]. ’Ezra [19].

F [1041, 595, 361, 409, 317, 869, 963, 1055, 446, 473, 319, 895, 931, 83, 318, 20, 43, 300, 997, 1047, 363, 544, 545, 1074, 671, 564, 303, 185]. F.R.S. [986, 759]. Fabian [568]. Fabri [803]. Factors [129]. Facts [809]. Faculty [514]. Failure [982]. Fairy [380]. Fairy-Tales [380]. Faith [969, 238]. Fall [821, 673]. Family [31, 310]. Farber [622, 439, 750, 94]. Farbstoffe [419]. Farbtafel [694]. Farm [618]. Farming [587]. Farrand [397, 397]. Farrington [67]. Fate [51]. Father [390]. Fathers [969]. February [1058]. Federal [886, 182]. Federico [23]. Feldhaus [694]. Ferchault [906]. Ferguson [1084]. Fermat [962]. Fermentation [1040]. Fernel [65]. Fernow [244]. Fertility [372]. Festschriften [303]. Feuer [750]. Fever [320, 320]. Fifteenth [500, 54, 285, 806]. Fifteenth-Century [54, 285]. Fifth [16]. Fifty [932, 218]. Fighters [1107]. Filliozat [990]. Filosofia [639]. Filthaut [816]. Fin [59]. Finan [177]. Finch [238]. Finding [964]. Fine [693]. Firewalk [530]. First [391, 993, 664, 1066, 365, 49, 736, 141, 226, 326, 711, 1034, 981, 769]. Fischer [1057]. Fish [181]. Fisher [693]. Fission [34]. Fitch [408]. Five [534, 1106]. Fixed [829, 884]. Fleckenstein [1114]. Fleming [701, 243]. Fletcher [378]. Florence [876, 832, 1011]. Florian [755]. Florkin [1001]. Flour [1043, 438]. Flower [373]. Fobes [115]. Fockema [38]. Foley [679]. Folk [366]. Folk-Songs [366]. Folklore [380]. Fonds [429]. Fontenelle [796, 1012]. Foods [510]. Forbes [1110, 1109, 33, 174, 362, 811]. Forces [408]. Foreign [182]. Forestry [244]. Forgerons [994]. Forgotten [702]. Form [131]. Formal [586]. Formation [678, 74]. Formative [575]. Formerly [511]. Forms [382]. Forster [653]. Foster [590]. Fothergill [536]. Foulquié [443]. Foundation [494]. Foundations [255]. Founders [83]. Four [1095, 338, 185]. Fourteenth [270, 556, 806]. Fourth [1086]. Fr [696]. Frag [639]. Fragment [283]. fragmenta [291]. Fragments [592]. France [502, 267]. Frances [756]. Francis [469, 998, 67]. Francisco [297]. Franciscvs [115]. Francois [566, 7]. Frank [408, 823, 256, 315]. Franklin [397, 507, 904, 532, 28, 50, 649, 430, 655, 665, 597, 4, 631, 390, 807, 548, 904]. Frans [1017]. Franz [650, 433, 694, 538]. franzosischen [962]. Fray [1090]. Frederic [381]. Frederick [896, 724, 617]. Free [576]. Freedom [508, 889]. French [804, 227, 110, 1012, 269, 559, 491, 1032]. Frère [228]. Frères [228]. Freud [575, 824]. Freunden [122]. Friderike [124]. Friedrich [1083, 476]. Fritz [358, 593]. Front [1, 47, 100, 139, 191, 264, 265, 279, 322, 334, 384, 415, 456, 520, 521, 553, 579, 600, 627, 652, 660, 682, 708, 739, 762, 771, 801, 827, 836, 878, 910, 939, 954, 980, 1006, 1028, 1061, 1088]. Frontal [318]. Fulton [595, 319, 318]. Function [92]. Functionibus [872]. fundamentos [19]. Furon [990]. Further [257]. Future [94].

G [160, 183, 442, 573, 622, 376, 780, 1045, 1019, 1121, 425, 378, 170, 867, 623, 473, 934, 931, 964, 62, 215, 307, 813, 990, 927, 592, 29, 932, 648, 178, 405, 536]. G. [337]. Gaade [255]. Gabriel [228]. Gabriel-Marie [228]. Gabrieli [23]. Gago [486]. Galdston [541]. gale [788]. Galen [858, 629, 674]. Galeni [291]. Galilei [572, 1000, 978]. Galileian [338]. Galileis [63]. Galileo [1000, 513, 572, 70, 765, 879, 913, 945, 985, 1066, 324, 709, 107, 388, 956, 726, 868, 978, 727, 70]. Gallagher [614]. Galvani [679, 797]. Galveston [320]. Gambling [501]. Gand [341]. Gandz [657]. Ganzenmüller [923]. Garbers [17]. Garden [374]. Gardner [482]. Garraghan [42]. Gas [326]. Gaspare [235]. Gassendi [1116]. Gate [545]. Gauss [1083]. Gauthier [18]. Gaylord [130, 310]. Gear [981]. Geburtstages [119]. Gegenwart [377]. Geheimnis [39]. geistige [853]. Gene [428]. General [793, 3, 378]. Générale [990]. Generation [907]. generatione [900]. Generations [786]. Genesis [742, 242]. Genetics [405]. Geneva [503]. Genève [357]. Gennadi [181]. Genres [132]. Geoffrey [356]. Geograaf [66]. Geographers [431]. Geography [546, 1009, 89]. Geological [892]. Geologie [855, 976]. geologischen [855]. Geology [242, 582, 1072]. geometric [804]. Géométrique [804]. Geometry [1103]. Georg [120, 317, 976, 375, 976]. George [961, 447, 26, 367, 399, 642, 432, 674, 866, 865, 130, 310, 653, 712, 940, 1029, 740, 844, 839, 840, 842, 841, 847, 845, 843]. Georges [1116, 169]. Georgius [872, 154, 976, 872, 977]. Gérard [443]. Gerhard [736]. Germain [696]. German [419, 656, 106, 145]. germanice [646]. Germany [267]. Gernsheim [993, 993]. Gerson [545]. Gesammelte [241, 1022]. Geschichte [369, 962, 693, 515, 1073, 540, 923, 373, 638, 962, 308, 537, 855, 368, 219, 410, 822]. Geschiedenis [38]. Gestalten [124]. gesuiti [70]. Geyer [816]. Geymonat [978, 1000]. Ghellinck [60]. Ghotul [366]. Gibbs [245]. Gibson [35]. Gilbert [42, 1047, 418, 1047]. Gillam [759]. Gillispie [242]. Gino [161]. Ginsburg [958]. Giordano [236]. Giorgio [483, 572, 726]. Giovanni [506]. Gisela [790]. Giuseppe [23]. Given [21]. Giyaseddin [1018]. Glasstone [255]. Glauber [869]. Gleanings [879, 913, 945, 985, 1066]. Glenn [644]. Glory [364]. Glossary [1072]. Glover [679]. Gnade [1057]. Gnudi [235].

God [670]. Gods [150, 445]. Goethe [210, 210, 547, 212]. Goethe. [210]. Goethezeit [211]. Gold [268]. Goldammer [352]. Golden [432, 703]. Goncalves [639]. Gone [117]. Goodman [314]. Goodrich [514]. Gordis [292]. Gordon [971, 1052, 353]. Gorilla [311]. Gösta [612]. Gottfried [119, 1114]. Gouby [702]. Goudsmit [73]. Government [156, 886, 182]. Grace [116]. Graeco [611]. Graeco-Roman [611]. Graecorum [609]. Graham [759]. Grain [1043]. Grain-Mills [1043]. Grant [611]. Granted [864]. Graphic [1008]. Grassi [407]. Gravitation [9]. Gray [986, 525, 526, 547]. Great [320, 177, 1101, 81, 242, 572, 319, 575, 245]. grecque [74]. Greece [432]. Greek [30, 431, 790, 457, 856]. Greeks [692, 402, 815, 968, 610]. Green [997, 797]. Gregor [192, 1056]. Gregory [311]. Grenzen [407]. Grete [620]. Gridley [896]. Grosse [124]. Grosseteste [643, 1094]. Grundproblem [375]. Grundsätze [931]. Grunhaldt [906, 351, 154]. Gudger [229]. Gudmund [587]. Guerlac [479]. Gugel [869]. Guide [447, 42]. Guitard [411]. Gunnery [523]. Gustav [646]. Guterman [296]. Gutmann [872]. Guttmacher [997]. Gynaecology [207]. Gynecology [997]. Gypsum [959]. Gyro [487]. Gyro-Compass [487].

H [749, 595, 622, 697, 780, 166, 722, 1073, 78, 97, 891, 359, 427, 425, 1111, 570, 546, 539, 692, 514, 817, 360, 487, 151, 45, 701, 810, 990, 1099, 549, 569, 594, 927, 724, 411, 533, 29, 434]. ha-Shamayim [545]. Haber [1055]. habiologia [446]. Hadas [210, 365]. Hagberg [475]. Hagen [812]. Hahn [255]. Hälfte [1023, 410, 822]. Hall [1042, 927, 678, 472, 673]. Hamburg [119]. Handasyde [693]. Handbook [790]. Handlin [701]. Handschriften [755]. Hanke [354]. Hans [573, 851, 21, 778, 593, 261, 577]. Hanson [1070]. hänsyn [571]. Happened [195]. Hard [1062]. Harding [39]. Hare [752]. Harold [1117, 380, 786]. Harris [400, 348, 117]. Harrison [360, 1108, 894]. Harry [969, 593, 303]. Hartmann [1022]. Hartner [77]. Harvard [80, 840, 5, 163]. Harvey [1118, 683, 144, 1015]. Haskins [313]. Hatt [587]. Hatton [360]. Haudricourt [990]. Hauled [1013]. Haven [964]. Haven-Finding [964]. Hayek [409]. Haynes [782]. Hays [1111]. Hazard [733, 616]. Health [541]. Heart [629, 350]. Heat [655, 280, 1034]. Heath [499]. Heathcote [622]. Heaven [545]. Heavenly [1053, 644]. Hebrew [251]. Hector [573]. Hegel [44]. Heilkunde [219, 410, 822, 77]. Heinz [442, 212]. Helen [650]. Hellenic [364]. Hellenism [250]. Helmut [705, 993, 976]. Henri [159, 1074, 59, 700]. Henricvs [115]. Henrik [1003]. Henry [479, 854, 1085, 926, 440, 167, 213, 234, 28, 116, 262, 558, 947, 27]. Henryk [787, 999]. Heraclitus [592]. Herbal [612]. Herbals [177, 1106]. Herbarium [1104]. Herbert [442, 79, 154, 403, 441, 429, 731, 440, 736, 134, 360]. Heredity [85]. Heretic [595]. hérétique [594]. Heritage [814, 182]. Herman [632]. Hermann [82]. Hermetic [401]. Heron [160, 696]. Herrlinger [471]. Herschel [732]. Herskovits [538]. Hertha [621]. Hesse [736, 748]. Hewson [486]. Heytesbury [898]. Hickman [671]. Hieroglyphic [252]. Highlander [366].

Highway [812]. Him [116]. Hindu [972, 28]. Hindus [28]. Hipocratismo [734]. Hirsch [353]. Hirst [568]. Histoire [205, 411, 159, 654, 788, 254, 162, 293, 566, 990]. Historia [639, 843, 76]. Historian [844]. Historians [688]. Historic [165]. Historica [931]. Historical [944, 787, 42, 720, 926, 691, 32, 582, 543, 536]. Historie [151]. Histories [888, 163]. Historiography [433]. historique [852, 74, 965]. historische [219, 410, 822]. History [232, 396, 563, 562, 635, 668, 636, 716, 849, 848, 876, 921, 961, 183, 260, 442, 286, 447, 513, 658, 854, 832, 834, 993, 995, 951, 987, 1086, 450, 376, 432, 949, 165, 920, 972, 719, 306, 222, 722, 842, 650, 259, 880, 1042, 142, 483, 886, 534, 963, 676, 628, 669, 33, 821, 395, 752, 371, 670, 141, 496, 10, 305, 486, 964, 96, 512, 717, 630, 16, 136, 229, 276, 301, 332, 433, 453, 565, 640, 673, 262, 810, 1062, 841, 174, 820, 927, 1106, 511, 564, 1065, 745, 171, 245, 666, 542]. History [919, 860, 672, 381, 382, 535, 275, 176, 180, 182, 587, 769, 1015, 718, 385, 896, 837, 922, 624, 896, 1109, 134, 887, 217, 310, 1071, 781, 589, 516, 408, 89, 972, 78, 1117, 1103, 1016, 588]. Hitchcock [401]. Hobbes [324]. Hochmittelalters [124]. Hoebel [695]. Hofmann [962, 638, 118, 368]. Hohenheim [1048]. Holmes [71, 507]. Holmyard [1042, 994, 673, 927]. Homer [445]. Hone [239]. Honoré [62]. Honour [565]. Hooke [225, 386, 9, 902]. Hoole [53]. Hoover [154]. Hope [240, 379]. Hopkins [1119, 569]. Hornof [875]. Horse [310]. Horses [310]. Horus [447]. Hospital [1119, 3, 649]. Hospitals [936]. Howard [1024, 1025, 115, 648]. Howell [1072]. Huang [249]. Huber [355]. Hubert [150]. Hugh [45]. Hull [317]. Human [748, 424, 470, 889, 186, 907, 188, 372]. Humanism [319]. Humanisme [813]. Humanist [595, 337]. Humanitarian [401]. Humboldt [705, 653, 51, 705]. Humphry [281]. Hundred [485, 255, 1093]. Hundredth [704]. Hungry [786]. Hunted [595]. Hunter [886, 1016]. Huntington [44, 398]. Husayn [859]. Hutchins [487, 468]. Hutton [41, 891]. Huttonian [874]. Huxley [735, 85]. Huygens [155, 1080, 155]. Hyde [87]. Hydraulics [1016]. hydrologie [411]. Hyndman [436].

Iago [541]. Ian [1052]. Ibn [1064, 1076, 883, 725, 19, 613, 18]. Ichthyology [229]. Idea [539, 189]. Ideas [184]. Identification [55]. Idria [935]. Ignorance [696]. ihre [693, 373]. ihrer [369]. II [1119, 962, 1081, 781, 9, 1105, 927, 410, 822, 860, 913, 884, 285, 221, 559, 914]. III [1042, 945]. Illtyd [736]. Illuminated [593]. Illusion [1033]. Illustrate [864]. Illustrated [32]. Illustration [284, 743, 373, 373]. Illustrations [152, 874]. Illustrators [373]. Illustratration [850]. illustrierten [693]. Ilza [249]. im [63, 935, 694, 77]. Imagery [828, 336]. Imagination [748]. Immanuel [90]. Immunity [309]. Impact [1089, 222]. impetus [462]. Incarnation [969]. Ince [1016]. Incidentally [554]. Including [151, 1045]. Incoherence [818]. Incorporating [972]. Independent [391]. Index [303, 540, 179]. India [972, 366]. Indian [539, 305, 1090, 641]. indischer [28]. Industrial [942, 1042, 963, 710, 881, 67]. Industrialist [361]. Industry [619, 880, 782, 1055, 32]. ineditos [486]. inédits [429]. Infectious [752]. Infinite [973, 236]. Influence [1025, 9, 640, 569, 180]. Influences [658, 834, 951, 1086, 769]. Information [137, 278, 413, 518, 552, 625, 706, 799, 908, 1004, 1126]. Ingo [1056]. initiatives [60]. Inquiry [904, 51, 380]. Inscripcões [125]. Inscriptions [338]. Inspiration [335]. Instauratae [21]. Institute [116]. Institution [97, 281]. Instrumente [925]. Instruments [55, 21, 80, 505]. Insufficiency [9]. Intellectual [516]. Interaction [215]. Internal [249]. International [876, 832, 834, 140]. Interpretation [246, 664, 1102]. Interpretations [188]. intorno [1000]. introducción [19]. Introduction [793, 584, 605, 78, 631, 852, 74, 1077, 182]. Introductory [677]. Inventaire [429]. Invention [391, 280, 417, 747]. Inventions [672]. Irénée [59]. Iris [72]. Iron [4, 906]. Irrational [402]. Irvine [735]. Isaac [474, 794, 1051, 426, 116, 645, 386, 1051]. Isadore [129]. Isis [459, 458, 109, 48, 460, 416]. Iskandar [613]. Islamic [1075, 713, 901]. Israel [251]. Italy [1019]. Itard [990]. Itineraries [1045]. itr [17]. IV [250, 985, 976]. Ivan [175]. IX [928]. Izquierdo [734, 753].

J [177, 749, 733, 1020, 1081, 1080, 469, 929, 42, 793, 993, 998, 89, 1109, 1024, 619, 38, 720, 1042, 732, 441, 779, 691, 754, 33, 473, 699, 400, 734, 135, 94, 152, 538, 486, 994, 1110, 300, 997, 966, 1105, 66, 252, 363, 424, 673, 990, 647, 174, 362, 927, 511, 118, 164, 368, 398, 616, 1108, 614, 1072, 887, 381, 1113, 871, 36, 1104]. Jack [128]. Jackson [1117]. Jacob [255, 1019]. Jacobi [296]. Jacopo [758]. Jaffe [34, 1100, 927]. Jahrbuch [40]. Jahrhundert [63, 935]. Jahrhunderts [755, 1023, 925, 219]. James [890, 358, 156, 408, 693, 597, 163, 692, 791, 172, 207, 811, 253, 1104, 991, 104, 711, 121]. Jammer [670]. Jan [875, 1076, 737, 760]. Janson [434]. January [658, 834, 951, 1086, 769]. Japan [215]. Jarmila [737]. Jean [42, 65, 169, 576, 341, 428, 1054]. Jean-Sylvain [428, 1054]. Jeannette [477, 398]. Jefferson [359, 618, 912]. Jeffries [193, 461]. Jekuthiel [958]. Jenner [356]. Jensen [510]. Jeremiah [238]. Jerome [235, 509]. Jerusalem [528]. Jessop [399]. Jesus [444]. Jesus-Marie [444]. Jet [342]. Jet-Propulsion [342]. Jewish [250]. Jirí [872]. Joachim [778, 1114]. Joanne [646]. Joaquín [753]. Johann [869, 1057]. Johannes [241, 339, 68]. John [502, 1020, 153, 474, 595, 896, 1002, 874, 246, 425, 319, 318, 300, 239, 759, 897, 503, 438, 185, 802, 4, 828, 604, 764, 776, 1063, 492, 1091, 356, 243]. Johns [1119, 569]. Johnstone [615]. Jolande [296]. Jonathan [50, 209]. Jonckheere [1017]. Jones [45, 590, 575, 824, 1117]. Jordan [508, 316]. José [19, 819, 753, 61, 76, 208]. Josef [335]. Joseph [637, 638, 962, 573, 677, 783, 1090, 60, 205, 759, 860, 829, 884, 597, 213]. Josiah [75, 710, 245]. Journal [645]. Journals [614, 871]. Jr. [178]. Judah [251]. Judío [819]. Julian [81, 359, 85]. Júlio [639]. Julius [335]. July [24, 16]. Justus [621].

Kāshī [285]. Kāshī’s [54]. Kaerntner [1048]. Kai [613]. Ka’ius [613]. Kalendern [571]. Karl [586, 210, 1057, 17, 786]. Karsten [374]. Kartografie [38]. Kashi [1018]. Kate [1049]. Kawakib [859]. Keele [350]. Keevil [966, 1105]. Keilschrifttext [656]. Keith [37, 158]. Kemp [699]. Kenelm [870]. Kennedy [1075]. Kenneth [597, 888, 810]. Kepler [821, 511, 241, 68, 511]. Kester [903]. Keuning [66]. Key [1018]. Keynes [356]. Khmer [248]. Kilgour [896]. Killer [135]. Kimiya [17]. Kind [913, 945]. King [311, 156, 583, 854]. Kingdoms [251]. Kings [251]. Kirby [896]. Kirk [592]. Kisch [702]. Kitab [17]. Kleine [778]. Klibansky [291]. Klickstein [731, 440]. Kline [607]. Klose [182]. Klosterneuburg [500]. Klyuch [1018]. Knapp [514]. Know [418]. knowledge [269, 166, 315]. Known [829]. Knut [475]. Kocher [570]. Koelln [358]. Koheleth [292]. Kongelige [21]. Konstitutionslehre [822]. kopernikowska [999]. Kopparberg [789, 298]. Körner [1102]. Koyré [1116, 973, 821]. Kraus [291]. Krumbiegel [1056]. Kruta [875, 872]. Ktesibios [160]. Kuhn [974]. Kulturvolker [419]. Kurt [755, 869, 189, 352, 355]. Kurze [1073]. Kymograph [1008].

L [978, 499, 993, 568, 510, 427, 404, 1055, 1043, 532, 699, 64, 20, 300, 997, 901, 487, 690, 990, 511, 614, 617, 36, 899]. l-Mas [899]. Labat [990]. Laboratory [942, 49]. Lacy [30]. Lamar [817]. Lance [576]. L’Ancien [293]. Lancisi [506]. Land [1080, 256]. Langfeld [442]. Language [802, 891, 380, 590]. Laplace [882]. Lapp [153]. Lapps [91]. Laribé [784]. Larsen [574]. l’Artillerie [882]. Lateinisches [93]. Latent [280]. Latin [469, 998, 506, 593, 420, 1030, 775, 914, 214]. Latinarum [544]. latine [930, 60]. Laurence [932]. Laurent [731, 227]. Lauritz [22]. Lavoisier [731, 52, 227, 584, 605, 527, 763, 1031, 730, 934]. Law [1115, 695, 140, 611, 724]. Lawrence [31, 304, 248]. Lazarus [351, 861]. Leaders [702, 97]. Leaf [509]. League [373, 373]. Leake [320]. Learned [696, 48]. Leben [122, 869, 694, 577]. Lebens [219, 410, 822]. Lebensbild [120]. Leblanc [830]. Lecture [5]. Lectures [225]. Ledyard [178]. Leeuwenhoek [1079, 25, 25, 53]. LeFanu [356]. Lefebvre [990]. Leff [971]. Legal [695, 44]. Leibnitz [647]. Leibniz [1114, 464, 119, 355]. Leibnizschen [118]. Leicester [440, 926]. Lejeune [930]. Leon [135, 18]. Leonard [408, 163, 692, 936, 406, 196]. Leonardo [1049, 792, 350, 424]. Lesky [935]. Lessing [733]. lessner [29]. L’Essor [60]. Letter [525, 365, 386, 1035, 464]. Letters [986, 193, 1051, 25, 845, 380, 1079, 427]. leurs [123]. levers [205]. Levy [613]. Lewis [733, 354, 750, 616, 325]. Ley [257, 785]. l’Hôpital [965]. Libellus [1112]. Liber [816, 816]. Libertad [517]. Liberty [87]. Library [116, 364, 51]. Libri [207, 976]. Libro [19, 725]. libros [900, 544]. Lichtenberg [337]. Lieberman [250]. Liebig [621]. Liège [1001]. Life [705, 574, 238, 595, 526, 361, 236, 213, 997, 764, 776, 239, 235, 575, 824, 1083, 436, 437, 809]. Lightning [50, 391, 430]. Lilley [719]. Limbo [195]. Linceo [23]. Lind [64]. Lindroth [298, 480, 789]. Lindsay [128]. Ling [677].

Linnoeus [475]. l’Institut [228]. Lionel [1107]. Lippmann [515]. List [497, 65]. Literary [250]. Literatur [93]. Lithberg [571]. littérature [60]. Little [829]. Live [892]. Liverpool [45]. Living [436]. Livingston [991]. LL.D [759]. Lloyd [622, 88]. Lobotomy [318]. Locke [502]. Lockemann [120]. Lockwood [294]. Lodestone [487]. Loeb [353]. L’oeuvre [307, 307]. Logan [711]. Logic [637, 586, 898, 586, 637, 478, 586]. logique [123]. logiques [169]. logistique [123]. lointaine [60]. Lombard [60]. Lomonosov [398]. London [45, 373]. Longitude [54]. Longitudes [285]. Looks [917]. L’optique [930]. Lore [28, 434]. Loren [1082]. Loria [161]. Lost [1031, 380]. Lot [385]. Lough [502]. Louhenry [154]. Louis [156, 254, 965, 580, 576]. Lowell [81]. lu [55]. Luce [399]. Ludovico [1000]. Ludwig [1008, 28, 873]. Luigi [679, 797]. Lull [756]. Luminescence [1015]. Lunar [883, 881]. Lungs [629]. Luther [465]. Luz [216]. Lyman [359]. Lynam [535]. Lynde [245]. Lynkeus [335]. Lynn [214, 995, 940].

M [1119, 787, 19, 586, 993, 507, 150, 926, 71, 440, 819, 400, 152, 994, 234, 639, 70, 424, 820, 611, 790, 825]. M.D [1052, 45]. ma [55]. Maccabees [365]. Machines [905]. Macklem [1115]. Macomber [116]. Macrobius. [348]. Made [514]. Maggid [353]. Magic [995, 493, 234, 41]. Magick [1020]. Magister [145]. Magnet [1047]. Magnete [1047, 418]. magnum [544]. Magnus [816, 212, 941]. Mahoma [61]. Maier [63]. Main [532, 564]. Maize [177]. Major [722]. Majumdar [305]. Maker [174]. Malady [615]. Malaria [895, 135]. Malaysia [221]. Malthus [807]. Malthusianism [786]. Mammals [742]. Man [121, 292, 576, 695, 29, 437, 438, 825, 742, 539, 580, 445, 510, 895, 174]. Manāteq [54, 285]. Mangold [577]. Manners [250]. Manpower [1101]. Manufacture [942]. Manuscript [986, 496]. Manuscripts [593, 775, 914, 1011, 871, 612]. Map [500]. Mapmaker [535]. Maps [88, 535]. Marāgha [55]. Marcel [1001]. March [1123, 276]. Marcus [150]. Margaret [679, 902]. Margenau [167]. Margerie [608]. Maria [506, 124, 76, 208, 694, 366]. Marie [228, 444]. Mariners [1107]. Mariotte [269]. Maritima [639]. maritimos [486]. Marjorie [240]. Marquardt [299]. Marrou [59]. Marshall [675, 856]. Marston [408, 217, 786]. Martha [299, 235]. Martin [359, 482]. martyr [594, 595]. Marvin [598, 750, 94]. Marxism [128]. Mary [534, 924, 748]. Mas [899]. Maschine [694]. Mason [564]. Mass [116]. Massachusetts [3]. Massignon [990]. Master [145]. Mastery [895]. Matangalila [28]. Matematiche [1000, 161]. Material [829, 116, 303]. Materialism [94]. Mates [478]. mathematica [474]. Mathematical [421, 867, 370, 344, 817, 898, 775, 914, 1011, 656]. Mathematician [325, 1003]. Mathematico [1010, 169]. Mathematico-Chymico-Mechanical [1010]. mathématico-logiques [169]. Mathematics [81, 961, 497, 82, 890, 1071, 1081, 607]. Mathematik [962, 118, 638, 368]. mathématique [307]. mathématiques [159, 654]. Mathematische [751]. mathematischen [656]. Mather [697]. Matousek [760]. Matter [46, 99, 138, 1, 47, 100, 190, 220, 263, 277, 139, 191, 264, 265, 321, 333, 383, 414, 279, 322, 334, 384, 455, 489, 519, 415, 456, 520, 551, 599, 626, 521, 553, 579, 600, 651, 659, 681, 707, 627, 652, 660, 682, 738, 761, 770, 800, 708, 739, 762, 771, 826, 835, 877, 909, 801, 827, 836, 878, 938, 953, 979, 1005, 910, 939, 954, 980, 1027, 1060, 1087, 1125, 1006, 1028, 1061, 1088]. Matthew [614]. Maturity [824]. Maurice [596, 730, 505]. Maurolico [831]. Max [397, 68, 1022, 670]. Maxime [74]. May [136, 1074, 733, 487, 616]. Maya [252]. Mayer [335, 901]. Mayo [31]. Mayor [428]. Mayow [764, 776, 1063]. McCrensky [1101]. McFarland [598]. McGill [94]. McKeen [121]. McKelvey [27]. McNall [508]. McQueen [361]. Meaning [130]. Means [6]. Measurement [1108]. mécanique [254, 795]. Mechanica [21]. Mechanical [672, 1010]. Mechanics [793]. mechanisering [164]. Med [872, 571]. Medal [712, 940, 1029]. Medecine [162, 813]. Médecins [1017]. medica [862]. Medical [306, 5, 319, 604, 565, 1106]. medicina [928]. medicinales [207]. Medicinalium [931]. Medicine [1119, 622, 842, 325, 565, 262, 702, 1077, 457, 810, 722, 249, 400, 28, 701, 810, 569, 966, 1105]. médico [1001]. médico-social [1001]. Medieval [982, 294, 844, 1089, 586, 972, 865, 675, 898, 1030, 1065, 1011, 75, 303, 283, 420, 497, 775, 914]. Médiévale [990]. medii [291]. Mediterranean [927, 1107, 43]. medivm [900]. Medizin [931, 540, 219, 410, 822]. Meeting [385]. Meetings [11]. Meier [593]. Mein [620]. meinem [122]. Meister [577]. Melanchthon [1035]. Mellitus [542]. Melville [538]. Memoir [1031]. Mémoires [357]. Memoirs [906, 397]. Memoriam [143]. Memories [255]. Men [320, 184, 866, 480, 640, 435, 313, 1100]. Mendel [1056, 192]. Mensch [873, 40]. Menshutkin [398]. Meridian [704]. Merlan [857]. Merrill [179]. Merton [133, 1091]. Mesopotamian [946, 959]. Metallica [154, 1046]. Metallurgical [1046]. Metallurgy [789, 362]. Metals [1046]. Meteorologica [556, 687]. Method [699, 400, 42]. Methoden [962]. Methodology [344]. Methods [179]. Metrology [543]. Metropolitan [790]. Meurerum [976]. México [753, 734]. Meyer [647, 211, 907]. Meyer-Abich [211]. Michael [469, 595, 1115, 595, 595, 736]. Michaels [408]. Michel [990, 594, 784, 159, 594]. Michurin [175]. Microfilm [497]. Microphotography [103]. Mid [340]. Mid-19th [340]. midden [38]. Middle [366, 927, 612, 593, 434, 435]. Middlemore [1019]. Mighty [509]. Mikhail [1122]. Milan [876, 832, 1011]. Millás [19, 819, 76, 208]. Miller [728, 209, 1122]. Million [310]. Mills [1043]. Milton [894, 1038, 1039, 903]. Mina [359]. Mind [728, 616, 484, 473, 245]. minérales [411]. Mineralogical [80, 915]. Mineralogie [976]. Mines [820]. Mining [1046, 284, 298, 789]. Minor [431]. Miracle [611]. Mirror [613]. Mirsky [477]. Misch [317, 301]. Misiak [787]. missionari [70]. Misstatements [956]. Mitchell [747]. Mittasch [308, 476, 578]. Mitte [755, 219]. Mittelalter [93]. M’koroth [446]. Mnemonics [802]. Model [883]. Modern [942, 678, 637, 781, 79, 642, 7, 891, 1039, 691, 473, 94, 167, 83, 45, 701, 702, 617, 310, 372]. modernen [822, 375]. modernes [796]. Modernization [151]. Moerbeke [339].

Mogenet [205]. Mojmir [872]. Moments [1099]. Momentum [342]. Monde [970]. mondes [796]. Monge [307, 729]. Mongrédien [1116]. Montagu [186]. Montaña [753]. Montesquieu [733]. Montpellier [965]. Montraville [797, 997]. Moody [675, 586]. Moon [967, 601, 54, 388]. Moore [892]. Moorhouse [512]. Moral [1115]. Morgan [210]. Moritz [1043]. Morley [1025]. Morris [607, 127, 618]. Morton [405]. Morus [640]. Moses [365, 210]. Mosher [381]. motibus [469]. Motion [967, 54, 797, 679]. Motions [629]. Mottelay [1047]. motu [797]. motus [816]. Mouvement [60, 1001, 205]. Movement [223, 350]. movimento [753]. Much [1046]. Muhammad [899]. Mukerjee [92]. Muller [441, 931]. Mullett [1041]. Munchner [755]. Munitz [1038, 1039]. Muñoz [61]. Murder [366]. Muria [366]. Muris [339]. Murray [720, 360]. Muscle [259]. Muscular [679, 797]. musculari [797]. Museum [27, 141, 790, 573]. Muslim [433]. Musse [122]. Mysteries [1078]. Mysterious [251]. Mystic [1054]. Myth [1039, 29]. Mythological [593]. Mythologies [380]. Myths [640].

N [1070, 20, 300, 997, 398, 542, 180, 309, 825]. nach [755]. nama [613]. Name [482]. Names [380, 179]. Naming [69]. Nan [404, 927]. Napoleon [772, 729]. národní [760]. Nasīr [200]. Nash [163, 692, 406]. National [630, 373, 373]. Natur [28]. Natura [1112, 816, 976]. Natural [1124, 463, 242, 880, 1051, 823, 1115, 82, 141, 496, 131, 611, 1065, 217, 1082, 1020]. Naturalia [723]. naturalis [474]. Naturalism [504, 1026]. Naturalist [257]. Naturalium [931]. Nature [150, 167, 86, 217, 664, 610]. Nature-Nurture [86]. naturelle [293]. Naturforschers [120]. Naturphilosoph [476]. Naturphilosophie [63]. Naturwissenschaft [540]. Naturwissenschaften [515]. Nautica [486]. Nautical [676]. navegador [486]. Navigation [676, 964, 486]. Navy [966, 1105, 955]. Near [172]. Nederland [38]. Needham [677, 990, 860]. Needs [779]. neerlandicorum [862]. Neglected [526]. Nei [249]. Nelson [314, 182]. Nemec [760]. Neoplatonism [857]. Nervosi [872]. Netboy [256]. neuen [962]. Neugebauer [967, 996, 302]. Neuman [365]. neunzehnten [1023]. Nevins [477]. Newman [890]. News [1069, 1037, 1098]. Newton [474, 904, 1051, 168, 9, 116, 1015, 655, 794, 1051, 1095, 386, 268, 426, 226, 821, 24]. Newtonian [904]. Nicholas [696, 86]. Nicolai [646]. Nicolas [803]. Nicolaus [644]. Nicolson [240]. niederländische [925]. Niels [622, 1003]. Nielsen [22]. Nietzsche [476]. Nilakantha [28]. Nils [571]. Nineteenth [184, 619, 1055, 789, 786, 496]. Ninth [453]. Nissen [693, 1106, 373]. Nitre [581]. Nobel [622, 622]. Nombre [271]. nombres [110]. Nomenclature [584, 605]. Norbert [296]. Norden [212]. Norman [591, 699]. North [864, 1055, 244, 102]. notas [19]. Note [804, 665, 52, 227, 463, 110, 763, 327, 745, 342, 110]. Notebook [362]. Notes [989, 1036, 113, 885, 777, 1097, 197, 662, 833, 667, 689, 802, 53, 985, 918, 331, 498, 560, 230, 808, 146, 56, 960, 1068, 12, 585, 714, 767, 828, 746, 196, 451, 290, 392, 48, 273, 467, 1014, 634, 531, 775, 914, 1011, 203, 950]. notis [646]. nous [654]. Nova [831]. November [977]. Nuclear [255, 34, 1108]. Numbers [251, 110]. Numerological [983]. nuncius [107]. Nuove [1000]. Nurture [86]. Nussbaum [617].

O [967, 1053, 760, 996, 515, 1045, 368, 932, 297]. ob [1018]. O’Beirne [1110]. Oberndorf [650]. Obituaries [15, 58, 112, 149, 204, 272, 289, 330, 346, 394]. Objectives [179]. Obscura [993]. Observation [1102]. Observations [1032, 145]. Observatory [340, 55]. Obstetrics [207]. Odd [385]. Oddities [1095]. Odysseus [964]. Oehser [97]. Oersted [490, 805]. OEuvre [1116, 1076]. Oeuvres [1021, 155]. Officers [871]. O’Flaherty [535]. Okruzhnosti [1018]. Olan [436]. Old [172, 387, 374, 641]. Oldham [823]. O’Leary [30]. Oliver [89, 507, 71, 896]. Olivieri [207]. Olmsted [400, 400]. O’Malley [424, 595, 152, 758]. Omnia [875, 816]. One [764, 825]. Only [786]. Opening [704]. Opera [875, 816]. Opinion [242]. Opinions [868]. Opponents [971]. Opportunities [779]. Optical [1094, 200]. Opticks [426]. Opuscula [862]. Oraciones [517]. Order [983, 481]. Ore [941, 501, 1003]. Ores [351]. Oresme [556]. Organic [536]. Organization [710]. Oriental [843]. Orientation [881, 596]. Orientations [677]. orígenes [753]. Origin [1090, 604, 741, 380, 309, 436, 559]. Original [871]. Origine [529, 559, 816]. Origines [990, 411]. Origins [942, 643, 79, 752, 514]. Orleans [327]. Ornament [870]. Ornithologie [377]. Ornithology [376]. Orr [45]. Orrery [648]. Orthopedic [45]. ortu [976]. Orville [598]. Oscar [701]. Osiris [367]. Ostwald [620, 620]. Other [615, 1120, 785]. Otho [697]. Otto [255, 783, 1114, 302, 577, 143]. Our [1007]. Over-Population [786]. Ovum [742]. Owen [45]. Owsei [997]. Oxford [239]. Oxygen [684]. Øystein [501, 1003].

P [358, 894, 650, 359, 237, 1018, 670, 1047, 116, 313, 990, 647, 721, 1099]. P.-H [990]. Pachter [234]. Padua [1066]. Pagan [1078]. Pagel [1077]. Palestine [250]. Palfyn [341]. Palissy [975]. Palmer [248]. Panofsky [727]. Pansophical [425]. Papaspyres [542]. Paper [736]. Papers [598, 1051, 359]. Paracelsus [296]. Paracelsus. [234, 352, 1077]. Parallax [713]. Parfait [271]. Paris [118, 145, 428]. Pariser [145]. Park [116]. Parker [247]. Parma [423]. Parr [615]. Parsons [364]. Part [605, 884, 764, 776, 775, 914]. Parva [723]. Pasquale [70]. passé [411]. Passed [30]. Past [441]. Pasteur [576, 1040]. Pastore [86]. Pasture [587]. Patent [864]. Patients [300]. Patrick [1026]. Patterns [1070]. Paul [733, 443, 159, 173, 97, 570, 539, 895, 729, 861, 219, 410, 822, 616, 641, 299]. Paul-Henri [159]. Paulus [291]. Pavlov [1120]. Pays [1001]. Payson [672]. Paz [517]. Paz-Soldan [517]. Peace [140, 221]. Péan [990]. Pearce [539]. Peck [698]. Peel [1042]. Peiresc [803]. Pelagians [971]. Pennsylvania [649]. pensée [123, 74, 795]. Pensees [664]. People [305, 181, 256, 786]. Percy [403]. Perfection [324]. Pergamon [674]. Period [967]. periodo [928]. Perkin [1024]. Permanent [955, 326]. Perry [209, 728]. Persia [306]. Persistent [490]. Person [258]. Personalia [14, 114, 147, 202, 231, 287, 329, 345, 449, 466]. Perspective [888, 316]. Pestilence [752]. Peter [171]. Petersson [870]. Petite [269]. Petran [872]. Petroleum [1109]. Petrus [66]. Pettegrove [358]. Petty [1042]. Peut [654]. Peut-on [654]. Pfeiffer [78]. Phaenomena [8]. Pharaonique [1017]. Phelps [245]. Phenology [745]. Philadelphia [430, 663, 711]. Philip [857, 721, 1099, 536, 1035]. Philipp [315]. Philippe [852]. Philippi [1096]. Philocrates [365]. Philolaos [195]. Philon [160]. Philosoph [873]. Philosopher [294, 823, 67, 173, 503]. Philosopher-Scientist [173]. Philosophers [688, 573]. Philosophiae [474]. Philosophical [802, 317, 1038, 699, 1077]. philosophie [852, 74, 357]. Philosophiques [1021]. Philosophorum [602]. Philosophy [358, 969, 834, 317, 127, 1051, 44, 95, 699, 167, 670, 131, 16, 136, 276, 332, 453, 932, 474, 591, 1026, 261, 185, 82, 94]. Phlogiston [684]. Photographer [993]. Photography [993]. Physical [165, 968, 167, 370, 721]. Physician [294]. Physicist [398]. Physics [497, 622, 685, 370, 957, 255, 891, 783, 167, 670, 817, 898]. Physiker [873]. Physiologie [537]. Physiology [622, 259, 218]. physique [596]. Pieper [863]. Pierce [235]. Pierre [970, 60, 74, 721, 1116]. Pierre-Maxime [74]. Piggott [157]. Pilgrimage [555]. Pioneer [360]. Pioneers [671]. Pitane [205]. Pittura [1049]. Pla [216]. Place [223, 602, 745, 666, 1091]. Plague [1041, 576, 568]. Plancius [66]. Planetarium [648]. Planetary [54, 285, 1064, 511, 974]. Planetis [754]. Planets [285, 967]. planifier [654]. Planimetria [804]. Plant [131, 87, 179, 182, 218]. Plants [178, 309, 406, 437]. Plato [44, 291]. platonicienne [74]. Platonicum [291]. Platonis [291]. Platonism [857]. Playfair [874]. Pline [293]. Plinius [151]. Plotinus [493]. Plough [587]. pluralité [796]. Pneumatic [632]. Pneumatics [160]. Poet [581, 210, 398]. Poffenberger [121]. Poincaré [700]. Points [4]. Polhem [298]. Policies [886]. Political [184]. Politics [724]. Pomp [752]. Pomponazzi [495]. Pontus [153]. Pope [1115]. Popper [335]. Popper-Lynkeus [335]. Population [786, 75]. por [486]. Porta [1020]. Porter [1052]. Portuguesa [639]. Possibility [765]. Post [882]. Potato [180]. Poteat [150]. Powell [704, 1002]. Poynter [300, 20]. Practica [755]. Practice [486, 565]. Practitioners [867]. Prafulla [972]. Prandtl [1023]. Pre [883]. Pre-Copernican [883]. Precious [151]. précurseur [462]. Precursor [7]. Precursors [192]. Prediction [270]. preeuclidiennes [159]. Preface [2, 48, 140, 221, 254]. Préfets. [1001]. Prehistórica [125]. Prehistory [342]. Premier [271]. Premiers [596]. préparation [60]. presalernitano [928]. Presidential [385]. prestigieux [411]. Presumed [1093]. Prevalence [786]. Price [104, 1020, 754]. Priest [463]. Priestley [759]. Primer [317]. Primitive [917, 539, 695]. Prince [1001]. Princes [613]. Princeton [648]. Principal [1081, 793]. Principally [472]. Principate [724]. Principia [226, 474]. principiis [816]. Principle [342, 1094]. Principles [783]. Printed [831]. Printer [757]. Printing [382]. Pritchard [312, 172]. Prize [622]. Probierbüchlein [154]. Problem [821].

Probleme [369]. Procedures [858]. Proceedings [1024]. Process [830]. processivi [816]. Prochaska [872, 872]. Procopius [390]. Products [782]. proemio [646]. Profession [1069, 1037, 1098]. Progress [233, 328, 396, 562, 636, 716, 848, 921, 255, 987]. Propagandist [354]. Prophet [508]. Proportionibus [817]. Proposed [459, 458, 710]. Propositional [586]. Propulsion [342]. Prosopographie [1017]. Protection [430]. Protective [391]. Protozoology [143]. Province [728]. Pseudonyma [381]. Psotnícková [737]. Psychiatrie [1073]. Psychiatry [1074]. Psychischen [39]. Psychoanalysis [650, 650, 750]. Psychological [49]. psychologie [443, 1076]. Psychology [442, 787, 1120, 183, 260, 589, 1074]. Psychopathology [661]. Ptolemaeus [106]. Ptolemaic [572]. Ptolémée [930]. Ptolemy [1009, 106, 468]. Public [1101, 912]. Publication [53]. Publications [1058, 1123, 847]. publiée [990]. Published [1046, 303, 179]. Pulkovo [340]. Pulmonary [683]. Purkyne [760, 875, 737]. Purpose [144]. Pusalker [305]. Putnam [294]. Pythagore [159]. Pythagoreans [195].

Q [210]. Qabus [613]. Qanunu’ [899]. quadrati [1032]. quae [291, 976]. Quaestiones [423, 816]. Quaker [663]. quas [646]. que [654]. Quecksilberbergwerk [935]. Quelques [1032, 850]. Queries [57, 148, 347, 948, 201, 274, 13, 988, 111, 288, 393, 452, 561, 633, 715, 606]. Query [268]. Quest [94]. Questions [687]. Quinn [864]. Qurra [662].

R [1070, 678, 890, 19, 814, 356, 573, 1109, 251, 317, 1042, 127, 359, 1025, 1111, 867, 33, 472, 752, 514, 305, 964, 64, 1110, 43, 239, 28, 673, 990, 647, 861, 174, 362, 927, 897, 468, 870, 402, 503, 871, 436]. Rabbi [545]. Rabelais [110]. Rabelaisiennes [933]. Races [580]. Radcliffe [239]. Radhakamal [92]. Radin [539]. Radioactivity [911]. Radius [323]. Raeder [21, 646]. Rafinesque [179]. Rafinesquianus [179]. Rahman [859]. Rain [199]. Rainbow [323, 773, 944, 982]. Raleigh [864]. Ralph [722]. Ramon [756]. Ramos [125]. Randall [408]. rapports [123, 60]. Rationalism [156]. Ratisbonensis [755]. Raven [567, 823]. Ray [356, 766, 802, 972]. Rayhan [899]. Raymond [1053, 357, 768]. Raymundus [291]. Re [154, 1046]. Reader [1035]. Readers [1007]. Readings [260]. Real [802]. Reality [167]. Reason [617, 409]. Réaumur [906, 906]. Received [1058, 1123]. Rechenbuch [755]. recherche [654]. recherches [60]. Reconstructed [612]. Reconstruction [251]. Record [343]. Recorde [224]. Redcliff [180]. Reductions [179]. Reference [472, 303]. References [957]. Reflection [200]. Reflections [1026, 700]. Reform [223]. Refraction [773, 200]. regarding [227, 490]. Regiomontanusp [197]. Registration [1008]. règne [1001]. Regula [1096]. Regulae [699]. Reichenbach [261]. Reinhold [931]. Rejection [765]. Related [1051, 1072]. Relating [116, 172]. Relation [655]. Relations [242, 55, 1115, 472, 32]. Relazioni [70]. Relief [1033]. Religion [228, 570, 243, 567]. Remak [702]. Remarks [266, 661, 337]. Reminiscences [5]. Renaissance [1019, 1042, 865, 694, 983, 866, 295, 1078, 1077, 434]. Rene [795, 906, 850, 473, 576, 307, 990, 254, 990]. Report [108]. Representation [676]. Reproduced [871]. Reproductions [497]. Republic [320]. Research [942, 361, 710, 133]. Researcher [361]. Researches [165]. Resolve [684]. Resources [496]. Respect [765]. Results [363]. Retrospective [829]. Reuben [613]. Reversible [1033]. Review [815, 967, 978, 1000, 369, 485, 705, 320, 1041, 1118, 35, 177, 187, 749, 1101, 299, 590, 733, 443, 649, 1049, 1070, 541, 1119, 160, 502, 567, 678, 973, 1020, 214, 637, 81, 159, 499, 607, 638, 751, 853, 890, 961, 962, 1071, 1081, 596, 781, 358, 1053, 153, 156, 240, 574, 864, 933, 121, 183, 260, 442, 589, 737, 760, 787, 872, 875, 1080, 903, 19, 238, 294, 430, 447, 469, 513, 643, 755, 814, 900, 929, 969, 970, 971, 42, 478, 586, 1116, 26, 27, 31, 79, 154, 173, 184, 254, 255, 356, 367, 397, 399, 403]. Review [474, 479, 516, 573, 595, 622, 793, 794, 851, 854, 991, 993, 995, 1079, 408, 896, 376, 377, 693, 894, 591, 857, 998, 1075, 63, 996, 89, 642, 432, 1109, 568, 1002, 1085, 34, 730, 162, 311, 874, 120, 122, 165, 308, 361, 476, 515, 578, 620, 1024, 972, 409, 477, 507, 548, 619, 697, 719, 747, 242, 780, 38, 166, 704, 306, 597, 722, 353, 650, 1073, 30, 78, 246, 247, 250, 251, 292, 317, 365, 598, 259, 1045, 1019, 720, 504, 796, 906, 968, 1021, 1042, 732, 735, 97, 677, 41, 127, 150, 186]. Review [236, 304, 441, 444, 445, 448, 510, 674, 779, 1100, 893, 73, 65, 429, 731, 850, 904, 1051, 210, 211, 249, 696, 679, 823, 21, 22, 68, 241, 297, 483, 615, 866, 1046, 1121, 778, 869, 891, 926, 1022, 1023, 1057, 80, 359, 427, 886, 1026, 888, 1120, 783, 873, 1038, 1039, 71, 44, 425, 378, 379, 404, 439, 440, 534, 621, 691, 736, 782, 924, 963, 1025, 1055, 170, 95, 128, 134, 237, 1052, 1111, 446, 819, 676, 1043, 117, 501, 570, 572, 754, 865, 867, 1115, 623, 1018, 546, 539, 33, 24, 69, 472, 473]. Review [532, 699, 821, 295, 319, 354, 400, 517, 734, 750, 753, 907, 169, 934, 135, 576, 752, 895, 82, 94, 163, 167, 168, 315, 316, 370, 371, 426, 670, 692, 748, 852, 931, 1114, 23, 508, 514, 423, 928, 675, 817, 209, 728, 123, 407, 83, 360, 1044, 37, 130, 152, 158, 188, 189, 305, 312, 318, 366, 538, 695, 703, 816, 898, 486, 964, 975, 994, 1076, 90, 644, 1040, 125, 64, 351, 646, 792, 797, 72, 1110, 213, 133, 20, 39, 40, 43, 93, 234, 296, 300, 350, 352, 401, 471, 506, 537, 540, 756, 791]. Review [923, 997, 1048, 1078, 1112, 96, 512, 613, 966, 1105, 901, 925, 88, 500, 543, 131, 124, 1084, 639, 487, 1047, 930, 727, 935, 999, 1001, 1050, 239, 119, 17, 18, 25, 28, 32, 59, 60, 61, 62, 66, 70, 74, 76, 91, 115, 116, 151, 155, 157, 161, 172, 205, 206, 207, 208, 215, 248, 252, 291, 293, 298, 301, 302, 307, 313, 348, 349, 357, 363, 364, 424, 431, 433, 480, 505, 544, 545, 565, 566, 608, 609, 640, 645, 673, 690, 725, 729, 789, 45, 235, 262, 470, 701, 702, 810, 813, 858]. Review [862, 965, 1017, 1077, 84, 85, 990, 1074, 855, 892, 902, 355, 547, 647, 575, 824, 671, 698, 721, 759, 905, 1099, 132, 757, 758, 1003, 1083, 700, 549, 569, 594, 861, 863, 976, 977, 174, 362, 428, 820, 927, 1102, 612, 1106, 897, 468, 511, 592, 593, 611, 724, 992, 1107, 564, 67, 870, 1117, 171, 118, 164, 245, 368, 398, 1054, 1103, 726, 402, 610, 723, 788, 694, 542, 795, 1016, 92, 212, 219, 410, 411, 577, 822, 75, 859, 503, 616, 77, 899, 533, 936, 1122, 1108, 29, 614, 860, 932, 261]. Review [314, 672, 811, 1072, 812, 617, 887, 216, 818, 303, 373, 380, 381, 382, 535, 790, 1113, 648, 871, 974, 36, 86, 87, 126, 129, 175, 176, 178, 179, 180, 181, 182, 217, 218, 243, 244, 253, 256, 257, 258, 309, 310, 372, 374, 375, 405, 406, 434, 435, 436, 437, 438, 475, 481, 482, 484, 509, 536, 587, 588, 618, 641, 784, 785, 786, 809, 825, 889, 1015, 1056, 1082, 1104, 185, 868, 1067, 571, 856]. Reviewing [48]. Revision [458]. Revisited [1053]. Revolution [678, 409, 1042, 404, 491, 974, 962, 227, 784, 647]. Revolutionary [428, 1054]. Revolutions [644]. Reymond [123]. rib [1032]. Ricci [614, 207]. Rice [648]. Richard [590, 896, 122, 697, 247, 569, 435]. Richards [897]. Richardus [291]. Richter [790, 790]. Ridlon [45]. Rights [188]. Rise [370, 898, 364, 701, 582, 261, 617]. Rittenhouse [648, 1033]. Ritter [703]. River [1002]. Roanoke [864]. Roback [589]. Robert [442, 255, 991, 78, 292, 796, 1051, 379, 797, 133, 93, 997, 45, 468, 611, 811, 372, 335, 463, 225, 386, 282, 471, 643, 120, 902, 757]. Rochd [18]. Rochot [1116]. Rockwood [1053]. Rocque [975]. Rod [391]. Roebuck [363]. Rogers [87, 244]. Roland [595, 594]. Rolf [977]. Roller [163, 692]. Rollo [1074]. Roman [611, 724, 788, 151]. Romance [361]. Romanell [1026]. Romanic [303]. Romano [819, 70, 819]. Romanoff [36, 36]. Romantic [257]. Rome [506, 724]. Romeinschen [38]. Ronald [752, 547]. Room [786]. Rose [588]. Rosen [69]. Rosenfelda [1018]. Rosenthal [433]. Rosinger [304]. Ross [650]. Roswell [27]. Rothschuh [537]. Rouse [1016]. Rovighi [423]. Roy [539, 94, 588]. Royal [24, 1110, 96, 757, 820, 1117, 281, 422]. Rudolf [1057, 316, 212, 549]. Rudolph [869]. Rule [703]. Rülein [863]. Rumford [1034, 574]. Runge [72, 72]. Runstaven [571]. Rupescissa [604]. Rush [26, 26, 427]. Russell [1082, 895, 75]. Russia [84, 398]. Russian [494]. Russo [566]. Ruth [623, 892].

S [442, 238, 1075, 719, 38, 1045, 1019, 968, 731, 440, 446, 1018, 252, 545, 1102, 592, 564, 542, 974, 179, 1082]. Sa [60, 1116]. Sacheverell [693]. Sacks [255]. Sacred [187]. Sadan [524]. Sailing [1013]. Saint [965, 59]. Saint-Eloi [965]. Salaman [180]. Salamanders [785]. Sallander [851]. Salpetersäure [578]. Salt [959, 284]. Salts [267]. Salud [517]. Salusbury [913, 572, 1092]. Sambursky [968]. Samuel [255, 73, 750, 670, 53]. Sanches [297]. Sansom [215]. Santillana [572, 726]. särskild [571]. Sarton [712, 940, 1029, 961, 447, 740, 844, 367, 839, 840, 642, 432, 842, 674, 866, 865, 846, 841, 837, 847, 845, 843]. Satan [444]. Saul [250]. Saunders [152, 424]. Savages [539]. savant [729]. Savasorda [344]. Savioz [357, 357]. Sax [786]. Saxl [43, 593]. Sceptical [557]. Scherz [646]. Schib [1057]. Schicksal [1056]. Schilpp [173]. Schmalhausen [129]. Schneider [375]. Schofield [1051]. Scholar [354, 401, 501]. School [1119, 5, 363, 1010, 1091]. Schriften [976, 352, 1048]. Schrödinger [610]. Schuhl [74]. Schuyler [991]. Schweitzer [95]. Science [232, 396, 563, 562, 635, 668, 636, 716, 849, 848, 876, 921, 718, 385, 121, 238, 286, 513, 643, 844, 1089, 403, 658, 832, 834, 995, 951, 987, 1086, 450, 802, 642, 432, 949, 920, 409, 223, 222, 30, 97, 779, 912, 1100, 904, 866, 80, 663, 1025, 865, 628, 669, 472, 395, 496, 675, 1012, 10, 717, 630, 955, 336, 16, 136, 140, 276, 332, 480, 453, 688, 84, 990, 492, 881, 1083, 841, 1030, 837, 457, 666, 922, 919, 523, 491, 617, 624, 772, 275, 126, 769, 856, 780, 221, 961, 903, 447, 79, 719]. Science [127, 1121, 888, 1039, 378, 128, 623, 576, 82, 163, 234, 67, 87, 243, 482, 187, 749, 567, 156, 479, 619, 747, 677, 886, 570, 748, 565, 690, 860, 887, 253, 481]. Sciences [998, 996, 302, 990, 850, 566, 494, 564, 1072]. Scientiarum [931]. Scientific [1101, 678, 1085, 242, 504, 763, 21, 80, 1038, 473, 168, 709, 565, 235, 1062, 564, 860, 261, 1101]. Scientifique [307, 850, 654, 1001]. scientifiques [505]. Scientist [101, 463, 337, 711, 503, 173, 359, 212]. Scientists [514, 1079, 96]. Scienze [1000]. scienziati [70]. Scipio [348]. scolastiques [795]. Scot [469]. Scott [473]. Scottish [1084]. scriptis [646]. Sculptures [790]. Sea [1107]. Seafarers [1107]. Sébastien [594]. secoli [928]. secolo [161]. Second [834, 704, 284, 48, 221]. Second-Century [284]. Sections [922]. Secundus [151]. Seele [39]. Segal [1018]. sein [72, 863, 577]. seine [976]. seinem [977]. seiner [1056]. seines [119]. Select [53]. selecta [862]. Selected [1101, 20, 296, 175]. Selection [1124, 1082]. Seleucid [967]. Sellars [94]. Selskab [21]. Sendino [61]. Senescence [893]. Sense [253]. Sentiment [435]. September [876, 1123, 832]. Servet [594]. Servetus [595, 595]. Service [1101]. ses [60]. Seventeenth [995, 472, 728, 300, 647, 282, 582]. Seventeenth-Century [647]. Seventh [276]. Seventy [534, 16, 136, 276, 332, 453]. Seventy-Eighth [332]. Seventy-Fifth [16]. Seventy-Five [534]. Seventy-Ninth [453]. Seventy-Seventh [276]. Seventy-Sixth [136]. Sexagenarium [198]. Shātir [1064, 883]. Shaar [545]. Shackleton [796]. Shamayim [545]. Shape [408]. Shaw [31]. Shell [1110]. Shelton [896]. Shepard [588]. Sherrington [65]. Sherwood [963, 83]. Ship [35]. Shlomon [545]. Short [810, 33]. Shryock [697, 569]. Sicile [930]. Sidereus [107]. Sidgwick [732]. Sidney [896, 365]. siècle [60, 795, 804, 60]. siècles [505]. Sifa’ [1076]. Sigerist [947, 262]. siglo [819]. Sigmund [575, 824]. Significance [817, 1062]. Sigrid [792]. Silva [125]. Silvanus [1047]. Simon [1081, 793, 990, 1016, 818]. Simpson [130, 310]. Sina [1076]. since [780]. Singer [1042, 673, 927, 236, 189, 32, 470, 858, 176, 565]. Sir [116, 45, 573, 281, 744, 238, 573, 237, 870]. Sisco [154, 351, 906]. sit [1032]. Sitwell [693]. Six [866]. Sixteenth [235, 614, 686]. Sixteenth-Century [686]. Sixth [136]. Sixty [310]. Sizzi [945]. Skelton [871]. Sketch [320]. Sketches [691]. Sleepwalkers [1067]. Sloane [573, 1104]. Slouka [533]. Small [269]. Smith [154, 699, 320, 1071, 445, 351, 1054]. Smithsonian [97]. sobre [76, 208]. Social [749, 719, 242, 888, 404, 133, 1084, 92, 180, 481, 753, 1001, 312, 133]. Societies [11, 313]. Society [534, 881, 718, 385, 24, 576, 422, 837, 922, 624, 96, 1117, 126]. Sociological [41]. Soda [959]. Sofia [423]. Solar [883]. Soldan [517]. Soldier [401]. Solidification [326]. Solomon [657]. Solvay [596]. Some [255, 879, 193, 340, 603, 829, 691, 767, 496, 1063, 957, 775, 845, 914, 1011, 419, 1032, 649, 80]. Sommer [990]. Son [1116, 1076]. Songs [366]. Sons [97]. Sooner [1046]. Soranu [997]. Source [440]. Sourcebook [255]. Sources [526]. sous [1001, 990]. South [419]. Soviet [628, 623, 405]. Sozialethische [352]. Sozialpolitische [352]. Space [670, 1038]. Space-Time [670]. Spätscholastik [63]. Special [420, 303]. Spectrum [199]. Speculation [805]. Speculative [904]. Speiser [751, 853]. Spemann [577]. Spenser [894]. Sphaera [283]. sphère [205]. Spheres [644]. Spiel [28]. Spirit [1095]. Sport [28]. Sprat [336, 1117]. Sprott [749]. square [1032]. St [755]. Staatsbibliothek [755]. Stahl [348]. Standard [422]. Standen [187]. Standing [786]. Standpoint [484]. Stanislaus [1050]. Stanley [154, 351]. Stark [696]. Starr [508, 309]. State [304]. Statement [200]. States [684, 807]. Statesman [549]. Staudt [787]. Stearn [373]. Stebbins [178]. Steel [906]. Stegner [704]. Steinitz [1049]. Sten [298, 480, 789]. Stenonis [646]. Stephen [525, 986, 526]. Stern [647]. Stevenson [622]. Stevin [1081, 1080, 793]. Stiernotte [670]. Stiftsbibliothek [755]. Stillman [572, 868]. Stimson [96]. Stoic [957, 478]. Stoïciens [123]. Stone [1108]. Stones [151, 897]. Stora [298, 789]. Storck [438]. Storia [170, 161]. Stories [320]. Störig [778]. Story [35, 184, 34, 735, 97, 88, 810, 892, 759, 87, 244, 310]. Stow [258]. Stresemann [377]. Stresser [990]. Stromgren [21, 21]. Structural [282]. Structure [408, 166, 133, 721, 314]. Struik [1081, 1080]. Stuart [867, 157, 544]. Student [361]. Students [5]. Studien [63]. Studies [409, 250, 699, 363, 303, 516, 1109, 127, 811]. Study [1101, 335, 160, 961, 971, 242, 917, 504, 1055, 676, 539, 472, 699, 319, 514, 846, 695, 194, 465, 215, 235, 1083, 382, 435, 484, 41, 779]. Stukeley [157]. Subject [401]. Subjects [194]. subterraneorum [976]. Succession [882]. südamerikanischer [419]. Sufi [859]. Suggested [1090]. Suicide [366]. suivie [205]. Sun [967, 54, 812]. Sunspots [199]. super [816]. Superintendents [374]. Surgeon [235]. Surgery [791, 45]. Survey [1075, 787, 720]. Survivals [382]. Suwaru’l [859]. Suwaru’l-Kawakib [859]. Svendsen [903]. Swartz [1099]. Swedish [480]. Sydney [151]. Syllabus [479]. Sylvain [428, 1054]. Symbolism [380]. Symcotts [300]. synopsis [291]. Synthesis [315]. System [459, 474, 421, 686]. Systematis [872]. Systeme [970, 970]. Systems [572, 511, 572].

T [637, 499, 121, 255, 399, 991, 165, 697, 359, 572, 407, 698, 870, 1117, 373]. Tūsī [200]. Tabaq [54, 285]. tablas [19]. Tables [774, 1075, 105]. Taddeo [423]. Tagliacozzi [235]. Tagung [119]. Tahafut [818]. Tales [380]. Taliaferro [468]. Tamburlaine [615]. Tar [942]. Taton [307, 990, 307, 850]. Taylor [381, 963, 867, 83, 964]. Teach [235]. Teacher [361]. Teaching [286, 450, 949, 669, 395, 10, 717, 919, 275]. Teague [438]. Technical [1110].

Technik [515, 540]. techniques [566]. Technologie [923]. Technologists [362]. Technology [403, 1042, 1046, 767, 946, 959, 32, 673, 174, 927, 896, 404, 811]. Tedesche [365]. Teeth [981]. Telescope [854, 69]. Telescopes [1066]. Tell [894]. Temkin [997]. Temperament [720]. Tercentenary [24]. terra [976, 705]. Tertiary [944]. Testament [78, 172]. Texas [320]. text [656, 506]. texte [205]. Texts [967, 172]. Th [218]. Thabit [662]. Thal [398]. Thales [511, 852]. Their [55, 366, 901, 373, 382]. Theodoric [791]. Théodoridès [990]. Theodosius [129, 889]. théologique [60]. Theology [463, 242]. Theoloog [66]. Theophrast [1048]. Theophrastus [530, 897]. Theoretical [101, 539]. théoricien [730]. théorie [462]. Theories [781, 371, 670, 684, 370, 1039]. Theory [982, 874, 713, 1064, 943, 37, 741, 133, 756, 9, 883, 911, 721, 436]. Thereof [904]. thermalisme [411]. Thermometer [422]. Thiele [251]. Things [436]. Thinker [210]. Third [951]. tholdoth [446]. Thomas [238, 894, 507, 359, 237, 213, 196, 1084, 45, 1117, 974, 744, 1092, 807, 823, 359, 817, 618]. Thompson [574, 1047, 2, 252, 932]. Thomson [89, 927]. Thoreau [745]. Thorndike [940, 214, 995]. Those [1045]. Thought [733, 184, 1085, 242, 236, 168, 709, 565, 1062, 564, 258, 184, 532, 611, 860, 974]. Ti [249]. tijd [38]. till [571]. Tilton [507, 71]. Timaei [291]. Time [1038, 670, 199, 185, 825, 1108]. Times [705, 78, 673, 235, 533, 1015, 1107]. ting [55]. Titan [1083]. Title [107]. Titled [511]. Tobacco [509]. Tobias [815, 700]. Today [533]. Todestag [977]. together [320]. Tome [990]. Tomus [875, 816]. Tools [80]. tot [38]. tra [70]. trabalhos [486]. Tractat [1044]. Tractatus [817]. Trade [32]. Tradicões [125]. Tradition [431]. Traditional [9]. traditionnelle [132]. Traditions [1030]. Traité [1032, 794, 52]. traités [205]. Traktat [1018]. Transformation [911]. Translation [469, 998, 8, 797]. Translations [339, 420]. Translator [53]. Transmission [250]. Trattato [1049]. Trauman [1049]. Travels [502]. Treatise [1018, 351, 1032]. Treatises [214]. Tree [62]. Trembley [503]. Tresbéns [1044]. Trevor [1042, 736, 927]. triad [804]. Triade [804]. Tribal [366]. Trinity [969]. Triumph [590, 982, 512, 617]. True [401, 666]. Truth [586]. Tswett [736]. Tuberculosis [576]. Tudor [867]. Tullius [150]. Tuning [720]. Turba [602]. Turnbull [425]. Twenty [650]. Two [225, 572, 776, 425]. Tyard [153]. Tycho [21, 22]. Type [743]. Types [382].

U.S.S.R. [1122]. Überblick [694]. ‘udi [899]. Uexküll [407]. Ugo [294]. Ulich [379]. Ulrich [863]. Unbelief [780]. Undersogelse [22]. Underwood [565]. UNESCO [188]. Uniformly [765]. Union [834]. Unique [569]. United [684, 807]. Universalismus [1114]. Universe [153, 236, 1039, 700, 973]. University [1119, 385, 239, 569, 840, 80, 514]. Unpublished [525, 879, 225, 386, 464]. unseres [407]. until [789, 724, 1015]. Updike [382]. upon [775, 914, 1011]. Ursprung [39]. Urzeit [694]. Use [339, 993]. Uses [441, 382]. Usher [672]. Utrum [1032].

V [622, 1018, 94, 207, 729, 976, 1072, 407]. Valentin [702]. Vallicrosa [819, 19, 76, 208]. Value [335]. Vandeul [429]. Vanni [423]. Varagnac [132]. Variation [309, 178]. Variations [804, 226]. Various [612]. Vartanian [504]. VAT [656, 656]. Vater [620]. Vatican [775, 914]. Vavilov [309]. vecchia [23]. Vedic [305]. Veith [249]. Velikovsky [90]. Velocity [765]. Venesection [327]. Venice [1045, 1066, 1011]. Venous [522]. Vercoutter [990]. Vererbungsgesetze [1056]. Verrazano [102]. Verrier [366]. Version [557, 930]. Versions [420]. Versionum [544]. versus [802]. Vesalius [152, 555, 554, 64, 470]. VI [875, 1066, 544]. Viêtnam [162]. Victor [117, 812]. Victorians [735]. Videnskabernes [21]. Vie [1116, 132]. Vienna [500]. Vienna-Klosterneuburg [500]. Viëtor [210, 210]. Views [461, 628, 871, 185]. VII [995, 1009]. VIII [995, 207]. Vinci [1049, 792, 350, 424]. Virchow [549]. Virginia [787]. viribus [797]. Virieux [123]. Virieux-Reymond [123]. Virolleaud [990]. Vistas [992]. vistos [486]. VIth [207]. Vladimirovich [175]. Vladislav [872, 875]. Vogel [755]. Vogelbücher [693]. Vol [1119, 962, 442, 367, 793, 359, 40, 540, 1050, 575, 617]. Vol.III [962]. Vol.VI [970]. Vol.VII [970]. Volcher [862, 471]. Volker [419, 694]. Volume [137, 278, 413, 518, 552, 625, 706, 799, 908, 1004, 1126, 1081, 781, 969, 1042, 677, 782, 305, 1105, 2, 48, 140, 221, 349, 363, 673, 262, 824, 927, 860, 871]. volumen [544, 291]. Volumes [995]. vom [822]. Vorläufer [63]. Vorträge [1022, 119]. Voyage [102, 871]. Voyages [864, 871].

W [749, 26, 255, 356, 896, 165, 166, 598, 445, 696, 300, 923, 487, 62, 229, 647, 1099, 398, 243, 434]. Wên. [249]. Waele [363]. Waerden [690]. während [118]. Waited [1046]. Waldo [1083]. Waley [236]. Walker [810]. Wallace [704, 1082]. Waller [851]. Walleriana [851]. Wallis [644]. Walter [864, 752, 1077, 438]. Walzer [291, 291]. Wang [677]. War [472, 955]. Ward [27]. Warshaw [135]. Was [391, 831, 51]. Wasser [1085]. Water [256]. Watson [703, 378]. Watt [280]. wattas‘idat [17]. Wayne [260]. Weatherwax [641]. Webster [398, 41, 235]. Wedgwood [710]. Wedkiewicz [1050]. Weeks [534, 924]. Weevers [218]. Weights [686, 675]. Weil [736]. Weissberg [448]. Welch [701]. Welfare [1084]. Weltbild [210]. Weltbildes [855]. Weltgeschichte [778]. Wendell [507, 71]. Wendler [977]. Wentworth [591, 2]. wereldbeeld [164]. Werk [869, 792, 72, 577]. Werke [241]. Werkmeister [166]. Werksarzt [935]. Werner [442]. Wesley [1002, 704, 492]. Wesleyan [514]. West [917, 704, 85]. Western [607, 184, 479, 532, 215, 974]. Weyl [82]. Wheeler [245]. Whether [1032]. White [576, 1024]. Whitfield [779]. Whitmore [698, 1117]. Whitney [477]. Whittaker [781, 371]. Who [320]. Wiener [721]. Wilbur [598]. Wild [435]. Wilfrid [373]. Wilhelm [120, 1023, 1114, 119, 863, 620]. Wilkins [802]. Wilkinson [1010, 698]. Willard [245]. William [255, 311, 735, 237, 907, 703, 20, 1047, 348, 373, 535, 243, 683, 144, 1118, 732, 898, 157, 701, 698]. Williams [1042, 782, 1025, 927, 736, 1025, 303]. Willstätter [122]. Willy [77, 257, 785]. Wilmer [506]. Wilsdorf [976]. Wilson [255, 747, 246, 898]. Wind [1078]. Wings [866]. Winners [622]. Winnett [45]. Winthrop [31]. Wirklichkeitsbildes [407]. Wissenschaft [210, 778]. wissenschaftlichen [119]. wissenschaftliches [72]. Withering [698]. Withington [896]. Wittkower [43]. Wolfson [115, 969]. Wonders [785]. Wood [780, 823, 94]. Word [913, 945]. Words [380]. Work [921, 987, 904, 21, 5, 473, 767, 764, 776, 1063, 575, 824, 255, 213, 1083]. Works [499, 1081, 662, 339, 399, 793, 731, 152, 116, 425, 901, 175, 53]. World [973, 890, 993, 477, 968, 1121, 572, 1115, 539, 151, 215, 364, 310, 292]. Worlds [236, 90]. Wretched [194]. Wright [506, 598]. Writing [512, 252]. Writings [20, 699, 65, 296, 997]. Written [565]. Wunder [124]. Wyman [193, 461].

X [928, 766]. X-Ray [766]. XI [928]. XII [816]. XIIe [60, 804]. XIV [156, 363]. XIVth [105]. XIX [161]. XV [40]. XVII [819]. XVIIe [505, 795]. XVIIIe [505].

Ya [934]. Yates [756]. Years [485, 255, 650, 1045, 575, 824, 343, 616, 932, 1093, 218, 310, 534]. Yellow [320, 249, 320]. Yerkes [442]. York [45]. Young [823, 4]. Yushkevich [1018].

Zdenek [875]. Zeichen [124]. Zeit [976]. Zeitlin [365]. Zellularpathologie [410, 822]. Ziel [39]. Zimmer [28]. Zinner [369, 925]. Zirkle [84]. Zivot [872]. Zorzi [1045]. zum [962]. zur [962, 377, 63, 515, 540, 923, 976, 694, 219, 410]. Zweig [124].

References

Anonymous:1950:FMa

[1] Anonymous. Front matter. Isis, 41(1):1–2, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

      /www.jstor.org/stable/226378.

Sarton:1950:PVD

[2] George Sarton. Preface to volume 41: D’Arcy Wentworth Thompson, 1860–1948. Isis, 41(1):3–8, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

      stable/226379.

Eaton:1950:CBM

[3] Leonard K. Eaton. Charles Bulfinch and the Massachusetts General Hospital. Isis, 41(1):8–11, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

      226380.

Haraszti:1950:YJA

[4] Zoltán Haraszti. Young John Adams on Franklin’s iron points. Isis, 41 (1):11–14, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/226381.

Krogh:1950:RWC

[5] August Krogh, Eugene M. Landis, and Cecil K. Drinker. Reminiscences of work on capillary circulation. a lecture to the students in the Harvard medical school, 1946. Isis, 41(1):14–20, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/226382.

Farber:1950:CDM

[6] Eduard Farber. Chemical discoveries by means of analogies. Isis, 41(1): 20–26, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226383.

Dufrenoy:1950:PMA

[7] M. L. Dufrenoy. A precursor of a modern anthropology: François Bernier (1620–1688). Isis, 41(1):27–29, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226384.

Honigmann:1950:ATA

[8] Ernest Honigmann. The Arabic translation of Aratus’ Phaenomena. Isis, 41(1):30–31, March 1950. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226385.

Patterson:1950:HGT

[9] Louise Diehl Patterson. Hooke’s gravitation theory and its influence on newton. ii: The insufficiency of the traditional estimate. Isis, 41(1):32– 45, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226386.

Meyer-Abich:1950:THS

[10] Adolf Meyer-Abich. Teaching of the history of science. Isis, 41(1):45– 46, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226387.

Anonymous:1950:MS

[11] Anonymous. Meetings & societies. Isis, 41(1):46–47, March 1950. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226388.

Honigmann:1950:NC

[12] Ernest Honigmann, O. Neugebauer, H. B. Van Hoesen, Lynn Thorndike, Wyndham D. Miles, and George Sarton. Notes and correspondence. Isis, 41(1):47–52, March 1950. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226389.

Keimer:1950:QA

[13] Ludwig Keimer, J. Pelseneer, Lynn Thorndike, and G. Sarton. Queries and answers. Isis, 41(1):52–55, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226390.

Anonymous:1950:Pa

[14] Anonymous. Personalia. Isis, 41(1):55, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226391.

Anonymous:1950:Oa

[15] Anonymous. Obituaries. Isis, 41(1):55–57, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/226392.

Sarton:1950:SFC

[16] George Sarton and Frances Siegel. Seventy-fifth critical bibliography of the history and philosophy of science and of the history of civilization (to July 1949). Isis, 41(1):58–103, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226393.

Sarton:1950:BRBa

[17] George Sarton. Book review: Kitab Kimiya’ al- ‘it.r wattas.‘idat by Karl Garbers. Isis, 41(1):103–104, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226394.

Sarton:1950:BRBb

[18] George Sarton. Book review: Ibn Rochd (Averroès) by Léon Gauthier. Isis, 41(1):104–106, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226395.

Carmody:1950:BRB

[19] Francis J. Carmody. Book review: El Libro de los fundamentos de las tablas astronómicas, de R. Abraham ibn ’Ezra, edición crítica con introducción y notas by José M. Millás Vallicrosa. Isis, 41(1):107–108, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226396.

Pagel:1950:BRBa

[20] W. Pagel. Book review: Selected Writings of William Clowes (1544– 1604) by F. N. L. Poynter. Isis, 41(1):109, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/226397.

Hellman:1950:BRBa

[21] C. Doris Hellman. Book review: [Det Kongelige Danske Videnskabernes Selskab]: Tycho Brahe’s Description of His Instruments and Scientific Work as Given in “Astronomiae Instauratae Mechanica” by Hans Raeder; Elis Stromgren; Bengt Stromgren. Isis, 41(1):109–112, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226398.

Hellman:1950:BRBb

[22] C. Doris Hellman. Book review: Tycho Brahes Bogtrykkeri. En bibliografisk-boghistorisk Undersogelse by Lauritz Nielsen. Isis, 41(1): 112–113, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/226399.

Loria:1950:BRB

[23] Gino Loria. Book review: Il carteggio Linceo della vecchia Accademia di Federico Cesi (1603–1630) by Giuseppe Gabrieli. Isis, 41(1):113– 114, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226400.

Koyre:1950:BRBa

[24] A. Koyré. Book review: Newton Tercentenary Celebrations 15–19 July 1946 by The Royal Society. Isis, 41(1):114–116, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226401.

Sarton:1950:BRBc

[25] George Sarton. Book review: The Collected Letters of Antoni van Leeuwenhoek by Antoni Van Leeuwenhoek. Isis, 41(1):116–117, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226402.

Cohen:1950:BRBa

[26] I. Bernard Cohen. Book review: The Autobiography of Benjamin Rush by Corner George W.; Benjamin Rush. Isis, 41(1):117–118, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226403.

Cohen:1950:BRBb

[27] I. Bernard Cohen. Book review: Henry A. Ward [1834–1906] Museum Builder to America by Roswell Ward; Blake McKelvey. Isis, 41(1):118– 119, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226404.

Sarton:1950:BRBd

[28] George Sarton. Book review: Hindu Medicine by Henry R. Zimmer; Ludwig Edelstein; Spiel um den Elefanten. Ein Buch von indischer Natur by Henry R. Zimmer; The Elephant-Lore of the Hindus. The Elephant-Sport (Matangalila) of Nilakantha by Franklin Edgerton. Isis, 41(1):120– 123, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226405.

Ware:1950:BRB

[29] James R. Ware. Book review: Confucius, The Man and the Myth by H[errlee] G[lessner] Creel. Isis, 41(1):123–125, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226406.

Gandz:1950:BRBa

[30] Solomon Gandz. Book review: How Greek Science Passed to the Arabs by De Lacy O’Leary. Isis, 41(1):125–127, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226407.

Cohen:1950:BRBc

[31] I. Bernard Cohen. Book review: The Winthrop Family in America by Lawrence Shaw Mayo. Isis, 41(1):127–128, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226408.

Sarton:1950:BRBe

[32] George Sarton. Book review: The Earliest Chemical Industry. An Essay in the Historical Relations of Economics & Technology Illustrated from the Alum Trade by Charles Singer. Isis, 41(1):128–131, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226409.

Koenig:1950:BRB

[33] Frederick O. Koenig. Book review: Short History of the Art of Distillation from the Beginnings up to the Death of Cellier Blumenthal by R. J. Forbes. Isis, 41(1):131–133, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226410.

Duveen:1950:BRB

[34] Denis Duveen. Book review: Crucibles: The Story of Chemistry from Ancient Alchemy to Nuclear Fission by Bernard Jaffe. Isis, 41(1):133– 134, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226411.

Albion:1950:BRB

[35] Robert G. Albion. Book review: The Story of the Ship by Charles E. Gibson. Isis, 41(1):134, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226412.

Zirkle:1950:BRBa

[36] Conway Zirkle. Book review: The Avian Egg by Alexis L. Romanoff; Anastasia J. Romanoff. Isis, 41(1):134–135, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/226413.

Montagu:1950:BRBa

[37] Ashley Montagu. Book review: A New Theory of Evolution by Arthur Keith. Isis, 41(1):135–136, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226414.

Forbes:1950:BRB

[38] R. J. Forbes. Book review: Geschiedenis der Kartografie van Nederland van den Romeinschen tijd tot het midden van de 19de eeuw by S. J. Fockema Andreae. Isis, 41(1):136–137, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226415.

Pagel:1950:BRBb

[39] Walter Pagel. Book review: Das Geheimnis der Seele. Ursprung und Ziel der Psychischen Energie by Esther Harding. Isis, 41(1):137–138, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226416.

Pagel:1950:BRBc

[40] Walter Pagel. Book review: Eranos-Jahrbuch 1947. Vol. XV: “Der Mensch”. Isis, 41(1):138, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226417.

Graubard:1950:BRBa

[41] Mark Graubard. Book review: Magic, a Sociological Study by Hutton Webster. Isis, 41(1):138–139, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226418.

Clark:1950:BRB

[42] Joseph T. Clark. Book review: A Guide to Historical Method by Gilbert J. Garraghan; Jean Delanglez. Isis, 41(1):139–143, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226419.

Pagel:1950:BRBd

[43] W. Pagel. Book review: British Art and the Mediterranean by F. Saxl; R. Wittkower. Isis, 41(1):143–144, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226420.

Howe:1950:BRB

[44] Mark DeWolfe Howe. Book review: Legal Philosophy from Plato to Hegel by Huntington Cairns. Isis, 41(1):144–145, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226421.

Saunders:1950:BRB

[45] John B. DeC. M. Saunders. Book review: On the Contributions of Hugh Owen Thomas of Liverpool, Sir Robert Jones of Liverpool and London, John Ridlon, M.D., of New York and Chicago, to Modern Orthopedic Surgery by H. Winnett Orr. Isis, 41(1):145–146, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226422.

Anonymous:1950:BMa

[46] Anonymous. Back matter. Isis, 41(1):??, March 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226423.

Anonymous:1950:FMb

[47] Anonymous. Front matter. Isis, 41(2):147–148, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227182.

Sarton:1950:SPV

[48] George Sarton. Second preface to volume 41 of Isis: Notes on the reviewing of learned books. Isis, 41(2):149–158, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227183.

Harper:1950:FPL

[49] Robert S. Harper. The first psychological laboratory. Isis, 41(2):158– 161, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227184.

Aldridge:1950:BFJ

[50] Alfred Owen Aldridge. Benjamin Franklin and Jonathan Edwards on lightning and earthquakes. Isis, 41(2):162–164, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227185.

VonHagen:1950:WFL

[51] Victor Wolfgang Von Hagen. Was this the fate of the library of Alexander von Humboldt? An inquiry. Isis, 41(2):164–167, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227186.

Duveen:1950:ALB

[52] Denis Duveen. Antoine Lavoisier’s Traité Élémentaire de Chimie. A bibliographical note. Isis, 41(2):168–171, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227187.

Dobell:1950:SHT

[53] Clifford Dobell. Samuel Hoole, translator of Leeuwenhoek’s Select Works: With notes on that publication. Isis, 41(2):171–180, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227188.

Kennedy:1950:FCP

[54] E. S. Kennedy. A Fifteenth-Century planetary computer: al-Kāshī’s “t.abaq al-Manāt.eq” i. motion of the Sun and Moon in longitude. Isis, 41(2):180–183, July 1950. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227189.

Hartner:1950:AIC

[55] Willy Hartner. The astronomical instruments of Cha-ma-lu-ting, their identification, and their relations to the instruments of the Observatory of Marāgha. Isis, 41(2):184–194, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227190.

Goodrich:1950:NC

[56] L. Carrington Goodrich, Alexandre Koyré, Lynn Thorndike, Martin Levey, Emmet Field Horine, G. S., and Brooke Hindle. Notes & correspondence. Isis, 41(2):194–198, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227191.

Abrams:1950:QA

[57] John W. Abrams, I. B. Cohen, George Sarton, Loren C. MacKinney, and Karl K. Darrow. Queries and answers. Isis, 41(2):198–201, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227192.

Anonymous:1950:Ob

[58] Anonymous. Obituaries. Isis, 41(2):201–202, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227193.

Sarton:1950:BRBf

[59] George Sarton. Book review: Saint Augustin et la Fin de la Culture Antique by Henri–Irénée Marrou. Isis, 41(2):202–204, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227194.

Sarton:1950:BRBg

[60] George Sarton. Book review: L’Essor de la littérature latine au XIIe siècle by Joseph De Ghellinck; Le Mouvement théologique du XIIe siècle. Sa préparation lointaine, avant et autour de Pierre Lombard, ses rapports avec les initiatives des Canonistes. Etudes, recherches et documents. Isis, 41(2):204–205, July 1950. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227195.

Sarton:1950:BRBh

[61] George Sarton. Book review: La Escala de Mahoma by José Muñoz Sendino. Isis, 41(2):206–207, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227196.

Sarton:1950:BRBi

[62] G. Sarton. Book review: The Tree of Battles by Honoré Bonet; G. W. Coopland. Isis, 41(2):207, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227197.

Dijksterhuis:1950:BRB

[63] E. J. Dijksterhuis. Book review: Die Vorläufer Galileis im 14. Jahrhundert. Studien zur Naturphilosophie der Spätscholastik by Anneliese Maier. Isis, 41(2):207–210, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227198.

OMalley:1950:BRB

[64] Charles D. O’Malley. Book review: The Epitome of Andreas Vesalius by L. R. Lind. Isis, 41(2):210–212, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227199.

Guerlac:1950:BRBa

[65] Henry Guerlac. Book review: The Endeavour of Jean Fernel, with a List of the Editions of His Writings by Charles Sherrington. Isis, 41(2): 212–213, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227200.

Sarton:1950:BRBj

[66] George Sarton. Book review: Petrus Plancius, Theoloog en Geograaf (1552–1622) by J. Keuning. Isis, 41(2):213–214, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227201.

Stimson:1950:BRB

[67] Dorothy Stimson. Book review: Francis Bacon, Philosopher of Industrial Science by Benjamin Farrington. Isis, 41(2):215–216, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227202.

Hellman:1950:BRBc

[68] C. Doris Hellman. Book review: Johannes Kepler by Max Caspar. Isis, 41(2):216–219, July 1950. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227203.

Koyre:1950:BRBb

[69] Alexandre Koyre. Book review: The Naming of the Telescope by Edward Rosen. Isis, 41(2):219–220, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227204.

Sarton:1950:BRBk

[70] George Sarton. Book review: Galileo in Cina. Relazioni attraverso il Collegio Romano tra Galileo e i gesuiti scienziati missionari in Cina (1610– 1640) by Pasquale M. D’Elia. Isis, 41(2):220–222, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227205.

Hornberger:1950:BRB

[71] Theodore Hornberger. Book review: Amiable Autocrat, a Biography of Dr. Oliver Wendell Holmes [1809–94] by Eleanor M. Tilton. Isis, 41 (2):222, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227206.

Oldenberg:1950:BRC

[72] Otto Oldenberg. Book review: Carl Runge, und sein wissenschaftliches Werk by Iris Runge. Isis, 41(2):222–224, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227207.

Guerlac:1950:BRA

[73] Henry Guerlac. Book review: Alsos by Samuel A. Goudsmit. Isis, 41(2): 224–227, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227208.

Sarton:1950:BRBl

[74] George Sarton. Book review: Essai sur la formation de la pensée grecque. Introduction historique à une étude de la philosophie platonicienne by Pierre-Maxime Schuhl. Isis, 41(2):227–228, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227209.

Usher:1950:BRB

[75] Abbott Payson Usher. Book review: British Medieval Population by Josiah Cox Russell. Isis, 41(2):228–229, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227210.

Sarton:1950:BRBm

[76] George Sarton. Book review: Estudios sobre historia de la ciencia española by José Maria Millás Vallicrosa. Isis, 41(2):229, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227211.

Veith:1950:BRB

[77] Ilza Veith. Book review: “Heilkunde im Alten China” by Willy Hartner. Isis, 41(2):230, July 1950. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227212.

Gandz:1950:BRBb

[78] Solomon Gandz. Book review: History of New Testament Times. With an Introduction to the Apocrypha by Robert H. Pfeiffer. Isis, 41(2):230– 231, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227213.

Cohen:1950:BRBd

[79] I. Bernard Cohen. Book review: The Origins of Modern Science by Herbert Butterfield. Isis, 41(2):231–233, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227214.

Hindle:1950:BRB

[80] Brooke Hindle. Book review: Some Early Tools of American Science: An Account of the Early Scientific Instruments and Mineralogical and Biological Collections in Harvard University by I. Bernard Cohen. Isis, 41(2):233–234, July 1950. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227215.

Boyer:1950:BRB

[81] Carl B. Boyer. Book review: The Mathematics of Great Amateurs by Julian Lowell Coolidge. Isis, 41(2):234–236, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227216.

Lenzen:1950:BRBa

[82] V. F. Lenzen. Book review: Philosophy of Mathematics and Natural Science by Hermann Weyl. Isis, 41(2):236–237, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227217.

Miles:1950:BRB

[83] Wyndham Miles. Book review: The Alchemists: Founders of Modern Chemistry by F. Sherwood Taylor. Isis, 41(2):237–238, July 1950. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227218.

Sax:1950:BRBa

[84] Karl Sax. Book review: The Death of a Science in Russia by Conway Zirkle. Isis, 41(2):238–239, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227219.

Sax:1950:BRBb

[85] Karl Sax. Book review: Heredity — East and West by Julian Huxley. Isis, 41(2):239–240, July 1950. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227220.

Zirkle:1950:BRBb

[86] Conway Zirkle. Book review: The Nature-Nurture Controversy by Nicholas Pastore. Isis, 41(2):240–242, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227221.

Zirkle:1950:BRBc

[87] Conway Zirkle. Book review: Liberty Hyde Bailey. A Story of Plant Science by Andrew Denny Rogers. Isis, 41(2):242–243, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227222.

Raisz:1950:BRB

[88] Erwin Raisz. Book review: The Story of Maps by Lloyd A. Brown. Isis, 41(2):243–244, July 1950. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227223.

Diller:1950:BRB

[89] Aubrey Diller. Book review: History of Ancient Geography by J. Oliver Thomson. Isis, 41(2):244–245, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227224.

Neugebauer:1950:BRB

[90] O. Neugebauer. Book review: Worlds in Collision by Immanuel Velikovsky. Isis, 41(2):245–246, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227225.

Sarton:1950:BRBn

[91] G. Sarton. Book review: The Lapps by Björn Collinder. Isis, 41(2):246– 247, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227226.

Ulich:1950:BRB

[92] Robert Ulich. Book review: The Social Function of Art by Radhakamal Mukerjee. Isis, 41(2):247, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227227.

Pagel:1950:BRBe

[93] Walter Pagel. Book review: Europaeische Literatur und Lateinisches Mittelalter by Ernst Robert Curtius. Isis, 41(2):247–248, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227228.

Lenzen:1950:BRBb

[94] V. F. Lenzen. Book review: Philosophy for the Future; The Quest of Modern Materialism by Roy Wood Sellars; V. J. McGill; Marvin Farber. Isis, 41(2):248–249, July 1950. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227229.

Jackman:1950:BRBa

[95] Sydney W. Jackman. Book review: The Philosophy of Civilization by Albert Schweitzer. Isis, 41(2):250, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227230.

Patterson:1950:BRB

[96] Louise Diehl Patterson. Book review: Scientists and Amateurs. A History of the Royal Society by Dorothy Stimson. Isis, 41(2):251–252, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227231.

Goldsmith:1950:BRB

[97] Albert R. Goldsmith. Book review: Sons of Science. The Story of the Smithsonian Institution and Its Leaders by Paul H. Oehser. Isis, 41(2): 252–253, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227232.

Pogo:1950:AD

[98] Alexander Pogo. Administrative documents. Isis, 41(2):253–256, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227233.

Anonymous:1950:BMb

[99] Anonymous. Back matter. Isis, 41(2):257, July 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227234.

Anonymous:1950:FMc

[100] Anonymous. Front matter. Isis, 41(3–4):259–260, December 1950. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227047.

Boas:1950:BTS

[101] Marie Boas. Boyle as a theoretical scientist. Isis, 41(3–4):261–268, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227048.

Hobbs:1950:VVA

[102] William Herbert Hobbs. Verrazano’s voyage along the North American coast in 1524. Isis, 41(3–4):268–277, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227049.

Luther:1950:EEM

[103] Frederic Luther. The earliest experiments in microphotography. Isis, 41 (3–4):277–281, December 1950. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227050.

Duveen:1950:JPC

[104] Denis Duveen. James price (1752–1783) chemist and alchemist. Isis, 41 (3–4):281–283, December 1950. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227051.

Thorndike:1950:BBX

[105] Lynn Thorndike. The Tables of Barcelona of the XIVth Century. Isis, 41 (3–4):283–285, December 1950. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227052.

Zinner:1950:CPA

[106] E. Zinner. Cl. Ptolemaeus und das Astrolab. (German) [Cl. Ptolemy and the astrolabe]. Isis, 41(3–4):286–287, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227053.

Rosen:1950:TGS

[107] Edward Rosen. The title of Galileo’s sidereus nuncius. Isis, 41(3–4): 287–289, December 1950. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227054.

Dieckmann:1950:ARD

[108] Herbert Dieckmann. The autopsy report on Diderot. Isis, 41(3–4):289– 290, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227055.

Sarton:1950:CBB

[109] George Sarton. The critical bibliographies of Isis. Isis, 41(3–4):291–298, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227056.

Francon:1950:NRN

[110] Marcel Françon. Note sur Rabelais et les nombres. (French) [Note on Rabelais and numbers]. Isis, 41(3–4):298–300, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227057.

Sarton:1950:QA

[111] George Sarton. Queries and answers. Isis, 41(3–4):300–301, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227058.

Anonymous:1950:Oc

[112] Anonymous. Obituaries. Isis, 41(3–4):301–303, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227059.

Beal:1950:NC

[113] Edwin C. Beal, Jr. and Louise D. Patterson. Notes and correspondence. Isis, 41(3–4):303–305, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227060.

Anonymous:1950:Pb

[114] Anonymous. Personalia. Isis, 41(3–4):305, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227061.

Sarton:1950:BRBo

[115] George Sarton. Book review: Corpvs commentariorvm Averrois in Aristotelem by Henricvs Avstryn Wolfson; David Baneth; Franciscvs Howard Fobes. Isis, 41(3–4):306–307, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227062.

Sarton:1950:BRBp

[116] G. Sarton. Book review: A Descriptive Catalogue of the Grace K. Babson Collection of the Works of Sir Isaac Newton and the Material Relating to Him in the Babson Institute Library, Babson Park, Mass by Henry P. Macomber. Isis, 41(3–4):307–308, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227063.

Johnson:1950:BRB

[117] Francis R. Johnson. Book review: All Coherence Gone by Victor Harris. Isis, 41(3–4):308–309, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227064.

Struik:1950:BRB

[118] D. J. Struik. Book review: Die Entwicklungsgeschichte der Leibnizschen Mathematik während des Aufenthaltes in Paris (1672–1676) by J. E. Hofmann. Isis, 41(3–4):309–310, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227065.

Rosenthal-Schneider:1950:BRB

[119] Ilse Rosenthal-Schneider. Book review: Vorträge der aus Anlass seines 300. Geburtstages in Hamburg abgehaltenen wissenschaftlichen Tagung by Gottfried Wilhelm Leibniz. Isis, 41(3–4):310–313, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227066.

Farber:1950:BRBa

[120] Eduard Farber. Book review: Robert Wilhelm Bunsen. Lebensbild eines deutschen Naturforschers by Georg Lockemann. Isis, 41(3–4):313–314, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227067.

Brozek:1950:BRB

[121] Josef Brožek. Book review: James McKeen Cattell, Man of Science, 1860–1944 by A. T. Poffenberger. Isis, 41(3–4):314–315, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227068.

Farber:1950:BRBb

[122] Eduard Farber. Book review: Aus meinem Leben. Von Arbeit, Musse und Freunden by Richard Willstätter. Isis, 41(3–4):315–316, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227069.

Mates:1950:BRB

[123] Benson Mates. Book review: La logique et l’épistémologie des Stoïciens, leurs rapports avec la logique d’Aristote, la logistique et la pensée contemporaines by Antoinette Virieux-Reymond. Isis, 41(3–4):316–317, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227070.

Rie:1950:BRB

[124] Robert Rie. Book review: Wunder und Zeichen. Grosse Gestalten des Hochmittelalters by Friderike Maria Zweig. Isis, 41(3–4):317–318, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227071.

Nykl:1950:BRB

[125] A. R. Nykl. Book review: Inscripções é Tradições da América Prehistórica especialmente do Brasil by Bernardo De Azevedo Da Silva Ramos. Isis, 41(3–4):318, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227072.

Zirkle:1950:BRBd

[126] Conway Zirkle. Book review: The Conflict of Science and Society by C. D. Darlington. Isis, 41(3–4):319, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227073.

Graubard:1950:BRBb

[127] Mark Graubard. Book review: Studies in Philosophy and Science by Morris R. Cohen. Isis, 41(3–4):319–320, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227074.

Jackman:1950:BRBb

[128] Sydney W. Jackman. Book review: Marxism & Contemporary Science by Jack Lindsay. Isis, 41(3–4):320, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227075.

Zirkle:1950:BRBe

[129] Conway Zirkle. Book review: Factors of Evolution by I. I. Schmalhausen; Isadore Dordick; Theodosius Dobzhansky. Isis, 41(3–4):320–321, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227076.

Montagu:1950:BRBb

[130] Ashley Montagu. Book review: The Meaning of Evolution by George Gaylord Simpson. Isis, 41(3–4):321–322, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227077.

Rickett:1950:BRB

[131] H. W. Rickett. Book review: The Natural Philosophy of Plant Form by Agnes Arber. Isis, 41(3–4):322–323, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227078.

Schrecker:1950:BRB

[132] Paul Schrecker. Book review: Civilisation traditionnelle et Genres de Vie by Andre Varagnac. Isis, 41(3–4):324–325, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227079.

Page:1950:BRB

[133] Charles H. Page. Book review: Social Theory and Social Structure. Toward the Codification of Theory and Research by Robert K. Merton. Isis, 41(3–4):325–326, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227080.

Jackman:1950:BRBc

[134] Sydney W. Jackman. Book review: Christianity and History by Herbert Butterfield. Isis, 41(3–4):326–327, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227081.

Leikind:1950:BRB

[135] Morris C. Leikind. Book review: Malaria, the Biography of a Killer by Leon J. Warshaw. Isis, 41(3–4):327, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227082.

Sarton:1950:SSC

[136] George Sarton and Frances Siegel. Seventy-sixth critical bibliography of the history and philosophy of science and of the history of civilization (to may 1950). Isis, 41(3–4):328–424, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227083.

Anonymous:1950:VI

[137] Anonymous. Volume information. Isis, 41(3–4):425–431, December 1950. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227046.

Anonymous:1950:BMc

[138] Anonymous. Back matter. Isis, 41(3–4):432–434, December 1950. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227084.

Anonymous:1951:FMa

[139] Anonymous. Front matter. Isis, 42(1):1–51, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/226657.

Sarton:1951:PVS

[140] George Sarton. Preface to volume 42: Science and peace. The development of international law. Isis, 42(1):3–9, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226658.

Levey:1951:FAM

[141] Martin Levey. The first American museum of natural history. Isis, 42(1): 10–12, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226659.

Gwei-Djen:1951:CHC

[142] Lu Gwei-Djen and Joseph Needham. A contribution to the history of Chinese dietetics. Isis, 42(1):13–20, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226660.

Dobell:1951:MOB

[143] Clifford Dobell. In memoriam: Otto Bütschli (1848–1920) “architect of protozoology”. Isis, 42(1):20–22, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226661.

Pagel:1951:WHP

[144] Walter Pagel. William Harvey and the purpose of circulation. Isis, 42(1): 22–38, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226662.

Zinner:1951:MAD

[145] E. Zinner. Magister Alard von Diest und die Pariser Beobachtungen von 1312–15. (German) [Master Alard von Diest and the Paris observations 1312–15]. Isis, 42(1):38–40, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226663.

Goldstein:1951:NC

[146] Herbert Goldstein, Erika von Erhardt-Siebold, G. S., Richard L. Schoenwald, and I. B. C. Notes and correspondence. Isis, 42(1):41–46, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226664.

Anonymous:1951:Pa

[147] Anonymous. Personalia. Isis, 42(1):46, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226665.

Cohen:1951:QA

[148] I. B. Cohen and George Sarton. Queries & answers. Isis, 42(1):47– 48, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226666.

Anonymous:1951:Oa

[149] Anonymous. Obituaries. Isis, 42(1):48–49, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226667.

Graubard:1951:BRBa

[150] Mark Graubard. Book review: Brutus: On the Nature of the Gods; On Divination; On Duties by Marcus Tullius Cicero; Hubert M. Poteat. Isis, 42(1):51–52, April 1951. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226668.

Sarton:1951:BRBa

[151] George Sarton. Book review: Roman Book on Precious Stones. Including on English Modernization of the 37th Booke of the Historie of the World by Plinius Secundus by Sydney H. Ball. Isis, 42(1):52–53, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226669.

Montagu:1951:BRBa

[152] M. F. Ashley Montagu. Book review: The Illustrations from the Works of Andreas Vesalius of Brussels by J. B. DeC. M. Saunders; Charles D. O’Malley. Isis, 42(1):53, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226670.

Brown:1951:BRBa

[153] Harcourt Brown. Book review: The Universe of Pontus de Tyard by John C. Lapp. Isis, 42(1):53–54, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226671.

Cohen:1951:BRBa

[154] I. Bernard Cohen. Book review: Bergwerk- und Probierbüchlein by Anneliese Grünhaldt Sisco; Cyril Stanley Smith; De Re Metallica by Herbert Clark Hoover; Louhenry Hoover; Georgius Agricola. Isis, 42(1):54–56, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226672.

Sarton:1951:BRBb

[155] George Sarton. Book review: Oeuvres complètes de Christiaan Huygens by Christiaan Huygens. Isis, 42(1):56–57, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226673.

Brown:1951:BRBb

[156] Harcourt Brown. Book review: Science and Rationalism in the Government of Louis XIV, 1661–1683 by James E. King. Isis, 42(1):57–58, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226674.

Sarton:1951:BRBc

[157] George Sarton. Book review: William Stukeley. An Eighteenth-Century Antiquary by Stuart Piggott. Isis, 42(1):58–60, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226675.

Montagu:1951:BRBb

[158] M. F. Ashley Montagu. Book review: An Autobiography by Arthur Keith. Isis, 42(1):61, April 1951. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226676.

Boyer:1951:BRB

[159] Carl B. Boyer. Book review: De Pythagore à Euclide. Contribution à l’histoire des mathématiques preeuclidiennes by Paul-Henri Michel. Isis, 42(1):61–63, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/226677.

Boas:1951:BRB

[160] Marie Boas. Book review: Ktesibios Philon and Heron. A Study in Ancient Pneumatics by A. G. Drachmann. Isis, 42(1):63, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226678.

Sarton:1951:BRBd

[161] George Sarton. Book review: Storia delle matematiche dall’alba della cività al secolo XIX by Gino Loria. Isis, 42(1):63–64, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226679.

Embree:1951:BRB

[162] John F. Embree. Book review: Histoire de la Medecine du Viêtnam by Bành, Du’O’Ng Bá. Isis, 42(1):64–65, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226680.

Lenzen:1951:BRBa

[163] V. F. Lenzen. Book review: Harvard Case Histories in Experimental Science by James Bryant Conant; Duane Roller; Leonard K. Nash. Isis, 42(1):65–66, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/226681.

Struik:1951:BRBa

[164] D. J. Struik. Book review: De mechanisering van het wereldbeeld by E. J. Dijksterhuis. Isis, 42(1):66–67, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226682.

Farber:1951:BRB

[165] Eduard Farber. Book review: Historic Researches. Chapters in the History of Physical and Chemical Discovery by T. W. Chalmers. Isis, 42(1): 67–68, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226683.

Frank:1951:BRB

[166] Philipp Frank. Book review: The Basis and Structure of Knowledge by W. H. Werkmeister. Isis, 42(1):68–69, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226684.

Lenzen:1951:BRBb

[167] V. F. Lenzen. Book review: The Nature of Physical Reality. A Philosophy of Modern Physics by Henry Margenau. Isis, 42(1):69–70, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226685.

Lenzen:1951:BRBc

[168] V. F. Lenzen. Book review: The Evolution of Scientific Thought. From Newton to Einstein by A. d’Abro. Isis, 42(1):70–71, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226686.

Leblanc:1951:BRB

[169] Hugues Leblanc. Book review: Le déclin des absolus mathématicologiques by Georges Bouligand; Jean Desgranges. Isis, 42(1):71–72, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226687.

Jacchia:1951:BRB

[170] L. Jacchia. Book review: Storia dell’Astronomia by G. Abetti. Isis, 42 (1):72, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226688.

Streeter:1951:BRB

[171] John W. Streeter. Book review: A Concise History of Astronomy by Peter Doig. Isis, 42(1):73–75, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226689.

Sarton:1951:BRBe

[172] George Sarton. Book review: Ancient Near Eastern Texts Relating to the Old Testament by James B. Pritchard. Isis, 42(1):75–76, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226690.

Cohen:1951:BRBb

[173] I. Bernard Cohen. Book review: Albert Einstein: Philosopher-Scientist by Paul Arthur Schilpp. Isis, 42(1):76–79, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226691.

Smith:1951:BRB

[174] Cyril Stanley Smith. Book review: Man the Maker. A History of Technology and Engineering by R. J. Forbes. Isis, 42(1):79–80, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226692.

Zirkle:1951:BRBa

[175] Conway Zirkle. Book review: Selected Works by Ivan Vladimirovich Michurin. Isis, 42(1):80–81, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226693.

Zirkle:1951:BRBb

[176] Conway Zirkle. Book review: A History of Biology by Charles Singer. Isis, 42(1):82, April 1951. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226694.

Arber:1951:BRB

[177] Agnes Arber. Book review: Maize in the Great Herbals by J. J. Finan. Isis, 42(1):82–83, April 1951. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226695.

Zirkle:1951:BRBc

[178] Conway Zirkle. Book review: Variation and Evolution in Plants by G. Ledyard Stebbins, Jr. Isis, 42(1):83–84, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226696.

Zirkle:1951:BRBd

[179] Conway Zirkle. Book review: Index Rafinesquianus. The Plant Names Published by C. S. Rafinesque with Reductions, and a Consideration of his Methods, Objectives and Attainments by Elmer D. Merrill. Isis, 42 (1):84, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226697.

Zirkle:1951:BRBe

[180] Conway Zirkle. Book review: The History and Social Influence of the Potato by Redcliff N. Salaman. Isis, 42(1):85, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226698.

Zirkle:1951:BRBf

[181] Conway Zirkle. Book review: A People’s Academy by Gennadi Fish. Isis, 42(1):85–86, April 1951. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226699.

Zirkle:1951:BRBg

[182] Conway Zirkle. Book review: America’s Crop Heritage: The History of Foreign Plant Introduction by the Federal Government by Nelson Klose.

Isis, 42(1):86, April 1951. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226700.

Brozek:1951:BRBa

[183] Josef Brožek. Book review: A History of Experimental Psychology by Edwin G. Boring. Isis, 42(1):87–88, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226701.

Cohen:1951:BRBc

[184] I. Bernard Cohen. Book review: Ideas & Men, The Story of Western Thought by Crane Brinton; English Political Thought in the Nineteenth Century. Isis, 42(1):88–89, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226702.

deSantillana:1951:BRB

[185] G. de Santillana. Book review: Four Views of Time in Ancient Philosophy by John F. Callahan. Isis, 42(1):89–90, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226703.

Graubard:1951:BRBb

[186] Mark Graubard. Book review: On Being Human by Ashley Montagu. Isis, 42(1):90–91, April 1951. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226704.

Barber:1951:BRB

[187] Bernard Barber. Book review: Science Is a Sacred Cow by Anthony Standen. Isis, 42(1):91–92, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226705.

Montagu:1951:BRBc

[188] M. F. Ashley Montagu. Book review: Human Rights. Comments and Interpretations by UNESCO. Isis, 42(1):92, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/226706.

Montagu:1951:BRBd

[189] M. F. Ashley Montagu. Book review: The Idea of Conflict by Kurt Singer. Isis, 42(1):92–93, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226707.

Anonymous:1951:BMa

[190] Anonymous. Back matter. Isis, 42(1):94, April 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226708.

Anonymous:1951:FMb

[191] Anonymous. Front matter. Isis, 42(2):95–96, June 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/226963.

Zirkle:1951:GMH

[192] Conway Zirkle. Gregor Mendel & his precursors. Isis, 42(2):97–104, June 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226964.

Dupree:1951:SLC

[193] A. Hunter Dupree. Some letters from Charles Darwin to Jeffries Wyman. Isis, 42(2):104–110, June 1951. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226965.

Neugebauer:1951:SWS

[194] O. Neugebauer. The study of wretched subjects. Isis, 42(2):111, June 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226966.

deSantillana:1951:PLW

[195] George de Santillana and Walter Pitts. Philolaos in limbo, or: What happened to the Pythagoreans? Isis, 42(2):112–120, June 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226967.

Patterson:1951:LTD

[196] Louise Diehl Patterson. Leonard and Thomas Digges: Biographical notes. Isis, 42(2):120–121, June 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226968.

Carmody:1951:RNA

[197] Francis J. Carmody. Regiomontanusp’ notes on al-Bit.rûjî’s astronomy. Isis, 42(2):121–130, June 1951. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226969.

Thorndike:1951:S

[198] Lynn Thorndike. Sexagenarium. Isis, 42(2):130–133, June 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226970.

Schove:1951:SAB

[199] D. Justin Schove. Sunspots, aurorae and blood rain: The spectrum of time. Isis, 42(2):133–138, June 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226971.

Winter:1951:SOR

[200] H. J. J. Winter and W. ’Arafat. A statement on optical reflection and “refraction” attributed to Nas.īr ud–Dīn at.-T .ūsī. Isis, 42(2):138–142, June 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226972.

Harvey:1951:QAa

[201] E. Newton Harvey. Queries & answers. Isis, 42(2):142–143, June 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226973.

Anonymous:1951:Pb

[202] Anonymous. Personalia. Isis, 42(2):143, June 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226974.

Yutaka:1951:NC

[203] Hirata Yutaka and L. Carrington Goodrich. Notes and correspondence. Isis, 42(2):143–145, June 1951. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226975.

Anonymous:1951:Ob

[204] Anonymous. Obituaries. Isis, 42(2):145–147, June 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/226976.

Sarton:1951:BRBf

[205] George Sarton. Book review: Autolycus de Pitane. Histoire du texte suivie de l’édition critique des traités de la sphère en mouvement et des levers et couchers by Joseph Mogenet. Isis, 42(2):147–148, June 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226977.

Sarton:1951:BRBg

[206] George Sarton and I. E. Drabkin. Book review: Caelius Aurelianus. On Acute Diseases and on Chronic Diseases by I. E. Drabkin. Isis, 42(2): 148–150, June 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226978.

Sarton:1951:BRBh

[207] George Sarton. Book review: Aetii Amideni Libri medicinales V–VIII by Alexander Olivieri; Aëtios of Amida: The Gynaecology and Obstetrics of the VIth Century A.D. by James V. Ricci. Isis, 42(2):150–152, June 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226979.

Sarton:1951:BRBi

[208] George Sarton. Book review: Estudios sobre Azarquiel by José Maria Millás Vallicrosa. Isis, 42(2):152–153, June 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226980.

Marx:1951:BRB

[209] Leo Marx. Book review: Jonathan Edwards by Perry Miller. Isis, 42(2): 153–156, June 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226981.

Hartner:1951:BRBa

[210] Willy Hartner and I. Bernard Cohen. Book review: Goethe. Dichtung, Wissenschaft, Weltbild by Karl Viëtor; Goethe the Poet. Goethe the Thinker by Karl Viëtor; Moses Hadas; Bayard Q. Morgan. Isis, 42(2): 156–157, June 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226982.

Hartner:1951:BRBb

[211] Willy Hartner. Book review: Biologie der Goethezeit by Adolf Meyer-Abich. Isis, 42(2):157–160, June 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226983.

Urdang:1951:BRBa

[212] George Urdang. Book review: Goethe as a Scientist by Rudolf Magnus; Heinz Norden. Isis, 42(2):160, June 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226984.

Owen:1951:BRB

[213] Gwilym E. Owen. Book review: Joseph Henry. His Life and Work by Thomas Coulson. Isis, 42(2):160–162, June 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226985.

Bobrovnikoff:1951:BRB

[214] N. T. Bobrovnikoff. Book review: Latin Treatises on Comets between 1238 and 1368 A.D. by Lynn Thorndike. Isis, 42(2):162–163, June 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226986.

Sarton:1951:BRBj

[215] George Sarton. Book review: The Western World and Japan. A Study in the Interaction of European and Asiatic Cultures by G. B. Sansom. Isis, 42(2):163–164, June 1951. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226987.

Winter:1951:BRB

[216] H. J. J. Winter. Book review: El Enigma de la Luz by Cortes Pla. Isis, 42 (2):164, June 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226988.

Zirkle:1951:BRBh

[217] Conway Zirkle. Book review: The Nature of Natural History by Marston Bates. Isis, 42(2):164–165, June 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226989.

Zirkle:1951:BRBi

[218] Conway Zirkle. Book review: Fifty Years of Plant Physiology by Th. Weevers. Isis, 42(2):165, June 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226990.

Urdang:1951:BRBb

[219] George Urdang. Book review: Geschichte der Medizin. Die historische Entwicklung der Heilkunde und des ärztlichen Lebens. I. Band: Von den Anfängen der Medizin bis zur Mitte des 18. Jahrhunderts by Paul Diepgen. Isis, 42(2):166–167, June 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226991.

Anonymous:1951:BMb

[220] Anonymous. Back matter. Isis, 42(2):168–172, June 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226992.

Sarton:1951:SPV

[221] George Sarton. Second preface to volume 42. Science and peace II. The botanical exploration of Malaysia. Isis, 42(3):173–176, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226555.

Fulton:1951:ISA

[222] John F. Fulton. The impact of science on American history. Isis, 42 (3):176–191, October 1951. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226556.

Foote:1951:PSB

[223] George A. Foote. The place of science in the British reform movement 1830–50. Isis, 42(3):192–208, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226557.

Patterson:1951:RC

[224] Louise Diehl Patterson. Recorde’s cosmography, 1556. Isis, 42(3):208– 218, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226558.

Hall:1951:TUL

[225] A. R. Hall. Two unpublished lectures of Robert Hooke. Isis, 42(3):219– 230, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226559.

Macomber:1951:CVE

[226] Henry P. Macomber. A comparison of the variations and errors in copies of the first edition of Newton’s principia, 1687. Isis, 42(3):230–232, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226560.

Duveen:1951:ALL

[227] Denis Duveen. Antoine Laurent Lavoisier (1743–1794): A note regarding his domicile during the French Revolution. Isis, 42(3):233–234, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226561.

Garnier:1951:EBR

[228] Frère Garnier. Edmond Brunhes (1834–1916) (en religion, Frère Gabriel-Marie, de l’Institut des Frères des Écoles Chrétiennes). Isis, 42(3):234– 237, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226562.

Sarton:1951:BDW

[229] George Sarton. Bibliography of Dr e. w. Gudger’s contributions to the history of ichthyology (1905–1951). Isis, 42(3):237–242, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226563.

Francon:1951:NC

[230] Marcel Françon and George Sarton. Notes and correspondence. Isis, 42 (3):242–243, October 1951. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226564.

Anonymous:1951:Pc

[231] Anonymous. Personalia. Isis, 42(3):243, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226565.

Anonymous:1951:CHS

[232] Anonymous. Courses in the history of science. Isis, 42(3):244, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226567.

Anonymous:1951:DP

[233] Anonymous. Dissertations in progress. Isis, 42(3):244, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226566.

Pagel:1951:BRB

[234] Walter Pagel. Book review: Paracelsus. Magic into Science by Henry M. Pachter. Isis, 42(3):244–246, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226568.

Saunders:1951:BRBa

[235] J. B. deC. M. Saunders. Book review: The Life and Times of Gaspare Tagliacozzi, Surgeon of Bologna, 1545–1599, with a Documented Study of the Scientific and Cultural Life of Bologna in the Sixteenth Century by Martha Teach Gnudi; Jerome Pierce Webster. Isis, 42(3):246–247, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226569.

Graubard:1951:BRBc

[236] Mark Graubard. Book review: Giordano Bruno, His Life and Thought. On the Infinite Universe and Worlds by Dorothea Waley Singer. Isis, 42(3):247–248, October 1951. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226570.

Jackman:1951:BRB

[237] Sydney W. Jackman. Book review: Sir Thomas Browne by William P. Dunn. Isis, 42(3):248–249, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226571.

Chalmers:1951:BRB

[238] Gordon Keith Chalmers. Book review: Sir Thomas Browne, a Doctor’s Life of Science and Faith by Jeremiah S. Finch. Isis, 42(3):249–250, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226572.

Rosenberg:1951:BRB

[239] Albert Rosenberg. Book review: The Life of Dr. John Radcliffe 1652– 1714, Benefactor of the University of Oxford by Campbell R. Hone. Isis, 42(3):250–251, October 1951. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226573.

Brown:1951:BRBc

[240] Harcourt Brown. Book review: The Breaking of the Circle by Marjorie Hope Nicolson. Isis, 42(3):251–252, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226574.

Hellman:1951:BRB

[241] C. Doris Hellman. Book review: Gesammelte Werke by Johannes Kepler. Isis, 42(3):252–255, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226575.

Foote:1951:BRB

[242] George A. Foote. Book review: Genesis and Geology. A Study in the Relations of Scientific Thought, Natural Theology, and Social Opinion in Great Britain, 1790–1850 by Charles Coulston Gillispie. Isis, 42(3):255– 256, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226576.

Zirkle:1951:BRBj

[243] Conway Zirkle. Book review: John William Draper and the Religion of Science by Donald Fleming; W. E. Calvert. Isis, 42(3):256–257, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226577.

Zirkle:1951:BRBk

[244] Conway Zirkle. Book review: Bernard Eduard Fernow, a Story of North American Forestry by Andrew Denny Rogers. Isis, 42(3):257, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226578.

Struik:1951:BRBb

[245] D. J. Struik. Book review: Josiah Willard Gibbs: The History of a Great Mind by Lynde Phelps Wheeler. Isis, 42(3):257–259, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226579.

Gandz:1951:BRBa

[246] Solomon Gandz. Book review: The Burden of Egypt. An Interpretation of Ancient Egyptian Culture by John A. Wilson. Isis, 42(3):259–260, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226580.

Gandz:1951:BRBb

[247] Solomon Gandz. Book review: The Calendars of Ancient Egypt by Richard A. Parker. Isis, 42(3):260–263, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226581.

Sarton:1951:BRBk

[248] George Sarton. Book review: The Ancient Khmer Empire by Lawrence Palmer Briggs. Isis, 42(3):263–265, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226582.

Hartner:1951:BRBc

[249] Willy Hartner. Book review: Huang Ti Nei Ching Su Wên. The Yellow Emperor’s Classic of Internal Medicine by Ilza Veith. Isis, 42(3):265– 266, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226583.

Gandz:1951:BRBc

[250] Solomon Gandz. Book review: Hellenism in Jewish Palestine. Studies in Literary Transmission, Beliefs and Manners of Palestine in the I Century B.C.E–IV Century C.E. by Saul Lieberman. Isis, 42(3):266–267, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226584.

Gandz:1951:BRBd

[251] Solomon Gandz. Book review: The Mysterious Numbers of the Hebrew Kings. A Reconstruction of the Chronology of the Kingdoms of Israel and Judah by Edwin R. Thiele. Isis, 42(3):267–268, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226585.

Sarton:1951:BRBl

[252] George Sarton. Book review: Maya Hieroglyphic Writing by J. Eric S. Thompson. Isis, 42(3):268–269, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226586.

Zirkle:1951:BRBl

[253] Conway Zirkle. Book review: Science and Common Sense by James B. Conant. Isis, 42(3):269–271, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226587.

Cohen:1951:BRBd

[254] I. Bernard Cohen. Book review: Histoire de la mécanique. Préface de Louis de Broglie by Réne Dugas. Isis, 42(3):271–272, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226588.

Cohen:1951:BRBe

[255] I. Bernard Cohen. Book review: Sourcebook on Atomic Energy by Samuel Glasstone; Foundations of Nuclear Physics by Robert T. Beyer; The Atom at Work by Jacob Sacks; New Atoms, Progress and Some Memories by Otto Hahn; W. Gaade; A Hundred Years of Physics by William Wilson. Isis, 42(3):272–273, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226589.

Zirkle:1951:BRBm

[256] Conway Zirkle. Book review: Water, Land and People by Bernard Frank; Anthony Netboy. Isis, 42(3):273–274, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226590.

Zirkle:1951:BRBn

[257] Conway Zirkle. Book review: Dragons in Amber. Further Adventures of a Romantic Naturalist by Willy Ley. Isis, 42(3):274–275, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226591.

Zirkle:1951:BRBo

[258] Conway Zirkle. Book review: Evolutionary Thought in America by Stow Person. Isis, 42(3):275, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226592.

Gelfan:1951:BRB

[259] Samuel Gelfan. Book review: The History of Muscle Physiology by E. Bastholm. Isis, 42(3):276, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226593.

Brozek:1951:BRBb

[260] Josef Brožek. Book review: Readings in the History of Psychology by Wayne Denis. Isis, 42(3):276–277, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226594.

Werkmeister:1951:BRB

[261] W. H. Werkmeister. Book review: The Rise of Scientific Philosophy by Hans Reichenbach. Isis, 42(3):277–278, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/226595.

Saunders:1951:BRBb

[262] J. B. deC. M. Saunders. Book review: A History of Medicine. Volume I by Henry E. Sigerist. Isis, 42(3):278–281, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226596.

Anonymous:1951:BMc

[263] Anonymous. Back matter. Isis, 42(3):282, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/226597.

Anonymous:1951:FMc

[264] Anonymous. Front matter. Isis, 42(3):??, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/226554.

Anonymous:1951:FMd

[265] Anonymous. Front matter. Isis, 42(4):283, December 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226805.

Baker:1951:RDC

[266] John R. Baker. Remarks on the discovery of cell-division. Isis, 42(4): 285–287, December 1951. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226806.

Reilly:1951:SAA

[267] Desmond Reilly. Salts, acids & alkalis in the 19th Century. A comparison between advances in France, England & Germany. Isis, 42(4):287–296, December 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226807.

Kuhn:1951:NQD

[268] Thomas S. Kuhn. Newton’s “31st query” and the degradation of gold. Isis, 42(4):296–298, December 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226808.

Pelseneer:1951:PCC

[269] Jean Pelseneer. Petite contribution à la connaissance de Mariotte. (French) [Small contribution to the knowledge of Mariotte]. Isis, 42 (4):299–301, December 1951. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226809.

Thorndike:1951:PEF

[270] Lynn Thorndike. Prediction of eclipses in the Fourteenth Century. Isis, 42(4):301–302, December 1951. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226810.

Francon:1951:APN

[271] Marcel Françon. Ausone et le premier nombre parfait. Isis, 42(4):302– 303, December 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226811.

Anonymous:1951:Oc

[272] Anonymous. Obituaries. Isis, 42(4):304–306, December 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226812.

Shryock:1951:NC

[273] Richard H. Shryock and Frans Verdoorn. Notes & correspondence. Isis, 42(4):306–307, December 1951. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226813.

Harvey:1951:QAb

[274] E. Newton Harvey. Queries & answers. Isis, 42(4):307, December 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226814.

Wyckoff:1951:THS

[275] Dorothy Wyckoff, Jane Oppenheimer, and Aaron J. Ihde. Teaching the history of science. Isis, 42(4):308, December 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226815.

Sarton:1951:SSC

[276] George Sarton and Frances Siegel. Seventy-seventh critical bibliography of the history and philosophy of science and of the history of civilization (to March 1951). Isis, 42(4):309–395, December 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226816.

Anonymous:1951:BMd

[277] Anonymous. Back matter. Isis, 42(4):396–397, December 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226817.

Anonymous:1951:VI

[278] Anonymous. Volume information. Isis, 42(4):399–404, December 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226804.

Anonymous:1952:FMa

[279] Anonymous. Front matter. Isis, 43(1):1–2, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227127.

Fleming:1952:LHI

[280] Donald Fleming. Latent heat and the invention of the Watt engine. Isis, 43(1):3–5, April 1952. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227128.

Foote:1952:SHD

[281] George A. Foote. Sir Humphry Davy and his audience at the Royal Institution. Isis, 43(1):6–12, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227129.

Kuhn:1952:RBS

[282] Thomas S. Kuhn. Robert Boyle and structural chemistry in the Seventeenth Century. Isis, 43(1):12–36, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227130.

Clagett:1952:MFB

[283] Marshall Clagett. A Medieval fragment of the De Sphaera et Cylindro of Archimedes. Isis, 43(1):36–38, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227131.

Rudolph:1952:SCC

[284] Richard C. Rudolph. A second-century Chinese illustration of salt mining. Isis, 43(1):39–41, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227132.

Kennedy:1952:FCP

[285] E. S. Kennedy. A Fifteenth-Century planetary computer: al-Kāshī’s “T .abaq al-Manāt.eq”. II. Longitudes, distances, and equations of the planets. Isis, 43(1):42–50, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227133.

Clagett:1952:THS

[286] Marshall Clagett. Teaching the history of science. Isis, 43(1):51–53, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227134.

Anonymous:1952:Pa

[287] Anonymous. Personalia. Isis, 43(1):54, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227136.

Sarton:1952:QAa

[288] George Sarton. Queries & answers. Isis, 43(1):54, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227135.

Anonymous:1952:Oa

[289] Anonymous. Obituaries. Isis, 43(1):55–56, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227137.

Raisz:1952:NC

[290] Erwin Raisz. Notes & correspondence. Isis, 43(1):56–57, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227138.

Sarton:1952:BRBa

[291] George Sarton. Book review: Corpus Platonicum medii aevi by Raymundus Klibansky; Plato Arabus by Richardus Walzer; Volumen I. Galeni compendium Timaei Platonis aliorumque dialogorum synopsis quae extant fragmenta by Paulus Kraus; Richardus Walzer. Isis, 43(1): 57–58, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227139.

Gandz:1952:BRBa

[292] Solomon Gandz. Book review: Koheleth — The Man and His World by Robert Gordis. Isis, 43(1):58, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227140.

Sarton:1952:BRBb

[293] George Sarton. Book review: Histoire naturelle by Pline L’Ancien. Isis, 43(1):58–60, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/227141.

Clagett:1952:BRB

[294] Marshall Clagett. Book review: Ugo Benzi, Medieval Philosopher and Physician, 1376–1439 by Dean Putnam Lockwood. Isis, 43(1):60–62, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227142.

Lafeuille:1952:BRB

[295] G. Lafeuille. Book review: Bibliothèque d’Humanisme et Renaissance. Isis, 43(1):62–64, April 1952. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227143.

Pagel:1952:BRBa

[296] Walter Pagel. Book review: Selected Writings by Paracelsus; Jolande Jacobi; Norbert Guterman. Isis, 43(1):64, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227144.

Hellman:1952:BRB

[297] C. Doris Hellman. Book review: O cometa do ano de 1577 by Francisco Sanches. Isis, 43(1):64–65, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227145.

Sarton:1952:BRBc

[298] George Sarton. Book review: Christopher Polhem and the Mining Company Stora Kopparberg by Sten Lindroth. Isis, 43(1):65–66, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227146.

Bauer:1952:BRB

[299] Hugo Bauer. Book review: Paul Ehrlich by Martha Marquardt. Isis, 43 (1):66, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227148.

Pagel:1952:BRBb

[300] Walter Pagel. Book review: A Seventeenth Century Doctor and His Patients: John Symcotts (1592(?)-1662) by F. N. L. Poynter; W. J.

Bishop. Isis, 43(1):66, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227147.

Sarton:1952:BRBd

[301] George Sarton. Book review: A History of Autobiography in Antiquity by George Misch. Isis, 43(1):67–69, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227149.

Sarton:1952:BRBe

[302] George Sarton and Francis J. Carmody. Book review: The Exact Sciences in Antiquity by Otto Neugebauer. Isis, 43(1):69–73, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227150.

Wohl:1952:BRBa

[303] Sonia S. Wohl. Book review: An Index of Medieval Studies Published in Festschriften, 1865–1946, with Special Reference to Romanic Material by Harry F. Williams. Isis, 43(1):73–74, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227151.

Graubard:1952:BRB

[304] Mark Graubard. Book review: The State of Asia by Lawrence K. Rosinger. Isis, 43(1):74–75, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227152.

Montagu:1952:BRBa

[305] M. F. Ashley Montagu. Book review: The History and Culture of the Indian People. Volume I, The Vedic Age by R. C. Majumdar; A.D. Pusalker. Isis, 43(1):75–76, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227153.

Frye:1952:BRB

[306] Richard N. Frye. Book review: A Medical History of Persia and the Eastern Caliphate by Cyril Elgood. Isis, 43(1):76–77, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227154.

Sarton:1952:BRBf

[307] George Sarton. Book review: L’oeuvre mathématique de G. Desargues by René Taton; L’oeuvre Scientifique de Monge by René Taton. Isis, 43(1): 77–79, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227155.

Farber:1952:BRBa

[308] Eduard Farber. Book review: Geschichte der Ammoniaksynthesc by Alwin Mittasch. Isis, 43(1):79–80, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227156.

Zirkle:1952:BRBa

[309] Conway Zirkle. Book review: The Origin, Variation, Immunity and Breeding of Cultivated Plants by N. I. Vavilov; K. Starr Chester. Isis, 43 (1):80, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227157.

Zirkle:1952:BRBb

[310] Conway Zirkle. Book review: Horses: The Story of the Horse Family in the Modern World through Sixty Million Years of History by George Gaylord Simpson. Isis, 43(1):80–81, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227158.

Erikson:1952:BRB

[311] G. E. Erikson. Book review: The Anatomy of the Gorilla by William King Gregory. Isis, 43(1):81–83, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227159.

Montagu:1952:BRBb

[312] M. F. Ashley Montagu. Book review: Social Anthropology by E. E. Evans-Pritchard. Isis, 43(1):83–84, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227160.

Sarton:1952:BRBg

[313] George Sarton. Book review: Of Societies and Men by Caryl P. Haskins. Isis, 43(1):84–85, April 1952. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227161.

White:1952:BRB

[314] Morton White. Book review: The Structure of Appearance by Nelson Goodman. Isis, 43(1):85–87, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227162.

Lenzen:1952:BRBa

[315] V. F. Lenzen. Book review: Contributions to the Analysis and Synthesis of Knowledge by Philipp Frank. Isis, 43(1):87–88, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227163.

Lenzen:1952:BRBb

[316] V. F. Lenzen. Book review: The New Perspective by Rudolf Jordan. Isis, 43(1):88–89, April 1952. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227164.

Gandz:1952:BRBb

[317] Solomon Gandz. Book review: The Dawn of Philosophy. A Philosophical Primer by Georg Misch; R. F. C. Hull. Isis, 43(1):89, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227165.

Montagu:1952:BRBc

[318] M. F. Ashley Montagu. Book review: Frontal Lobotomy and Affective Behavior by John F. Fulton. Isis, 43(1):89–90, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227166.

Leake:1952:BRBa

[319] C. D. Leake. Book review: The Great Medical Bibliographers: A Study in Humanism by John F. Fulton. Isis, 43(1):90–91, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227167.

Ackerkneciit:1952:BRB

[320] Erwin H. Ackerkneciit. Book review: Yellow Fever in Galveston, Republic of Texas, 1839. An Account of the Great Epidemic together with a Biographical Sketch by Chauncey D. Leake and Stories of the Men Who Conquered Yellow Fever by Ashbel Smith. Isis, 43(1):91, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227168.

Anonymous:1952:BMa

[321] Anonymous. Back matter. Isis, 43(1):92, April 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227169.

Anonymous:1952:FMb

[322] Anonymous. Front matter. Isis, 43(2):93–94, July 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227170.

Boyer:1952:DRR

[323] Carl B. Boyer. Descartes and the radius of the rainbow. Isis, 43(2): 95–98, July 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227171.

Mintz:1952:GHC

[324] Samuel I. Mintz. Galileo, Hobbes, and the circle of perfection. Isis, 43(2): 98–100, July 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227172.

Kibre:1952:LCD

[325] Pearl Kibre. Lewis of Caerleon, doctor of medicine, astronomer, and mathematician (d. 1494?). Isis, 43(2):100–108, July 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227173.

Roller:1952:TFS

[326] Duane H. D. Roller and M. Thilorier. Thilorier and the first solidification of a “permanent” gas (1835). Isis, 43(2):109–113, July 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227174.

Guerriero:1952:NVN

[327] Henry Guerriero. A note on venesection in the New Orleans area. Isis, 43(2):114–116, July 1952. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227175.

Anonymous:1952:DDP

[328] Anonymous. Doctoral dissertations in progress. Isis, 43(2):116–117, July 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227176.

Anonymous:1952:Pb

[329] Anonymous. Personalia. Isis, 43(2):117, July 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227177.

Anonymous:1952:Ob

[330] Anonymous. Obituaries. Isis, 43(2):117–119, July 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227178.

Edelstein:1952:NC

[331] Ludwig Edelstein, Giorgio de Santillana, Walter Pitts, Marie Boas, Thomas S. Kuhn, Herbert Reichner, Louise Patterson, and George Sarton. Notes & correspondence. Isis, 43(2):119–127, July 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227179.

Sarton:1952:SEC

[332] George Sarton and Frances Siegel. Seventy-eighth critical bibliography of the history and philosophy of science and of the history of civilization (to December 1951). Isis, 43(2):128–208, July 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227180.

Anonymous:1952:BMb

[333] Anonymous. Back matter. Isis, 43(2):??, July 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227181.

Anonymous:1952:FMc

[334] Anonymous. Front matter. Isis, 43(3):209–210, September 1952. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227459.

Bluh:1952:VIS

[335] Otto Blüh. The value of inspiration. A study on Julius Robert Mayer and Josef Popper-Lynkeus. Isis, 43(3):211–220, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227460.

Rosenberg:1952:BSS

[336] Albert Rosenberg. Bishop sprat on science and imagery. Isis, 43(3): 220–222, September 1952. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227461.

Mautner:1952:RGC

[337] Franz H. Mautner and Franklin Miller, Jr. Remarks on G. C. Lichtenberg, humanist–scientist. Isis, 43(3):223–231, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227462.

Suter:1952:FGI

[338] Rufus Suter. Four Galileian inscriptions. Isis, 43(3):231–236, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227463.

Clagett:1952:UMT

[339] Marshall Clagett. The use of the Moerbeke translations of Archimedes in the works of Johannes de Muris. Isis, 43(3):236–242, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227464.

Dvoichenko-Markov:1952:POS

[340] Eufrosina Dvoichenko-Markov. The Pulkovo Observatory and some American astronomers of the mid-19th Century. Isis, 43(3):243–246, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227465.

VandeVelde:1952:JPC

[341] A. J. J. Van de Velde. Jean Palfyn, 1650–1730. Chirurgien de Gand. Isis, 43(3):246–250, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227466.

Tasch:1952:CMA

[342] Paul Tasch. Conservation of momentum in antiquity. a note on the prehistory of the principle of jet-propulsion. Isis, 43(3):251–252, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227467.

Thorndike:1952:REY

[343] Lynn Thorndike. A record of eclipses for the years 1478 to 1506. Isis, 43 (3):252–256, September 1952. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227468.

Levey:1952:EAS

[344] Martin Levey. The encyclopedia of Abraham Savasorda: A departure in mathematical methodology. Isis, 43(3):257–264, September 1952. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227469.

Anonymous:1952:Pc

[345] Anonymous. Personalia. Isis, 43(3):264–265, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227470.

Anonymous:1952:Oc

[346] Anonymous. Obituaries. Isis, 43(3):265–266, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227471.

Dick:1952:QA

[347] Hugh G. Dick, John C. Eaton, Edward Fueter, and Paul Schrecker. Queries & answers. Isis, 43(3):266–267, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227472.

Sarton:1952:BRBh

[348] George Sarton. Book review: Macrobius. Commentary on the Dream of Scipio by William Harris Stahl. Isis, 43(3):267–268, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227473.

Sarton:1952:BRBi

[349] George Sarton. Book review: Al-Biruni Commemoration Volume. A.H. 362–A.H. 1362. Isis, 43(3):268–270, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227474.

Pagel:1952:BRBc

[350] Walter Pagel. Book review: Leonardo da Vinci on Movement of the Heart and Blood by K. D. Keele. Isis, 43(3):270–271, September 1952.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227475.

OMalley:1952:BRB

[351] Charles D. O’Malley. Book review: Lazarus Ercker’s Treatise on Ores and Assaying by Anneliese Grunhaldt Sisco; Cyril Stanley Smith. Isis, 43 (3):271–272, September 1952. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227476.

Pagel:1952:BRBd

[352] Walter Pagel. Book review: Paracelsus. Sozialethische und Sozialpolitische Schriften by Kurt Goldammer. Isis, 43(3):272, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227477.

Galdston:1952:BRB

[353] Iago Galdston. Book review: The Maggid of Caro by Hirsch Loeb Gordon. Isis, 43(3):272–274, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227478.

Leake:1952:BRBb

[354] Chauncey D. Leake. Book review: Bartolomé de Las Casas: Bookman, Scholar and Propagandist by Lewis Hanke. Isis, 43(3):274–275, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227479.

Schneider:1952:BRB

[355] Heinrich Schneider. Book review: Leibniz by Kurt Huber. Isis, 43(3): 275–276, September 1952. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227480.

Cohen:1952:BRBa

[356] I. Bernard Cohen. Book review: John Ray, a Bibliography by Geoffrey Keynes; A Bio-Bibliography of Edward Jenner, 1749–1823 by W. R. LeFanu. Isis, 43(3):276–277, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227481.

Sarton:1952:BRBj

[357] George Sarton. Book review: Mémoires autobiographiques de Charles Bonnet de Genève by Charles Bonnet; Raymond Savioz; La philosophie de Charles Bonnet de Genève by Raymond Savioz. Isis, 43(3):277–280, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227482.

Brinton:1952:BRB

[358] Crane Brinton. Book review: The Philosophy of the Enlightenment by Ernst Cassirer; Fritz C. A. Koelln; James P. Pettegrove. Isis, 43(3): 280–281, September 1952. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227483.

Hindle:1952:BRB

[359] Brooke Hindle. Book review: Thomas Jefferson: Scientist by Edwin T. Martin; The Papers of Thomas Jefferson. Vol. I (1760–1776) by Julian P. Boyd; Lyman H. Butterfield; Mina R. Bryan. Isis, 43(3):281–282, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227484.

Miles:1952:BRB

[360] Wyndham Miles. Book review: Herbert H. Dow; Pioneer in Creative Chemistry by Murray Campbell; Harrison Hatton. Isis, 43(3):282–283, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227485.

Farber:1952:BRBb

[361] Eduard Farber. Book review: A Romance in Research. The Life of Charles F. Burgess: Student — Teacher — Researcher – Industrialist by Alexander McQueen. Isis, 43(3):283, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227486.

Smith:1952:BRB

[362] Cyril Stanley Smith. Book review: Metallurgy in Antiquity. A Notebook for Archaeologists and Technologists by R. J. Forbes. Isis, 43(3):283–285, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227487.

Sarton:1952:BRBk

[363] George Sarton. Book review: Corinth. Results of Excavations Conducted by the American School of Classical Studies at Athens. Volume XIV. by Carl Roebuck; F. J. de Waele. Isis, 43(3):285–286, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227488.

Sarton:1952:BRBl

[364] George Sarton. Book review: The Alexandrian Library, Glory of the Hellenic World. Its Rise, Antiquities, and Destructions by Edward Alexander Parsons. Isis, 43(3):286–287, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227489.

Gandz:1952:BRBc

[365] Solomon Gandz. Book review: The First Book of Maccabees by Solomon Zeitlin; Abraham A. Neuman; Sidney Tedesche; Aristeas to Philocrates (Letter of Aristeas) by Moses Hadas. Isis, 43(3):287–288, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227490.

Montagu:1952:BRBd

[366] Ashley Montagu. Book review: Folk-Songs of Chhattisgarh by Verrier Elwin; The Muria and Their Ghotul by Verrier Elwin; Maria Murder and Suicide by Verrier Elwin; Bondo Highlander by Verrier Elwin; The Tribal Art of Middle India by Verrier Elwin. Isis, 43(3):288–289, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227491.

Cohen:1952:BRBb

[367] I. Bernard Cohen. Book review: Osiris, Vol. 9 by George Sarton. Isis, 43 (3):289–290, September 1952. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227492.

Struik:1952:BRB

[368] D. J. Struik. Book review: Geschichte der Mathematik by O. Becker; J. E. Hofmann. Isis, 43(3):290–291, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227493.

Abrams:1952:BRB

[369] J. W. Abrams. Book review: Astronomie. Geschichte ihrer Probleme by Ernest Zinner. Isis, 43(3):291–292, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227494.

Lenzen:1952:BRBc

[370] V. F. Lenzen. Book review: The Rise of the New Physics: Its Mathematical and Physical Theories by A. D’Abro. Isis, 43(3):292–293, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227495.

Lenzen:1952:BRBd

[371] V. F. Lenzen. Book review: A History of the Theories of Aether and Electricity by Edmund Whittaker. Isis, 43(3):293–294, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227496.

Zirkle:1952:BRBc

[372] Conway Zirkle. Book review: Human Fertility: The Modern Dilemma by Robert C. Cook. Isis, 43(3):294–295, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227497.

Wohl:1952:BRBb

[373] Sonia S. Wohl. Book review: The Art of Botanical Illustration by Wilfrid Blunt; William T. Stearn; National Book League, London: Flower Books and Their Illustrators: Catalogue of an Exhibition Arranged for the National Book League by Wilfrid Blunt; Die botanische Illustration; ihre Geschichte und Bibliographie by Claus Nissen. Isis, 43(3):295–296, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227498.

Zirkle:1952:BRBd

[374] Conway Zirkle. Book review: The Old Company’s Garden at the Cape and Its Superintendents by Mia C. Karsten. Isis, 43(3):296–297, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227499.

Zirkle:1952:BRBe

[375] Conway Zirkle. Book review: Die Evolutionstheorie, das Grundproblem der modernen Biologie by Georg Schneider. Isis, 43(3):297–298, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227500.

Cottrell:1952:BRBa

[376] G. W. Cottrell, Jr. Book review: The History of American Ornithology before Audubon by Elsa G. Allen. Isis, 43(3):298, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227501.

Cottrell:1952:BRBb

[377] G. W. Cottrell, Jr. Book review: Die Entwicklung der Ornithologie von Aristoteles bis zur Gegenwart by Erwin Stresemann. Isis, 43(3):298–300, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227502.

Ihde:1952:BRBa

[378] Aaron J. Ihde. Book review: General Education in Science by I. Bernard Cohen; Fletcher G. Watson. Isis, 43(3):300, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227503.

Ihde:1952:BRBb

[379] Aaron J. Ihde. Book review: Crisis and Hope in American Education by Robert Ulich. Isis, 43(3):301, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227504.

Wohl:1952:BRBc

[380] Sonia S. Wohl. Book review: The Lost Language of Symbolism; An Inquiry into the Origin of Certain Letters, Words, Names, Fairy-Tales, Folklore, and Mythologies by Harold Bayley. Isis, 43(3):301–302, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227505.

Wohl:1952:BRBd

[381] Sonia S. Wohl. Book review: The Bibliographical History of Anonyma and Pseudonyma by Archer Taylor; Frederic J. Mosher. Isis, 43(3):302– 303, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/227506.

Wohl:1952:BRBe

[382] Sonia S. Wohl. Book review: Printing Types, Their History, Forms, and Uses; A Study in Survivals by Daniel Berkeley Updike. Isis, 43(3): 303–304, September 1952. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227507.

Anonymous:1952:BMc

[383] Anonymous. Back matter. Isis, 43(3):??, September 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227508.

Anonymous:1952:FMd

[384] Anonymous. Front matter. Isis, 43(4):305–306, December 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227382.

Brown:1952:OLP

[385] Harcourt Brown. An odd lot. Presidential address delivered at the close of the meeting of the History of Science Society at Brown university, 5 April 1952. Isis, 43(4):307–311, December 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227383.

Koyre:1952:ULR

[386] Alexandre Koyré. An unpublished letter of Robert Hooke to Isaac Newton. Isis, 43(4):312–337, December 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227384.

Sisco:1952:HOB

[387] Anneliese Sisco. How old is the Bergbüchlein? Isis, 43(4):337–343, December 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227385.

Rosen:1952:GDB

[388] Edward Rosen. Galileo on the distance between the Earth and the Moon. Isis, 43(4):344–348, December 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227386.

Stahlman:1952:DCA

[389] William D. Stahlman. On the date of a comet ascribed to A.D. 1238. Isis, 43(4):348–351, December 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227387.

Hujer:1952:FPD

[390] Karel Hujer. Father Procopius Diviš — the European Franklin. Isis, 43 (4):351–357, December 1952. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227388.

Cohen:1952:DDE

[391] I. Bernard Cohen and Robert Schofield. Did Diviš erect the first European protective lightning rod, and was his invention independent? Isis, 43(4):358–364, December 1952. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227389.

S:1952:NC

[392] G. S., Thomas S. Kuhn, and C. W. Adams. Notes & correspondence. Isis, 43(4):364–366, December 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227390.

Sarton:1952:QAb

[393] George Sarton. Queries & answers. Isis, 43(4):366, December 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227391.

Anonymous:1952:Od

[394] Anonymous. Obituaries. Isis, 43(4):366–368, December 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227392.

Leake:1952:THS

[395] Chauncey D. Leake. Teaching the history of science. Isis, 43(4):368, December 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227393.

Anonymous:1952:DPH

[396] Anonymous. Dissertations in progress in the history of science. Isis, 43 (4):368–369, December 1952. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227394.

Cohen:1952:BRBc

[397] I. Bernard Cohen. Book review: Memoirs by Benjamin Franklin; Max Farrand; Autobiography by Benjamin Franklin; Max Farrand. Isis, 43 (4):369–371, December 1952. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227395.

Sutton:1952:BRB

[398] Robert F. Sutton. Book review: Russia’s Lomonosov, Chemist, Courtier, Physicist, Poet by Boris N. Menshutkin; Jeannette Eyre Thal; Edward J. Webster; W. Chapin Huntington. Isis, 43(4):371–373, December 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227396.

Cohen:1952:BRBd

[399] I. Bernard Cohen. Book review: The Works of George Berkeley, Bishop of Cloyne by A. A. Luce; T. E. Jessop. Isis, 43(4):373–374, December 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227397.

Leake:1952:BRBc

[400] Chauncey D. Leake. Book review: Claude Bernard and the Experimental Method in Medicine by J. M. D. Olmsted; E. Harris Olmsted. Isis, 43 (4):374, December 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/227398.

Pagel:1952:BRBe

[401] Walter Pagel. Book review: Ethan Allen Hitchcock. Soldier — Humanitarian — Scholar. Discoverer of the “True Subject” of the Hermetic Art by I. Bernard Cohen. Isis, 43(4):374–375, December 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227399.

Temkin:1952:BRB

[402] Owsei Temkin. Book review: The Greeks and the Irrational by E. R. Dodds. Isis, 43(4):375–377, December 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227400.

Cohen:1952:BRBe

[403] I. Bernard Cohen. Book review: A Century of Science, 1851–1951 by Herbert Dingle; A Century of Technology, 1851–1951 by Percy Dunsheath. Isis, 43(4):377–378, December 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227401.

Ihde:1952:BRBc

[404] Aaron J. Ihde. Book review: The Chemical Revolution. A Contribution to Social Technology by Archibald Clow; Nan L. Clow. Isis, 43(4):378–379, December 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227402.

Zirkle:1952:BRBf

[405] Conway Zirkle. Book review: Soviet Genetics by Alan G. Morton. Isis, 43 (4):379–380, December 1952. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227403.

Zirkle:1952:BRBg

[406] Conway Zirkle. Book review: Plants and the Atmosphere by Leonard K. Nash. Isis, 43(4):380–381, December 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227404.

Meyer:1952:BRB

[407] Heinrich Meyer. Book review: Die Einheit unseres Wirklichkeitsbildes und die Grenzen der Einzelwissenschaften by E. Grassi; T. v. Uexküll. Isis, 43(4):381–382, December 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227405.

Condit:1952:BRB

[408] Carl W. Condit. Book review: History of the Development of Building Construction in Chicago by Frank A. Randall; American Building. The Forces That Shape It by James Marston Fitch; Contemporary Structure in Architecture by Leonard Michaels. Isis, 43(4):382–383, December 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227406.

Fleming:1952:BRB

[409] Donald Fleming. Book review: The Counter-Revolution of Science. Studies on the Abuse of Reason by F. A. Hayek. Isis, 43(4):383–385, December 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227407.

Urdang:1952:BRBa

[410] George Urdang. Book review: Geschichte der Medizin. Die historische Entwicklung der Heilkunde und des ärztlichen Lebens. II. Bd. I. Hälfte: Von der Medizin der Aufklärung bis zur Begründung der Zellularpathologie (ca. 1740–ca. 1858) by Paul Diepgen. Isis, 43(4):385–386, December 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227408.

Urdang:1952:BRBb

[411] George Urdang. Book review: Le prestigieux passé des eaux minérales. Histoire du thermalisme et de l’hydrologie des origines à 1950. by E.- H. Guitard. Isis, 43(4):386, December 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227409.

Boas:1952:AD

[412] Marie Boas, I. Bernard Cohen, and Frederick G. Kilgour. Administrative documents. Isis, 43(4):387–393, December 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227410.

Anonymous:1952:VI

[413] Anonymous. Volume information. Isis, 43(4):394–400, December 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227381.

Anonymous:1952:BMd

[414] Anonymous. Back matter. Isis, 43(4):401–402, December 1952. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227411.

Anonymous:1953:FMa

[415] Anonymous. Front matter. Isis, 44(1–2):1–2, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227634.

Stimson:1953:EB

[416] Dorothy Stimson. The editing of Isis. Isis, 44(1–2):3–4, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227635.

Rosen:1953:CDI

[417] Edward Rosen. Carlo Dati on the invention of eyeglasses. Isis, 44(1–2): 4–10, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227636.

Roller:1953:DBK

[418] Duane H. D. Roller. Did Bacon know Gilbert’s De Magnete? Isis, 44 (1–2):10–13, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/227637.

Fester:1953:EFS

[419] G. A. Fester. Einige Farbstoffe südamerikanischer Kulturvolker. (German) [Some dyes of the South American Volker culture]. Isis, 44(1–2): 13–16, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227638.

Clagett:1953:MLT

[420] Marshall Clagett. The Medieval Latin translations from the Arabic of the Elements of Euclid, with special emphasis on the versions of Adelard of Bath. Isis, 44(1–2):16–42, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227639.

DeGrazia:1953:MDE

[421] Alfred De Grazia. Mathematical derivation of an election system. Isis, 44(1–2):42–51, June 1953. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227640.

Patterson:1953:RSS

[422] Louise Diehl Patterson. The Royal Society’s standard thermometer, 1663–1709. Isis, 44(1–2):51–64, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227641.

MacClintock:1953:BRB

[423] Stuart MacClintock. Book review: Le Quaestiones de Anima di Taddeo da Parma by Sofia Vanni Rovighi. Isis, 44(1–2):64–65, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227642.

Sarton:1953:BRBa

[424] George Sarton. Book review: Leonardo da Vinci on the Human Body by Charles D. O’Malley; J. B. de C. M. Saunders. Isis, 44(1–2):65–66, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227643.

Hujer:1953:BRB

[425] Karel Hujer. Book review: Two Pansophical Works by John Amos Comenius; G. H. Turnbull. Isis, 44(1–2):66–68, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227644.

Lenzen:1953:BRB

[426] V. F. Lenzen. Book review: Opticks by Isaac Newton. Isis, 44(1–2): 68–69, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227645.

Hindle:1953:BRB

[427] Brooke Hindle. Book review: Letters of Benjamin Rush by L. H. Butterfield. Isis, 44(1–2):69–70, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227646.

Smith:1953:BRB

[428] Edwin Burrows Smith. Book review: Jean-Sylvain Bailly, Revolutionary Mayor of Paris by Gene A. Brucker. Isis, 44(1–2):70–72, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227647.

Guerlac:1953:BRB

[429] Henry Guerlac. Book review: Inventaire du Fonds Vandeul et inédits de Diderot by Herbert Dieckmann. Isis, 44(1–2):72–74, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227648.

Clagett:1953:BRBa

[430] Marshall Clagett. Book review: Lightning Protection — 200th Anniversary of the “Philadelphia Experiment” by Benjamin Franklin. Isis, 44 (1–2):74, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/227649.

Sarton:1953:BRBb

[431] George Sarton. Book review: The Tradition of the Minor Greek Geographers by Aubrey Diller. Isis, 44(1–2):75, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227650.

Drabkin:1953:BRB

[432] I. E. Drabkin. Book review: A History of Science. Ancient Science through the Golden Age of Greece by George Sarton. Isis, 44(1–2):75– 78, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227651.

Sarton:1953:BRBc

[433] George Sarton. Book review: A History of Muslim Historiography by Franz Rosenthal. Isis, 44(1–2):78–79, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227652.

Zirkle:1953:BRBa

[434] Conway Zirkle. Book review: Apes and Ape Lore in the Middle Ages and the Renaissance by H. W. Janson. Isis, 44(1–2):79–80, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227653.

Zirkle:1953:BRBb

[435] Conway Zirkle. Book review: Wild Men in the Middle Ages: A Study in Art, Sentiment and Demonology by Richard Bernheimer. Isis, 44(1– 2):80, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227654.

Zirkle:1953:BRBc

[436] Conway Zirkle. Book review: The Origin of Life and the Evolution of Living Things. An Environmental Theory by Olan R. Hyndman. Isis, 44 (1–2):81, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227655.

Zirkle:1953:BRBd

[437] Conway Zirkle. Book review: Plants, Man and Life by Edgar Anderson. Isis, 44(1–2):81–82, June 1953. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227656.

Zirkle:1953:BRBe

[438] Conway Zirkle. Book review: Flour for Man’s Bread by John Storck; Walter Dorwin Teague. Isis, 44(1–2):82–83, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227657.

Ihde:1953:BRBa

[439] Aaron J. Ihde. Book review: The Evolution of Chemistry by Eduard Farber. Isis, 44(1–2):83, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227658.

Ihde:1953:BRBb

[440] Aaron J. Ihde. Book review: A Source Book in Chemistry 1400–1900 by Henry M. Leicester; Herbert S. Klickstein. Isis, 44(1–2):84, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227659.

Graubard:1953:BRBa

[441] Mark Graubard. Book review: The Uses of the Past by Herbert J. Muller. Isis, 44(1–2):84–86, June 1953. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227660.

Brozek:1953:BRB

[442] Josef Brožek. Book review: A History of Psychology in Autobiography. Vol. 4 by Edwin G. Boring; Herbert S. Langfeld; Heinz Werner; Robert Yerkes. Isis, 44(1–2):86–87, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227661.

Beebe-Center:1953:BRB

[443] J. G. Beebe-Center. Book review: La psychologie contemporaine by Paul Foulquié; Gérard Deledalle. Isis, 44(1–2):87, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227662.

Graubard:1953:BRBb

[444] Mark Graubard. Book review: Satan by Bruno De Jesus-Marie. Isis, 44 (1–2):88, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227663.

Graubard:1953:BRBc

[445] Mark Graubard. Book review: Man and His Gods by Homer W. Smith. Isis, 44(1–2):88–89, June 1953. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227664.

Jammer:1953:BRB

[446] Max Jammer. Book review: M’koroth l’tholdoth habiologia by F. S. Bodenheimer. Isis, 44(1–2):89–91, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227665.

Clagett:1953:BRBb

[447] Marshall Clagett. Book review: Horus — A Guide to the History of Science by George Sarton. Isis, 44(1–2):91–93, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227666.

Graubard:1953:BRBd

[448] Mark Graubard. Book review: The Accused by Alexander Weissberg. Isis, 44(1–2):93–94, June 1953. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227667.

Anonymous:1953:Pa

[449] Anonymous. Personalia. Isis, 44(1–2):95, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227668.

Condit:1953:THS

[450] Carl W. Condit. Teaching the history of science. Isis, 44(1–2):95–96, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227669.

Pelseneer:1953:NC

[451] J. Pelseneer, Denis Duveen, Lynn Thorndike, E. J. Dijksterhuis, R. J. Forbes, and Aaron J. Ihde. Notes & correspondence. Isis, 44(1–2):96– 99, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227670.

Sarton:1953:QA

[452] George Sarton and C. W. Adams. Queries & answers. Isis, 44(1–2): 99–101, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227671.

Sarton:1953:SNC

[453] George Sarton and Frances Siegel. Seventy-ninth critical bibliography of the history and philosophy of science and of the history of civilization (to 31 August 1952). Isis, 44(1–2):102–204, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227672.

Kilgour:1953:AD

[454] Frederick G. Kilgour, Charles C. Gillespie, and Marie Boas. Administrative documents. Isis, 44(1–2):205–210, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227673.

Anonymous:1953:BMa

[455] Anonymous. Back matter. Isis, 44(1–2):??, June 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227674.

Anonymous:1953:FMb

[456] Anonymous. Front matter. Isis, 44(3):211–212, September 1953. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227085.

Temkin:1953:GMS

[457] Owsei Temkin. Greek medicine as science and craft. Isis, 44(3):213–225, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227086.

Guerlac:1953:PRB

[458] Henry Guerlac. A proposed revision of the Isis critical bibliography. Isis, 44(3):226–228, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227087.

Anonymous:1953:PSC

[459] Anonymous. Proposed system of classification for Isis critical bibliography. Isis, 44(3):229–231, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227088.

Sarton:1953:WB

[460] George Sarton. Why Isis? Isis, 44(3):232–242, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227089.

Dupree:1953:JWV

[461] A. Hunter Dupree. Jeffries Wyman’s views on evolution. Isis, 44(3): 243–246, September 1953. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227090.

Pines:1953:PBT

[462] S. Pines. Un précurseur bagdadien de la théorie de l’impetus. Isis, 44 (3):247–251, September 1953. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227091.

Fisch:1953:SPN

[463] Harold Fisch. The scientist as priest: A note on Robert Boyle’s natural theology. Isis, 44(3):252–265, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227092.

Schneider:1953:ULC

[464] Heinrich Schneider. An unpublished letter concerning Leibniz. Isis, 44 (3):266–272, September 1953. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227093.

Norlind:1953:CLC

[465] Wilhelm Norlind. Copernicus and luther: A critical study. Isis, 44(3): 273–276, September 1953. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227094.

Anonymous:1953:Pb

[466] Anonymous. Personalia. Isis, 44(3):276, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227095.

Stahlman:1953:NC

[467] W. D. Stahlman and L. Carrington Goodrich. Notes & correspondence. Isis, 44(3):277, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227096.

Stahlman:1953:BRBa

[468] William D. Stahlman. Book review: The Almagest, by Ptolemy by Hutchins, Robert; R. Catesby Taliaferro. Isis, 44(3):278–280, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227097.

Clagett:1953:BRBc

[469] Marshall Clagett. Book review: Al-Bitruji de motibus celorum. Critical Edition of the Latin Translation of Michael Scot by Francis J. Carmody. Isis, 44(3):280–281, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227098.

Saunders:1953:BRB

[470] J. B. deC. M. Saunders. Book review: Vesalius on the Human Brain by Charles Singer. Isis, 44(3):281–283, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227099.

Pagel:1953:BRBa

[471] Walter Pagel. Book review: Volcher Coiter, 1534–1576 by Herrlinger, Robert. Isis, 44(3):283, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227100.

Kuhn:1953:BRBa

[472] Thomas S. Kuhn. Book review: Ballistics in the Seventeenth Century, a Study in the Relations of Science and War with Reference Principally to England by A. R. Hall. Isis, 44(3):284–285, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227101.

Kuhn:1953:BRBb

[473] Thomas S. Kuhn. Book review: The Scientific Work of René Descartes (1596–1650) by J. F. Scott; Descartes and the Modern Mind by Albert G. A. Balz. Isis, 44(3):285–287, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227102.

Cohen:1953:BRBa

[474] I. Bernard Cohen. Book review: Philosophiae naturalis principia mathematica by Isaac Newton; A New System of Chemical Philosophy by John Dalton. Isis, 44(3):287–288, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227103.

Zirkle:1953:BRBf

[475] Conway Zirkle. Book review: Carl Linnoeus by Knut Hagberg; Alan Blair. Isis, 44(3):288–289, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227104.

Farber:1953:BRBa

[476] Eduard Farber. Book review: Friedrich Nietzsche als Naturphilosoph by Alwin Mittasch. Isis, 44(3):289–290, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227105.

Fleming:1953:BRBa

[477] Donald Fleming. Book review: The World of Eli Whitney by Jeannette Mirsky; Allan Nevins. Isis, 44(3):290–291, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227106.

Clark:1953:BRB

[478] Joseph T. Clark. Book review: Stoic Logic by Benson Mates. Isis, 44 (3):291–293, September 1953. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227107.

Cohen:1953:BRBb

[479] I. Bernard Cohen. Book review: Science in Western Civilization: A Syllabus by Henry Guerlac. Isis, 44(3):293–295, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227108.

Sarton:1953:BRBd

[480] George Sarton. Book review: Swedish Men of Science, 1650–1950 by Sten Lindroth. Isis, 44(3):295–296, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227109.

Zirkle:1953:BRBg

[481] Conway Zirkle. Book review: Science and the Social Order by Bernard Barber. Isis, 44(3):296–298, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227110.

Zirkle:1953:BRBh

[482] Conway Zirkle. Book review: In the Name of Science by Martin Gardner. Isis, 44(3):298, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227111.

Hellman:1953:BRB

[483] C. Doris Hellman. Book review: The History of Astronomy by Giorgio Abetti; Betty Burr Abetti. Isis, 44(3):298–300, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227112.

Zirkle:1953:BRBi

[484] Conway Zirkle. Book review: The Mind and the Eye: A Study of the Biologist’s Standpoint by Agnes Arber. Isis, 44(3):300–301, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227113.

Ackerknecht:1953:BRB

[485] Erwin H. Ackerknecht. Book review: A Hundred Years of Biology by Ben Dawes. Isis, 44(3):301, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227114.

Morison:1953:BRB

[486] S. E. Morison. Book review: A History of the Practice of Navigation by J. B. Hewson; A Nautica dos Descobrimentos. Os Descobrimentos maritimos vistos por un navegador. Colectanea de artigos, conferencias e trabalhos ineditos by Almirante Gago Coutinho. Isis, 44(3):301–302, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227115.

Roller:1953:BRB

[487] Duane H. D. Roller. Book review: From Lodestone to Gyro-Compass by H. L. Hutchins; W. E. May. Isis, 44(3):303, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227116.

Anonymous:1953:AD

[488] Anonymous. Administrative documents. Isis, 44(3):304–306, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227117.

Anonymous:1953:BMb

[489] Anonymous. Back matter. Isis, 44(3):??, September 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227118.

Stauffer:1953:PER

[490] R. C. Stauffer. Persistent errors regarding Oersted’s discovery of electromagnetism. Isis, 44(4):307–310, December 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226995.

Williams:1953:SEF

[491] L. Pearce Williams. Science, education and the French revolution. Isis, 44 (4):311–330, December 1953. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226996.

Schofield:1953:JWS

[492] Robert E. Schofield. John Wesley and science in 18th Century England. Isis, 44(4):331–340, December 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226997.

Merlan:1953:PM

[493] Philip Merlan. Plotinus and magic. Isis, 44(4):341–348, December 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226998.

Lipski:1953:FRA

[494] Alexander Lipski. The foundation of the Russian Academy of Sciences. Isis, 44(4):349–354, December 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226999.

Wilson:1953:PCC

[495] Curtis Wilson. Pomponazzi’s criticism of calculator. Isis, 44(4):355–362, December 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227000.

Lurie:1953:SMR

[496] Edward Lurie. Some manuscript resources in the history of Nineteenth Century American natural science. Isis, 44(4):363–370, December 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227001.

Clagett:1953:MMP

[497] Marshall Clagett. Medieval mathematics and physics: A check list of microfilm reproductions. Isis, 44(4):371–381, December 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227002.

Fester:1953:NC

[498] G. A. Fester, J. Pelseneer, and L. J. M. Coleby. Notes & correspondence. Isis, 44(4):382, December 1953. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227003.

Boyer:1953:BRB

[499] Carl B. Boyer. Book review: The Works of Archimedes by T. L. Heath. Isis, 44(4):383–384, December 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227004.

Raisz:1953:BRB

[500] Erwin Raisz. Book review: The Vienna-Klosterneuburg Map Corpus of the Fifteenth Century by Dana Bennett Durand. Isis, 44(4):384–385, December 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227005.

Johnson:1953:BRB

[501] Francis R. Johnson. Book review: Cardano: The Gambling Scholar by Øystein Øre. Isis, 44(4):385–387, December 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227006.

Boas:1953:BRB

[502] Marie Boas. Book review: Locke’s Travels in France 1675–1679 by John Lough. Isis, 44(4):387, December 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227007.

Vartanian:1953:BRB

[503] Aram Vartanian. Book review: Abraham Trembley of Geneva: Scientist and Philosopher by John R. Baker. Isis, 44(4):387–389, December 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227008.

Gillispie:1953:BRB

[504] Charles C. Gillispie. Book review: Diderot and Descartes, a Study of Scientific Naturalism in the Enlightenment by Aram Vartanian. Isis, 44 (4):389–391, December 1953. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227009.

Sarton:1953:BRBe

[505] George Sarton. Book review: Les instruments scientifiques aux XVIIe et XVIIIe siècles by Maurice Daumas. Isis, 44(4):391–392, December 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227010.

Pagel:1953:BRBb

[506] Walter Pagel. Book review: De Aneurysmatibus. Aneurysms. The Latin text of Rome, 1745 by Lancisi,Giovanni Maria; Wilmer Cave Wright. Isis, 44(4):392–393, December 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227011.

Fleming:1953:BRBb

[507] Donald Fleming. Book review: A Bibliography of Oliver Wendell Holmes by Thomas Franklin Currier; Eleanor M. Tilton. Isis, 44(4):393–394, December 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227012.

Lurie:1953:BRBa

[508] Edward Lurie. Book review: David Starr Jordan: Prophet of Freedom by Edward McNall Burns. Isis, 44(4):394, December 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227013.

Zirkle:1953:BRBj

[509] Conway Zirkle. Book review: The Mighty Leaf: Tobacco through the Centuries by Jerome E. Brooks. Isis, 44(4):394–395, December 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227014.

Graubard:1953:BRBe

[510] Mark Graubard. Book review: Man’s Foods by L. B. Jensen. Isis, 44 (4):395–396, December 1953. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227015.

Stahlman:1953:BRBb

[511] William D. Stahlman. Book review: A History of Astronomy from Thales to Kepler. Formerly Titled History of the Planetary Systems from Thales to Kepler by J. L. E. Dreyer. Isis, 44(4):396–397, December 1953. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227016.

Pfeiffer:1953:BRB

[512] Robert H. Pfeiffer. Book review: The Triumph of the Alphabet: A History of Writing by A. C. Moorhouse. Isis, 44(4):397–398, December 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227017.

Clagett:1953:BRBd

[513] Marshall Clagett. Book review: Augustine to Galileo. The History of Science A.D. 400–1650 by A. C. Crombie. Isis, 44(4):398–403, December 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227018.

Lurie:1953:BRBb

[514] Edward Lurie. Book review: Origins of American Scientists: A Study Made under the Direction of a Committee of the Faculty of Wesleyan University by R. H. Knapp; H. B. Goodrich. Isis, 44(4):403–404, December 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227019.

Farber:1953:BRBb

[515] Eduard Farber. Book review: Beiträge zur Geschichte der Naturwissenschaften und der Technik by Edmund O. Von Lippmann. Isis, 44(4): 404, December 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/227020.

Cohen:1953:BRBc

[516] I. Bernard Cohen. Book review: Studies in Intellectual History. Isis, 44 (4):404–405, December 1953. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227021.

Leake:1953:BRB

[517] Chauncey D. Leake. Book review: Salud y Libertad: Oraciones de un Creyente by C. E. Paz-Soldan. Isis, 44(4):405, December 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227022.

Anonymous:1953:VI

[518] Anonymous. Volume information. Isis, 44(4):406–409, December 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226993.

Anonymous:1953:BMc

[519] Anonymous. Back matter. Isis, 44(4):??, December 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227023.

Anonymous:1953:FMc

[520] Anonymous. Front matter. Isis, 44(4):??, December 1953. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226994.

Anonymous:1954:FMa

[521] Anonymous. Front matter. Isis, 45(1):1–2, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227676.

Farber:1954:CVB

[522] Eduard Farber. The color of venous blood. Isis, 45(1):3–9, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227677.

Webb:1954:SGE

[523] Henry J. Webb. The science of gunnery in Elizabethan England. Isis, 45(1):10–21, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/227678.

Thorndike:1954:AS

[524] Lynn Thorndike. Albumasar in sadan. Isis, 45(1):22–32, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227679.

Chipman:1954:ULS

[525] R. A. Chipman. An unpublished letter of Stephen Gray on electrical experiments, 1707–1708. Isis, 45(1):33–40, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227680.

Cohen:1954:NSL

[526] I. Bernard Cohen. Neglected sources for the life of Stephen Gray (1666 or 1667–1736). Isis, 45(1):41–50, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227681.

Guerlac:1954:LHB

[527] Henry Guerlac. Lavoisier and his biographers. Isis, 45(1):51–62, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227682.

Sarton:1954:JC

[528] George Sarton. The Jerusalem Congress. Isis, 45(1):63–77, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227683.

Shu-hua:1954:OB

[529] Li Shu-hua. Origine de la boussole. Isis, 45(1):78–94, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227684.

Coutant:1954:TF

[530] Victor Coutant. Theophrastus and the firewalk. Isis, 45(1):95–97, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227685.

Thorndike:1954:NC

[531] Lynn Thorndike, W. Norlind, and Heinrich Meyer. Notes & correspondence. Isis, 45(1):98–99, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227686.

Kuhn:1954:BRB

[532] Thomas S. Kuhn. Book review: Main Currents of Western Thought by Franklin L. Baumer. Isis, 45(1):100, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227687.

Vetter:1954:BRB

[533] Quido Vetter. Book review: Astronomy in Czechoslovakia from the Earliest Times to Today by H. Slouka. Isis, 45(1):100–101, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227688.

Ihde:1954:BRBa

[534] Aaron J. Ihde. Book review: A History of the American Chemical Society — Seventy-Five Eventful Years by Charles Albert Browne; Mary Elvira Weeks. Isis, 45(1):101–102, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227689.

Wohl:1954:BRB

[535] Sonia S. Wohl. Book review: The Mapmaker’s Art: Essays on the History of Maps by Edward William O’Flaherty Lynam. Isis, 45(1):102–103, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227690.

Zirkle:1954:BRBa

[536] Conway Zirkle. Book review: Historical Aspects of Organic Evolution by Philip G. Fothergill. Isis, 45(1):103–105, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227691.

Pagel:1954:BRBa

[537] Walter Pagel. Book review: Geschichte der Physiologie by Rothschuh, K. E. Isis, 45(1):105–106, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227692.

Montagu:1954:BRB

[538] Ashley Montagu. Book review: Franz Boas by Melville J. Herskovits. Isis, 45(1):106–107, May 1954. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227693.

Kluckhohn:1954:BRB

[539] Clyde Kluckhohn. Book review: The Savages of America. A Study of the Indian and the Idea of Civilization by Roy H. Pearce; Theoretical Anthropology by David Bidney; The World of Primitive Man by Paul Radin. Isis, 45(1):107–109, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227694.

Pagel:1954:BRBb

[540] Walter Pagel. Book review: Index zur Geschichte der Medizin, Naturwissenschaft und Technik, Vol. I by Walter Artelt. Isis, 45(1):109, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227695.

Blake:1954:BRB

[541] John B. Blake. Book review: The Epidemiology of Health by Iago Galdston. Isis, 45(1):110, May 1954. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227696.

Tolstoi:1954:BRB

[542] Edward Tolstoi. Book review: The History of Diabetes Mellitus by N. S. Papaspyres. Isis, 45(1):110–111, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227697.

Richeson:1954:BRB

[543] A. W. Richeson. Book review: Historical Metrology by A. E. Berriman. Isis, 45(1):111–112, May 1954. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227698.

Sarton:1954:BRBa

[544] George Sarton. Book review: Corpus Commentariorum Averrois in Aristotelem. Versionum Latinarum volumen VI, I. Commentarium magnum in Aristotelis de Anima libros by F. Stuart Crawford. Isis, 45(1):112– 113, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227699.

Sarton:1954:BRBb

[545] George Sarton. Book review: The Gate of Heaven (Shaar ha-Shamayim) by Rabbi Gerson Ben Shlomon d’Arles; F. S. Bodenheimer. Isis, 45(1): 113–115, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227700.

Kibre:1954:BRB

[546] Pearl Kibre. Book review: The Domesday Geography of Eastern England by H. C. Darby. Isis, 45(1):115–117, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227701.

Schneider:1954:BRB

[547] Heinrich Schneider. Book review: Goethe the Alchemist by Ronald D. Gray. Isis, 45(1):117–119, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227702.

Fleming:1954:BRBa

[548] Donald Fleming. Book review: Benjamin Franklin by I. Bernard Cohen. Isis, 45(1):119–120, May 1954. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227703.

Sigerist:1954:BRBa

[549] Henry E. Sigerist. Book review: Rudolf Virchow, Doctor, Statesman, Anthropologist by Erwin H. Ackerknecht. Isis, 45(1):121–122, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227704.

Boas:1954:AD

[550] Marie Boas and Charles C. Gillispie. Administrative documents. Isis, 45(1):123–126, May 1954. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227705.

Anonymous:1954:BMa

[551] Anonymous. Back matter. Isis, 45(1):127, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227706.

Anonymous:1954:VIa

[552] Anonymous. Volume information. Isis, 45(1):??, May 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227675.

Anonymous:1954:FMb

[553] Anonymous. Front matter. Isis, 45(2):129–130, July 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227355.

Sarton:1954:DBV

[554] George Sarton. The death and burial of Vesalius, and, incidentally, of Cicero. Isis, 45(2):131–137, July 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227356.

OMalley:1954:AVP

[555] C. Donald O’Malley. Andreas Vesalius’ pilgrimage. Isis, 45(2):138–144, July 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227357.

Thorndike:1954:OFC

[556] Lynn Thorndike. Oresme and Fourteenth Century commentaries on the meteorologica. Isis, 45(2):145–152, July 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227358.

Boas:1954:EVB

[557] Marie Boas. An early version of Boyle’s: Sceptical Chymist. Isis, 45(2): 153–168, July 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227359.

Knowlton:1954:HC

[558] A. A. Knowlton. Henry Crew (1859–1953). Isis, 45(2):169–174, July 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227360.

Shu-hua:1954:OBI

[559] Li Shu-hua. Origine de la Boussole II. Aimant et Boussole. (French) [Origin of Boussole. II. Aimant and Boussole]. Isis, 45(2):175–196, July 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227361.

Forbes:1954:NC

[560] R. J. Forbes and A. Hunter Dupree. Notes & correspondence. Isis, 45(2): 197–198, July 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227362.

Sarton:1954:QA

[561] George Sarton and R. C. Archibald. Queries & answers. Isis, 45(2): 199, July 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227363.

Anonymous:1954:DPH

[562] Anonymous. Dissertations in progress in the history of science. Isis, 45(2):199–200, July 1954. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227364.

Anonymous:1954:DHS

[563] Anonymous. Doctorates in the history of science. Isis, 45(2):200, July 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227365.

Stauffer:1954:BRB

[564] Robert C. Stauffer. Book review: Main Currents of Scientific Thought. A History of the Sciences by S. F. Mason. Isis, 45(2):201–202, July 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227366.

Sarton:1954:BRBc

[565] George Sarton. Book review: Science, Medicine, and History. Essays on the Evolution of Scientific Thought and Medical Practice, Written in Honour of Charles Singer by E. Ashworth Underwood. Isis, 45(2):202– 204, July 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227367.

Sarton:1954:BRBd

[566] George Sarton. Book review: Histoire des sciences et des techniques. Bibliographie by François Russo. Isis, 45(2):204–205, July 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227368.

Boas:1954:BRB

[567] Marie Boas. Book review: Science and Religion by Charles E. Raven. Isis, 45(2):205–206, July 1954. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227369.

Dubos:1954:BRB

[568] Rene J. Dubos. Book review: The Conquest of Plague by L. Fabian Hirst. Isis, 45(2):206–208, July 1954. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227370.

Sigerist:1954:BRBb

[569] Henry E. Sigerist. Book review: The Unique Influence of the Johns Hopkins University on American Medicine by Richard H. Shryock. Isis, 45(2):208–209, July 1954. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227371.

Johnson:1954:BRBa

[570] Francis R. Johnson. Book review: Science and Religion in Elizabethan England by Paul H. Kocher. Isis, 45(2):209–212, July 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227372.

vanWijk:1954:BRB

[571] W. E. van Wijk. Book review: Computus med särskild hänsyn till Runstaven och den Borgerliga Kalendern by Nils Lithberg. Isis, 45(2):212– 213, July 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227373.

Johnson:1954:BRBb

[572] Francis R. Johnson. Book review: Dialogue on the Great World Systems by Galileo Galilei; T. Salusbury; Giorgio de Santillana; Dialogue concerning the Two Chief World Systems — Ptolemaic and Copernican by Galileo Galilei; Stillman Drake. Isis, 45(2):213–215, July 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227374.

Cohen:1954:BRBa

[573] I. Bernard Cohen. Book review: Sir Hans Sloane and the British Museum by G. R. De Beer; Sir Joseph Banks, the Autocrat of the Philosophers by Hector Charles Cameron. Isis, 45(2):215–218, July 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227375.

Brown:1954:BRB

[574] Sanborn C. Brown. Book review: An American in Europe. The Life of Benjamin Thompson, Count Rumford by Egon Larsen. Isis, 45(2):218– 220, July 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227376.

Schoenwald:1954:BRB

[575] Richard L. Schoenwald. Book review: The Life and Work of Sigmund Freud. Vol. I: The Formative Years and the Great Discoveries 1856–1900 by Ernest Jones. Isis, 45(2):220–221, July 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227377.

Leikind:1954:BRB

[576] Morris C. Leikind. Book review: Louis Pasteur, Free Lance of Science by René Dubos; The White Plague. Tuberculosis, Man and Society by Jean Dubos; René Dubos. Isis, 45(2):221–223, July 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227378.

Urdang:1954:BRB

[577] George Urdang. Book review: Hans Spemann, Ein Meister der Entwicklungsphysiologie, sein Leben und sein Werk by Otto Mangold. Isis, 45(2): 223–224, July 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227379.

Farber:1954:BRBa

[578] Eduard Farber. Book review: Salpetersäure aus Ammoniak by Alwin Mittasch. Isis, 45(2):224, July 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227380.

Anonymous:1954:FMc

[579] Anonymous. Front matter. Isis, 45(3):225–226, September 1954. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226709.

Lurie:1954:LAR

[580] Edward Lurie. Louis Agassiz and the races of Man. Isis, 45(3):227–242, September 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226710.

Guerlac:1954:PN

[581] Henry Guerlac. The poets’ nitre. Isis, 45(3):243–255, September 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226711.

Schneer:1954:RHG

[582] Cecil Schneer. The rise of historical geology in the Seventeenth Century. Isis, 45(3):256–268, September 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226712.

Clagett:1954:KAB

[583] Marshall Clagett. King Alfred and the Elements of Euclid. Isis, 45(3): 269–277, September 1954. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226713.

Duveen:1954:ILCa

[584] Denis I. Duveen and Herbert S. Klickstein. The introduction of Lavoisier’s chemical nomenclature into America. Isis, 45(3):278–292, September 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226714.

Leake:1954:NC

[585] Chauncey D. Leake and Maria Rooseboom. Notes & correspondence. Isis, 45(3):293, September 1954. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226715.

Clark:1954:BRB

[586] Joseph T. Clark. Book review: Ancient Formal Logic by I. M. Bocheński; The Propositional Logic of Boethius by Karl Dürr; Truth and Consequence in Medieval Logic by Ernest A. Moody. Isis, 45(3):294–301, September 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226716.

Zirkle:1954:BRBb

[587] Conway Zirkle. Book review: Plough and Pasture: The Early History of Farming by E. Cecil Curwen; Gudmund Hatt. Isis, 45(3):301–302, September 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226717.

Zirkle:1954:BRBc

[588] Conway Zirkle. Book review: History of the Rose by Roy E. Shepard. Isis, 45(3):302–303, September 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226718.

Brozek:1954:BRB

[589] Joseph Brožek. Book review: History of American Psychology by A. A. Roback. Isis, 45(3):303–304, September 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226719.

Baumer:1954:BRB

[590] Franklin L. Baumer. Book review: The Triumph of the English Language by Richard Foster Jones. Isis, 45(3):304–305, September 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226720.

DeLacy:1954:BRB

[591] Phillip De Lacy. Book review: Epicurus and His Philosophy by Norman Wentworth DeWitt. Isis, 45(3):306–308, September 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226721.

Stahlman:1954:BRBa

[592] William D. Stahlman. Book review: Heraclitus, The Cosmic Fragments by G. S. Kirk. Isis, 45(3):308–309, September 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226722.

Stahlman:1954:BRBb

[593] William D. Stahlman. Book review: Catalogue of Astrological and Mythological Illuminated Manuscripts of the Latin Middle Ages by Fritz Saxl; Hans Meier; Harry Bober. Isis, 45(3):309–311, September 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226723.

Sigerist:1954:BRBc

[594] Henry E. Sigerist. Book review: Michel Servet, hérétique et martyr, 1553–1953 by Roland H. Bainton; Autour de Michel Servet et de Sébastien Castellion by B. Becker. Isis, 45(3):311–313, September 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226724.

Cohen:1954:BRBb

[595] I. Bernard Cohen. Book review: Michael Servetus, Humanist and Martyr by John F. Fulton; Hunted Heretic. The Life and Death of Michael Servetus, 1511–1553 by Roland H. Bainton; Michael Servetus by Charles Donald O’Malley. Isis, 45(3):313–314, September 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226725.

Bridgman:1954:BRB

[596] P. W. Bridgman. Book review: Les Premiers congrès de physique Solvay et l’orientation de la physique depuis 1911 by Maurice de Broglie. Isis, 45 (3):314–315, September 1954. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226726.

Fulton:1954:BRB

[597] John F. Fulton. Book review: Joseph Barcroft, 1872–1947 by Kenneth James Franklin. Isis, 45(3):315–316, September 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226727.

Gardner:1954:BRB

[598] Lester D. Gardner. Book review: The Papers of Wilbur and Orville Wright by Marvin W. McFarland. Isis, 45(3):316–318, September 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226728.

Anonymous:1954:BMb

[599] Anonymous. Back matter. Isis, 45(3):??, September 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226729.

Anonymous:1954:FMd

[600] Anonymous. Front matter. Isis, 45(4):321–322, December 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226778.

Harrison:1954:BM

[601] Thomas P. Harrison. Birds in the Moon. Isis, 45(4):323–330, December 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226779.

Plessner:1954:PTP

[602] M. Plessner. The place of the turba philosophorum in the development of alchemy. Isis, 45(4):331–338, December 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226780.

Greene:1954:SAA

[603] John C. Greene. Some aspects of American astronomy 1750–1815. Isis, 45(4):339–358, December 1954. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226781.

Multhauf:1954:JRO

[604] Robert P. Multhauf. John of Rupescissa and the origin of medical chemistry. Isis, 45(4):359–367, December 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226782.

Duveen:1954:ILCb

[605] Denis I. Duveen and Herbert S. Klickstein. The introduction of Lavoisier’s chemical nomenclature into america: Part 2. Isis, 45(4): 368–382, December 1954. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226783.

Ward:1954:QA

[606] Eric Ward, Harold R. Wainerdi, and George Sarton. Queries & answers. Isis, 45(4):383–384, December 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226784.

Boyer:1954:BRB

[607] Carl B. Boyer. Book review: Mathematics in Western Culture by Morris Kline. Isis, 45(4):385–387, December 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226785.

Sarton:1954:BRBe

[608] George Sarton. Book review: Etudes Américaines by Emanuel de Margerie. Isis, 45(4):387–388, December 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226786.

Sarton:1954:BRBf

[609] George Sarton. Book review: Catalogus Codicum Astrologorum Graecorum. Isis, 45(4):388–389, December 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226787.

Temkin:1954:RBG

[610] Owsei Temkin. Book review: Nature and the Greeks by Erwin Schrödinger. Isis, 45(4):389–390, December 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226788.

Stahlman:1954:BRBc

[611] William D. Stahlman. Book review: Miracle and Natural Law in Graeco-Roman and Early Christian Thought by Robert M. Grant. Isis, 45(4): 390–391, December 1954. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226789.

Sonnedecker:1954:BRB

[612] Glenn Sonnedecker. Book review: Agnus Castus, a Middle English Herbal Reconstructed from Various Manuscripts by Gösta Brodin. Isis, 45(4): 392–393, December 1954. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226790.

Plessner:1954:BRB

[613] M. Plessner. Book review: A Mirror for Princes. The Qabus nama by Kai’ Ka’ius ibn Iskandar; Reuben Levy. Isis, 45(4):393–395, December 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226791.

Ware:1954:BRB

[614] James R. Ware. Book review: China in the Sixteenth Century: The Journals of Matthew Ricci: 1583–1610 by L. J. Gallagher. Isis, 45(4):

395–397, December 1954. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226792.

Hellman:1954:BRB

[615] C. Doris Hellman. Book review: Tamburlaine’s Malady and Other Essays on Astrology in Elizabethan Drama by Johnstone Parr. Isis, 45(4):398– 399, December 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226793.

Vartanian:1954:BRB

[616] Aram Vartanian. Book review: The European Mind. The Critical Years (1680–1715) by Paul Hazard; J. Lewis May. Isis, 45(4):399–400, December 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226794.

Williams:1954:BRB

[617] L. Pearce Williams. Book review: The Triumph of Science and Reason, 1660–1685. The Rise of Modern Europe, vol. 6 by Frederick L. Nussbaum. Isis, 45(4):400–402, December 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226795.

Zirkle:1954:BRBd

[618] Conway Zirkle. Book review: Thomas Jefferson’s Farm Book by Edwin Morris Betts. Isis, 45(4):402–403, December 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226796.

Fleming:1954:BRBb

[619] Donald Fleming. Book review: Science and Industry in the Nineteenth Century by J. D. Bernal. Isis, 45(4):403–405, December 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226797.

Farber:1954:BRBb

[620] Eduard Farber. Book review: Wilhelm Ostwald, Mein Vater by Grete Ostwald. Isis, 45(4):405–406, December 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226798.

Ihde:1954:BRBb

[621] Aaron J. Ihde. Book review: Justus von Liebig by Hertha Von Dechend. Isis, 45(4):406–407, December 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226799.

Cohen:1954:BRBc

[622] I. Bernard Cohen. Book review: Nobel Prize Winners in Physics, 1901– 1950 by Niels H. de V. Heathcote; Nobel Prize Winners in Medicine and Physiology, 1901–1950 by Lloyd G. Stevenson; Nobel Prize Winners in Chemistry, 1901–1950 by Eduard Farber. Isis, 45(4):407–408, December 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226800.

Joravsky:1954:BRB

[623] David Joravsky. Book review: Soviet Science by Ruth G. Christman. Isis, 45(4):408–409, December 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226801.

Williams:1954:HSS

[624] L. Pearce Williams, Carl B. Boyer, and Max H. Fisch. The History of Science Society. Isis, 45(4):410–412, December 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226802.

Anonymous:1954:VIb

[625] Anonymous. Volume information. Isis, 45(4):413–417, December 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226777.

Anonymous:1954:BMc

[626] Anonymous. Back matter. Isis, 45(4):??, December 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226803.

Anonymous:1955:FMa

[627] Anonymous. Front matter. Isis, 46(1):1–2, March 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/226818.

Joravsky:1955:SVH

[628] David Joravsky. Soviet views on the history of science. Isis, 46(1):3– 13, March 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226819.

Fleming:1955:GMB

[629] Donald Fleming. Galen on the motions of the blood in the heart and lungs. Isis, 46(1):14–21, March 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226820.

Reingold:1955:NAH

[630] Nathan Reingold. The national archives and the history of science in America. Isis, 46(1):22–28, March 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226821.

Heathcote:1955:FIE

[631] N. H. de V. Heathcote. Franklin’s introduction to electricity. Isis, 46(1): 29–35, March 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226822.

Kerker:1955:HBD

[632] Milton Kerker. Herman Boerhaave and the development of pneumatic chemistry. Isis, 46(1):36–49, March 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226823.

Sarton:1955:QA

[633] George Sarton, I. B. Cohen, and Irving Massey. Queries & answers. Isis, 46(1):50–52, March 1955. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226824.

Taton:1955:NC

[634] René Taton, T. D. Phillips, Lynn Thorndike, Charles W. David, Claude K. Deischer, and Harvey P. Hall. Notes & correspondence. Isis, 46(1):53–55, March 1955. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226825.

Anonymous:1955:DHSa

[635] Anonymous. Doctorates in the history of science. Isis, 46(1):56, March 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226827.

Anonymous:1955:DPH

[636] Anonymous. Dissertations in progress in the history of science. Isis, 46 (1):56, March 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226826.

Boone:1955:BRB

[637] William W. Boone. Book review: Conventional Logic and Modern Logic by Joseph T. Clark. Isis, 46(1):57, March 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226828.

Boyer:1955:BRB

[638] Carl B. Boyer. Book review: Geschichte der Mathematik by Joseph Ehrenfried Hofmann. Isis, 46(1):57–59, March 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226829.

Rogers:1955:BRB

[639] Francis M. Rogers. Book review: Filosofia da Historia Maritima Portuguesa by De Frag. M.; Júlio Gonçalves. Isis, 46(1):59–61, March 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226830.

Sarton:1955:BRBa

[640] George Sarton. Book review: Animals, Men and Myths. A History of the Influence of Animals on Civilization and Culture by Morus. Isis, 46(1): 61–64, March 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226831.

Zirkle:1955:BRB

[641] Conway Zirkle. Book review: Indian Corn in Old America by Paul Weatherwax. Isis, 46(1):64–65, March 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226832.

Diller:1955:BRB

[642] Aubrey Diller. Book review: Ancient Science and Modern Civilization by George Sarton. Isis, 46(1):65–66, March 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226833.

Clagett:1955:BRB

[643] Marshall Clagett. Book review: Robert Grosseteste and the Origins of Experimental Science 1100–1700 by A. C. Crombie. Isis, 46(1):66–69, March 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226834.

Neugebauer:1955:BRB

[644] O. Neugebauer. Book review: On the Revolutions of the Heavenly Spheres by Nicolaus Copernicus; Charles Glenn Wallis. Isis, 46(1):69–71, March 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226835.

Sarton:1955:BRBb

[645] George Sarton. Book review: Journal by Isaac Beeckman. Isis, 46(1): 71–72, March 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226836.

OMalley:1955:BRB

[646] Charles D. O’Malley. Book review: Epistolae et epistolae ad eum datae, quas cum proemio ac notis germanice scriptis by Nicolai Stenonis; Gustav Scherz; Joanne Raeder. Isis, 46(1):72–74, March 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226837.

Schneider:1955:BRB

[647] Heinrich Schneider. Book review: Leibnitz and the Seventeenth-Century Revolution by R. W. Meyer; J. P. Stern. Isis, 46(1):74–76, March 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226838.

Woolf:1955:BRB

[648] Harry Woolf. Book review: The Rittenhouse Orrery, Princeton’s Eighteenth-Century Planetarium, 1767–1954 by Howard G. Rice. Isis, 46(1):76–77, March 1955. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226839.

Bell:1955:BRB

[649] Whitfield J. Bell, Jr. Book review: Some Account of the Pennsylvania Hospital by Benjamin Franklin. Isis, 46(1):77–78, March 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226840.

Galdston:1955:BRB

[650] Iago Galdston. Book review: Twenty Years of Psychoanalysis by Franz Alexander; Helen Ross; A History of Psychoanalysis in America by C. P. Oberndorf. Isis, 46(1):78–80, March 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226841.

Anonymous:1955:BMa

[651] Anonymous. Back matter. Isis, 46(1):??, March 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226842.

Anonymous:1955:FMb

[652] Anonymous. Front matter. Isis, 46(2):81, June 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227119.

Ackerknecht:1955:GFA

[653] Erwin H. Ackerknecht. George Forster, Alexander von Humboldt, and ethnology. Isis, 46(2):83–95, June 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227120.

Pelseneer:1955:PPR

[654] Jean Pelseneer. Peut-on planifier la recherche scientifique? Ce que nous dit l’histoire des mathématiques. Isis, 46(2):95–98, June 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227121.

Cohen:1955:FBN

[655] I. Bernard Cohen. Franklin, Boerhaave, Newton, Boyle & the absorption of heat in relation to color. Isis, 46(2):99–104, June 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227122.

Huber:1955:EMK

[656] Peter Huber. Zu einem mathematischen Keilschrifttext (VAT 8512). (German) [On a mathematical cuneiform text (VAT 8512)]. Isis, 46(2): 104–106, June 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227123.

Levey:1955:SG

[657] Martin Levey. Solomon Gandz, 1884–1954. Isis, 46(2):107–110, June 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227124.

Cohen:1955:ECB

[658] I. Bernard Cohen and Katharine Strelsky. Eightieth critical bibliography of the history of science and its cultural influences (to 1 January 1954).

Isis, 46(2):111–220, June 1955. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227125.

Anonymous:1955:BMb

[659] Anonymous. Back matter. Isis, 46(2):??, June 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227126.

Anonymous:1955:FMc

[660] Anonymous. Front matter. Isis, 46(3):221–222, September 1955. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226340.

Drabkin:1955:RAP

[661] I. E. Drabkin. Remarks on ancient psychopathology. Isis, 46(3):223–234, September 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226341.

Carmody:1955:NAW

[662] Francis J. Carmody. Notes on the astronomical works of Thabit b. Qurra. Isis, 46(3):235–242, September 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226342.

Hindle:1955:QBS

[663] Brooke Hindle. The Quaker background and science in colonial Philadelphia. Isis, 46(3):243–250, September 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226343.

Dieckmann:1955:FED

[664] Herbert Dieckmann. The first edition of Diderot’s pensees sur l’interpretation de la nature. Isis, 46(3):251–267, September 1955. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226344.

Cohen:1955:NCD

[665] I. Bernard Cohen. A note concerning Diderot and Franklin. Isis, 46 (3):268–272, September 1955. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226345.

Thorndike:1955:TPA

[666] Lynn Thorndike. The true place of astrology in the history of science. Isis, 46(3):273–278, September 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226346.

Cohen:1955:NC

[667] I. Bernard Cohen, Roger Hahn, Lloyd Espenschied, Marshall Clagett, Bertha W. Rubinstein, George Sarton, Vasco Ronch, Bruno Boni, Chester G. Moore, Jane D. Oppenheimer, Vasco Ronchi, and Roberto Almagia. Notes & correspondence. Isis, 46(3):278–283, September 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226347.

Anonymous:1955:DHSb

[668] Anonymous. Doctorates in the history of science. Isis, 46(3):284, September 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226348.

Kerker:1955:THS

[669] Milton Kerker, Paul Gilbert, Carl W. Condit, and A. R. Hall. Teaching the history of science. Isis, 46(3):284–286, September 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226349.

Lenzen:1955:BRBa

[670] V. F. Lenzen. Book review: Concepts of Space. The History of Theories of Space in Physics by Max Jammer; God and Space-Time. Deity in the Philosophy of Samuel Alexander by Alfred P. Stiernotte. Isis, 46(3): 287–289, September 1955. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226350.

Schofield:1955:BRBa

[671] Robert E. Schofield. Book review: The English Pioneers of Anaesthesia (Beddoes, Davy, and Hickman) by F. F. Cartwright. Isis, 46(3):289–290, September 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226351.

White:1955:BRB

[672] Lynn White, Jr. Book review: A History of Mechanical Inventions by Abbott Payson Usher. Isis, 46(3):290–293, September 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226352.

Sarton:1955:BRBc

[673] George Sarton. Book review: A History of Technology. Volume I: From Early Times to the Fall of the Ancient Empires by Charles Singer; E. J. Holmyard; A. R. Hall. Isis, 46(3):294–296, September 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226353.

Green:1955:BRB

[674] Robert M. Green. Book review: Galen of Pergamon by George Sarton. Isis, 46(3):296–297, September 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226354.

Maier:1955:BRB

[675] Anneliese Maier. Book review: The Medieval Science of Weights by Ernest A. Moody; Marshall Clagett. Isis, 46(3):297–300, September 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226355.

Janssens:1955:BRB

[676] Emile Janssens. Book review: The Nautical Chart of 1424 and the Early Discovery and Cartographical Representation of America. A Study of the History of Early Navigation and Cartography by Armando Cortesao. Isis, 46(3):300–302, September 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226356.

Goodrich:1955:BRB

[677] L. Carrington Goodrich. Book review: Science and Civilisation in China. Volume I, Introductory Orientations by Joseph Needham; Wang Ling. Isis, 46(3):302–304, September 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226357.

Boas:1955:BRB

[678] Marie Boas. Book review: The Scientific Revolution, 1500–1800. The Formation of the Modern Scientific Attitude by A. R. Hall. Isis, 46(3): 304–305, September 1955. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226358.

Heathcote:1955:BRB

[679] N. H. de V. Heathcote. Book review: Commentary on the Effects of Electricity on Muscular Motion by Luigi Galvani; Margaret Glover Foley.

Isis, 46(3):305–309, September 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226359.

Boas:1955:AD

[680] Marie Boas, Charles C. Gillispie, and Stanley M. Loomis. Administrative documents. Isis, 46(3):310–316, September 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226360.

Anonymous:1955:BMc

[681] Anonymous. Back matter. Isis, 46(3):??, September 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226361.

Anonymous:1955:FMd

[682] Anonymous. Front matter. Isis, 46(4):317–318, December 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227573.

Fleming:1955:WHP

[683] Donald Fleming. William Harvey and the pulmonary circulation. Isis, 46 (4):319–327, December 1955. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227574.

Siegfried:1955:AUS

[684] Robert Siegfried. An attempt in the united states to resolve the differences between the oxygen and the phlogiston theories. Isis, 46(4): 327–336, December 1955. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227575.

Hesse:1955:ADC

[685] Mary B. Hesse. Action at a distance in classical physics. Isis, 46(4): 337–353, December 1955. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227576.

Smith:1955:SCD

[686] Cyril Stanley Smith. A Sixteenth-Century decimal system of weights. Isis, 46(4):354–357, December 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227577.

Thorndike:1955:MQM

[687] Lynn Thorndike. More questions on the meteorologica. Isis, 46(4):357– 360, December 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227578.

Sarton:1955:HPS

[688] George Sarton. “Historians and philosophers” of science. Isis, 46(4): 360–366, December 1955. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227579.

David:1955:NC

[689] Charles W. David, William P. D. Wightman, and L. Voet. Notes & correspondence. Isis, 46(4):367, December 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227580.

Sarton:1955:BRBd

[690] George Sarton. Book review: Science Awakening by B. L. Van der Waerden; Arnold Dresden. Isis, 46(4):368–369, December 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227581.

Ihde:1955:BRB

[691] Aaron J. Ihde. Book review: From Classical to Modern Chemistry. Some Historical Sketches by A. J. Berry. Isis, 46(4):369, December 1955. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227582.

Lenzen:1955:BRBb

[692] V. F. Lenzen. Book review: The Development of the Concept of the Electric Charge. Electricity from the Greeks to Coulomb by Duane Roller; Duane H. D. Roller; James Bryant Conant; Leonard K. Nash. Isis, 46 (4):369–370, December 1955. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227583.

Cottrell:1955:BRB

[693] G. W. Cottrell, Jr. Book review: Die illustrierten Vogelbücher, ihre Geschichte und Bibliographie by Claus Nissen; Fine Bird Books 1700– 1900 by Sacheverell Sitwell; Handasyde Buchanan; James Fisher. Isis, 46(4):370–372, December 1955. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227584.

Thorndike:1955:BRB

[694] Lynn Thorndike. Book review: Die Maschine im Leben der Völker: Ein Überblick von der Urzeit bis zur Renaissance, mit 205 Abbildungen und I Farbtafel by Franz Maria Feldhaus. Isis, 46(4):372, December 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227585.

Montagu:1955:BRBa

[695] Ashley Montagu. Book review: The Law of Primitive Man. A Study in Comparative Legal Dynamics by E. Adamson Hoebel. Isis, 46(4):373, December 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227586.

Hay:1955:BRB

[696] William H. Hay. Book review: Of Learned Ignorance by Nicholas Cusanus; Fr. Germain Heron; W. Stark. Isis, 46(4):373–374, December 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227587.

Fleming:1955:BRB

[697] Donald Fleming. Book review: Cotton Mather by Otho T. Beall,; Richard H. Shryock. Isis, 46(4):374–376, December 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227588.

Schofield:1955:BRBb

[698] Robert E. Schofield. Book review: William Withering of Birmingham by T. Whitmore Peck; K. Douglas Wilkinson. Isis, 46(4):376–377, December 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227589.

Kuhn:1955:BRB

[699] Thomas S. Kuhn. Book review: New Studies in the Philosophy of Descartes by Norman Kemp Smith; Descartes’ Philosophical Writings by Norman Kemp Smith; The Method of Descartes. A Study of the Regulae by L. J. Beck. Isis, 46(4):377–380, December 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227590.

Seeger:1955:BRB

[700] Raymond J. Seeger. Book review: Henri Poincaré, Critic of Crisis. Reflections on His Universe of Discourse by Tobias Dantzig. Isis, 46 (4):380–382, December 1955. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227591.

Saunders:1955:BRBa

[701] J. B. de C. M. Saunders. Book review: William H. Welch and the Rise of Modern Medicine by Donald Fleming; Oscar Handlin. Isis, 46(4): 382–383, December 1955. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227592.

Saunders:1955:BRBb

[702] J. B. de C. M. Saunders. Book review: Forgotten Leaders in Modern Medicine, Valentin, Gouby, Remak, Auerbach by Bruno Kisch. Isis, 46 (4):383–385, December 1955. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227593.

Montagu:1955:BRBb

[703] Ashley Montagu. Book review: Charles Darwin and the Golden Rule by William Emerson Ritter; Edna Watson Bailey. Isis, 46(4):385–386, December 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227594.

Freeman:1955:BRB

[704] John Finley Freeman. Book review: Beyond the Hundredth Meridian. John Wesley Powell and the Second Opening of the West by Wallace Stegner. Isis, 46(4):386–387, December 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227595.

Ackerknecht:1955:BRB

[705] Erwin H. Ackerknecht. Book review: Humboldt: The Life and Times of Alexander von Humboldt, 1769–1859 by Helmut de Terra. Isis, 46 (4):387–388, December 1955. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227596.

Anonymous:1955:VI

[706] Anonymous. Volume information. Isis, 46(4):389–394, December 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227572.

Anonymous:1955:BMd

[707] Anonymous. Back matter. Isis, 46(4):??, December 1955. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227597.

Anonymous:1956:FMa

[708] Anonymous. Front matter. Isis, 47(1):1–2, March 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227541.

Panofsky:1956:GCA

[709] Erwin Panofsky. Galileo as a critic of the arts: Aesthetic attitude and scientific thought. Isis, 47(1):3–15, March 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227542.

Schofield:1956:JWP

[710] Robert E. Schofield. Josiah Wedgwood and a proposed Eighteenth-Century industrial research organization. Isis, 47(1):16–19, March 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227543.

Tolles:1956:PFS

[711] Frederick B. Tolles. Philadelphia’s first scientist: James Logan. Isis, 47 (1):20–30, March 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/227544.

Anonymous:1956:GSM

[712] Anonymous. The George Sarton Medal. Isis, 47(1):31–32, March 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227545.

Kennedy:1956:PTI

[713] E. S. Kennedy. Parallax theory in Islamic astronomy. Isis, 47(1):33– 53, March 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227546.

Leake:1956:NC

[714] Chauncey D. Leake, Harry Alpert, W. H. Brindley, F. Willett, and O. Neugebauer. Notes & correspondence. Isis, 47(1):54–58, March 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227547.

Truesdell:1956:QA

[715] C. Truesdell, Paul W. Hagensick, Waclaw Slabczynski, L. A. Mayer, and George Sarton. Queries & answers. Isis, 47(1):59–60, March 1956.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227548.

Anonymous:1956:DPH

[716] Anonymous. Dissertations in progress in the history of science. Isis, 47 (1):61, March 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227549.

Pinkus:1956:THS

[717] L. F. Pinkus, Duane H. D. Roller, Charles C. Gillispie, John C. Greene, and L. E. Klopfer. Teaching the history of science. Isis, 47(1):61–63, March 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227550.

Boyer:1956:HSS

[718] Carl B. Boyer, L. Pearce Williams, and Max H. Fisch. The History of Science Society. Isis, 47(1):64–65, March 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227551.

Fleming:1956:BRBa

[719] Donald Fleming. Book review: Essays on the Social History of Science by S. Lilley. Isis, 47(1):65–66, March 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227552.

Gilbert:1956:BRB

[720] W. Paul Gilbert. Book review: Tuning and Temperament. A Historical Survey by J. Murray Barbour. Isis, 47(1):66–67, March 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227553.

Schofield:1956:BRBa

[721] Robert E. Schofield. Book review: The Aim and Structure of Physical Theory by Pierre Duhem; Philip P. Wiener. Isis, 47(1):67–69, March 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227554.

Fulton:1956:BRB

[722] John F. Fulton. Book review: A History of Medicine by Ralph H. Major. Isis, 47(1):69–70, March 1956. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227555.

Temkin:1956:BRB

[723] Owsei Temkin. Book review: Parva Naturalia by Aristotle. Isis, 47(1): 70–71, March 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227556.

Stahlman:1956:BRB

[724] William D. Stahlman. Book review: Astrology in Roman Law and Politics. I. Astrology in Rome until the End of the Principate by Frederick H. Cramer. Isis, 47(1):71–74, March 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227557.

Sarton:1956:BRBa

[725] George Sarton. Book review: Libro de Agricultura by Ibn Bas.s.al. Isis, 47(1):74–77, March 1956. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227558.

Taylor:1956:BRB

[726] F. Sherwood Taylor. Book review: The Crime of Galileo by Giorgio de Santillana. Isis, 47(1):77–78, March 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227559.

Rosen:1956:BRB

[727] Edward Rosen. Book review: Galileo as a Critic of the Arts by Erwin Panofsky. Isis, 47(1):78–80, March 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227560.

Marx:1956:BRB

[728] Leo Marx. Book review: The New England Mind: The Seventeenth Century by Perry Miller; The New England Mind: From Colony to Province. Isis, 47(1):80–81, March 1956. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227561.

Sarton:1956:BRBb

[729] George Sarton. Book review: Monge, le savant ami de Napoléon Bonaparte, 1746–1818 by Paul V. Aubry. Isis, 47(1):81–84, March 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227562.

Duveen:1956:BRB

[730] Denis I. Duveen. Book review: Lavoisier, théoricien et expérimentateur by Maurice Daumas. Isis, 47(1):84–85, March 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227563.

Guerlac:1956:BRB

[731] Henry Guerlac. Book review: A Bibliography of the Works of Antoine Laurent Lavoisier, 1743–1794 by Denis I. Duveen; Herbert S. Klickstein. Isis, 47(1):85–88, March 1956. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227564.

Gingerich:1956:BRB

[732] Owen Gingerich. Book review: William Herschel by J. B. Sidgwick. Isis, 47(1):88–89, March 1956. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227565.

Baumer:1956:BRB

[733] Franklin L. Baumer. Book review: European Thought in the Eighteenth Century, from Montesquieu to Lessing by Paul Hazard; J. Lewis May. Isis, 47(1):89–90, March 1956. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227566.

Leake:1956:BRBa

[734] Chauncey D. Leake. Book review: El Hipocratismo en México by J. J. Izquierdo. Isis, 47(1):90–91, March 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227567.

Glass:1956:BRB

[735] Bentley Glass. Book review: Apes, Angels, & Victorians. The Story of Darwin, Huxley, and Evolution by William Irvine. Isis, 47(1):91– 93, March 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227568.

Ihde:1956:BRBa

[736] Aaron J. Ihde. Book review: Michael Tswett’s First Paper on Chromatography by Gerhard Hesse; Herbert Weil; The Elements of Chromatography by Trevor Illtyd Williams. Isis, 47(1):93–94, March 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227569.

Brozek:1956:BRBa

[737] Josef Brožek. Book review: Jan Evangelista Purkyne by Jarmila Psotnícková. Isis, 47(1):94, March 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227570.

Anonymous:1956:BMa

[738] Anonymous. Back matter. Isis, 47(1):95–96, March 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227571.

Anonymous:1956:FMb

[739] Anonymous. Front matter. Isis, 47(2):97–98, June 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227332.

Clagett:1956:GS

[740] Marshall Clagett, I. Bernard Cohen, I. E. Drabkin, John F. Fulton, Henry Guerlac, and Conway Zirkle. George Sarton 1884–1956. Isis, 47(2):99– 100, June 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227333.

Nash:1956:ODC

[741] Leonard K. Nash. The origin of Dalton’s chemical atomic theory. Isis, 47(2):101–116, June 1956. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227334.

vonBaer:1956:GOM

[742] Karl Ernst von Baer and Charles Donald O’Malley. On the Genesis of the ovum of mammals and of man. Isis, 47(2):117–153, June 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227335.

Stahlman:1956:ADA

[743] William D. Stahlman. Astronomical dating applied to a type of astrological illustration. Isis, 47(2):154–160, June 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227336.

Merton:1956:BST

[744] E. S. Merton. The botany of Sir Thomas Browne. Isis, 47(2):161–171, June 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227337.

Stoller:1956:NTP

[745] Leo Stoller. A note on Thoreau’s place in the history of phenology. Isis, 47(2):172–181, June 1956. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227338.

Panofsky:1956:NC

[746] Erwin Panofsky, P. H. Brans, Robert P. Multhaue, and Raoul Michel May. Notes & correspondence. Isis, 47(2):182–188, June 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227339.

Fleming:1956:BRBb

[747] Donald Fleming. Book review: Science and Invention by Mitchell Wilson. Isis, 47(2):188–190, June 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227340.

Lenzen:1956:BRB

[748] V. F. Lenzen. Book review: Science and the Human Imagination by Mary B. Hesse. Isis, 47(2):190–191, June 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227341.

Barber:1956:BRB

[749] Bernard Barber. Book review: Science and Social Action by W. J. H. Sprott. Isis, 47(2):191–192, June 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227342.

Leake:1956:BRBb

[750] Chauncey D. Leake. Book review: Psychoanalysis and Ethics by Lewis Samuel Feuer; Marvin Farber. Isis, 47(2):192–194, June 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227343.

Boyer:1956:BRB

[751] Carl B. Boyer. Book review: Die Mathematische Denkweise by Andreas Speiser. Isis, 47(2):194–195, June 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227344.

Leikind:1956:BRB

[752] Morris C. Leikind. Book review: The History and Conquest of Common Diseases by Walter R. Bett; Pomp and Pestilence, Infectious Disease, Its Origins and Conquest by Ronald Hare. Isis, 47(2):195–196, June 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227345.

Leake:1956:BRBc

[753] Chauncey D. Leake. Book review: Montaña y los orígenes del movimento social y científico de México by José Joaquín Izquierdo. Isis, 47(2):196– 197, June 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227346.

Johnson:1956:BRB

[754] Francis R. Johnson. Book review: The Equatorie of the Planetis by Derek J. Price. Isis, 47(2):197–198, June 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227347.

Clagett:1956:BRB

[755] Marshall Clagett. Book review: Die Practica des Algorismus Ratisbonensis. Ein Rechenbuch des Benediktinerklosters St. Emmeram aus der Mitte des 15. Jahrhunderts nach den Handschriften der Munchner Staatsbibliothek und der Stiftsbibliothek St. Florian by Kurt Vogel. Isis, 47(2): 198–199, June 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/227348.

Pagel:1956:BRBa

[756] Walter Pagel. Book review: The Art of Ramon Lull. An Approach to It through Lull’s Theory of the Elements by Frances A. Yates. Isis, 47(2): 199–200, June 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227349.

Schullian:1956:BRBa

[757] Dorothy M. Schullian. Book review: Robert Estienne, Royal Printer by Elizabeth Armstrong. Isis, 47(2):200–201, June 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227350.

Schullian:1956:BRBb

[758] Dorothy M. Schullian. Book review: Jacopo Aconcio by Charles Donald O’Malley. Isis, 47(2):201–203, June 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227351.

Schofield:1956:BRBb

[759] Robert E. Schofield. Book review: The Crucible. The Story of Joseph Priestley, LL.D., F.R.S. by John Graham Gillam. Isis, 47(2):203–204, June 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227352.

Brozek:1956:BRBb

[760] Josef Brožek. Book review: Jan Evangelista Purkyne, badatel — národní buditel by B. Nemec; O. Matousek. Isis, 47(2):204, June 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227353.

Anonymous:1956:BMb

[761] Anonymous. Back matter. Isis, 47(2):205–207, June 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227354.

Anonymous:1956:FMc

[762] Anonymous. Front matter. Isis, 47(3):209, September 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226887.

Guerlac:1956:NLS

[763] Henry Guerlac. A note on Lavoisier’s scientific education. Isis, 47(3): 211–216, September 1956. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226888.

Partington:1956:LWJa

[764] J. R. Partington. The life and work of John Mayow (1641–1679). Part one. Isis, 47(3):217–230, September 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226889.

Cohen:1956:GRP

[765] I. Bernard Cohen. Galileo’s rejection of the possibility of velocity changing uniformly with respect to distance. Isis, 47(3):231–235, September 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226890.

Manes:1956:DXR

[766] George I. Manes. The discovery of x-ray. Isis, 47(3):236–238, September 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226891.

Levey:1956:SNC

[767] Martin Levey, Miroslav Krek, and Husni Haddad. Some notes on the chemical technology in an Eleventh Century Arabic work on bookbinding. Isis, 47(3):239–243, September 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226892.

Stahlman:1956:RCA

[768] William D. Stahlman. Raymond Claire Archibald (1875–1955). Isis, 47 (3):244–246, September 1956. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226893.

Zirkle:1956:EFC

[769] Conway Zirkle, John F. Fulton, I. E. Drabkin, Carl B. Boyer, I. Bernard Cohen, and Katharine Strelsky. Eighty-first critical bibliography of the history of science and its cultural influences (to 1 January 1956). Isis, 47 (3):247–360, September 1956. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226894.

Anonymous:1956:BMc

[770] Anonymous. Back matter. Isis, 47(3):361–366, September 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226895.

Anonymous:1956:FMd

[771] Anonymous. Front matter. Isis, 47(4):367–368, December 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226628.

Williams:1956:SEN

[772] L. Pearce Williams. Science, education and Napoleon I. Isis, 47(4): 369–382, December 1956. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226629.

Boyer:1956:RRA

[773] Carl B. Boyer. Refraction and the rainbow in antiquity. Isis, 47(4): 383–386, December 1956. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226630.

vanderWaerden:1956:TEA

[774] B. L. van der Waerden. Tables for the Egyptian and Alexandrian calendar. Isis, 47(4):387–390, December 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226631.

Thorndike:1956:NUS

[775] Lynn Thorndike. Notes upon some Medieval Latin astronomical, astrological and mathematical manuscripts at the Vatican: Part I. Isis, 47 (4):391–404, December 1956. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226632.

Partington:1956:LWJb

[776] J. R. Partington. The life and work of John Mayow (1641–1679): Part two. Isis, 47(4):405–417, December 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226633.

C:1956:NC

[777] I. B. C., Thomas H. Grainger, P. H. Brans, and L. P. Williams. Notes & correspondence. Isis, 47(4):418–423, December 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226634.

Hiebert:1956:BRB

[778] Erwin N. Hiebert. Book review: Kleine Weltgeschichte der Wissenschaft by Hans Joachim Störig. Isis, 47(4):424–426, December 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226635.

Greene:1956:BRB

[779] John C. Greene. Book review: Early American Science, Needs and Opportunities for Study by Whitfield J. Bell. Isis, 47(4):426–427, December 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226636.

Foote:1956:BRB

[780] George A. Foote. Book review: Belief and Unbelief since 1850 by H. G. Wood; Science and Christian Belief by C. A. Coulson. Isis, 47(4): 427–428, December 1956. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226637.

Bridgman:1956:BRB

[781] P. W. Bridgman. Book review: History of the Theories of Aether and Electricity. Volume II, The Modern Theories, 1900–1926 by Edmund Whittaker. Isis, 47(4):428–430, December 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226638.

Ihde:1956:BRBb

[782] Aaron J. Ihde. Book review: American Chemical Industry. Volume 1, Background and Beginnings: 1608–1910 by Williams Haynes; American Chemical Industry. Volume 5, The Decade of New Products: 1930–1939. Isis, 47(4):430–431, December 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226639.

Holton:1956:BRB

[783] Gerald Holton. Book review: Principles and Applications of Physics by Otto Blüh; Joseph Denison Elder. Isis, 47(4):431–433, December 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226640.

Zirkle:1956:BRBa

[784] Conway Zirkle. Book review: La Révolution Agricole by Michel Augé-Laribé. Isis, 47(4):433, December 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226641.

Zirkle:1956:BRBb

[785] Conway Zirkle. Book review: Salamanders and Other Wonders by Willy Ley. Isis, 47(4):433–434, December 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226642.

Zirkle:1956:BRBc

[786] Conway Zirkle. Book review: The Prevalence of People by Marston Bates; Hungry Generations: The Nineteenth Century Case against Malthusianism by Harold A. Boner; Standing Room Only. The Challenge of Over-Population by Karl Sax. Isis, 47(4):434–438, December 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226643.

Brozek:1956:BRBc

[787] Josef Brožek. Book review: Catholics in Psychology. A Historical Survey by Henryk Misiak; Virginia M. Staudt. Isis, 47(4):438–439, December 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226644.

Theodorides:1956:BRB

[788] J. Théodoridès. Book review: “L’histoire de la gale et le roman de l’Acare” by B. Dujardin. Isis, 47(4):439–441, December 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226645.

Sarton:1956:BRBc

[789] George Sarton. Book review: Mining and Copper Metallurgy in Stora Kopparberg until the Beginning of the Nineteenth Century by Sten Lindroth. Isis, 47(4):441–443, December 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226646.

Wohl:1956:BRB

[790] Sonia S. Wohl. Book review: Handbook of the Greek Collection, Metropolitan Museum of Art by Gisela M. A. Richter; Catalogue of Greek Sculptures in the Metropolitan Museum of Art by Gisela M. A. Richter. Isis, 47(4):443–444, December 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226647.

Pagel:1956:BRBb

[791] Walter Pagel. Book review: The Surgery of Theodoric, ca. A.D. 1267 by Eldridge Campbell; James Colton. Isis, 47(4):444–445, December 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226648.

OMalley:1956:BRBa

[792] C. D. O’Malley. Book review: Leonardo da Vinci. Das anatomische Werk by Sigrid Esche. Isis, 47(4):445–447, December 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226649.

Cohen:1956:BRBa

[793] I. Bernard Cohen. Book review: The Principal Works of Simon Stevin. Vol. I: General Introduction — Mechanics by Simon Stevin; E. J. Dijksterhuis; C. Dikshoorn. Isis, 47(4):447–448, December 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226650.

Cohen:1956:BRBb

[794] I. Bernard Cohen. Book review: Traité d’optique by Isaac Newton. Isis, 47(4):448–449, December 1956. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226651.

Truesdell:1956:BRB

[795] C. Truesdell. Book review: La mécanique au XVIIe siècle (des antécédents scolastiques à la pensée classique) by Rene Dugas. Isis, 47 (4):449–452, December 1956. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226652.

Gillispie:1956:BRB

[796] Charles C. Gillispie. Book review: Entretiens sur la pluralité des mondes. Digressions sur les anciens et les modernes by Bernard le Bovier de Fontenelle; Robert Shackleton. Isis, 47(4):452–453, December 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226653.

OMalley:1956:BRBb

[797] C. D. O’Malley. Book review: Commentary on the Effect of Electricity on Muscular Motion. A Translation of De viribus electricitatis in motu musculari commentarius by Luigi Galvani; Robert Montraville Green. Isis, 47(4):454–455, December 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226654.

Kuhn:1956:AD

[798] Thomas S. Kuhn and Stanley M. Loomis. Administrative documents. Isis, 47(4):455–460, December 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226655.

Anonymous:1956:VI

[799] Anonymous. Volume information. Isis, 47(4):461–466, December 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226627.

Anonymous:1956:BMd

[800] Anonymous. Back matter. Isis, 47(4):??, December 1956. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226656.

Anonymous:1957:FMa

[801] Anonymous. Front matter. Isis, 48(1):1–2, March 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/226896.

DeMott:1957:SVM

[802] Benjamin DeMott. Science versus mnemonics: Notes on John Ray and on John Wilkins’ essay toward a real character, and a philosophical language. Isis, 48(1):3–12, March 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226897.

Chapin:1957:AAN

[803] Seymour L. Chapin. The astronomical activities of Nicolas Claude Fabri de Peiresc. Isis, 48(1):13–29, March 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226898.

Baron:1957:NVA

[804] Roger Baron. Note sur les variations au XIIe Siècle de la triade Géométrique altimetria, planimetria, cosmimetria. (French) [note on the variations in the 12th Century of the geometric triad, altimetria, planimetria, cosmimetria]. Isis, 48(1):30–32, March 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226899.

Stauffer:1957:SEB

[805] Robert C. Stauffer. Speculation and experiment in the background of Oersted’s discovery of electromagnetism. Isis, 48(1):33–50, March 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226900.

Thorndike:1957:EFF

[806] Lynn Thorndike. Eclipses in the Fourteenth and Fifteenth Centuries. Isis, 48(1):51–57, March 1957. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226901.

Zirkle:1957:BFT

[807] Conway Zirkle. Benjamin Franklin, Thomas Malthus and the United States Census. Isis, 48(1):58–62, March 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226902.

Fulton:1957:NC

[808] John F. Fulton, Jean F. Leroy, Stillman Drake, Edward Rosen, George A. Summent, and John B. Blake. Notes & correspondence. Isis, 48(1):63– 70, March 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226903.

Zirkle:1957:BRBa

[809] Conway Zirkle. Book review: The Facts of Life by C. D. Darlington. Isis, 48(1):71–73, March 1957. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226904.

Saunders:1957:BRBa

[810] J. B. DeC. M. Saunders. Book review: A Short History of Medicine by Erwin H. Ackerknecht; The Story of Medicine by Kenneth Walker. Isis, 48(1):73–76, March 1957. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226905.

White:1957:BRB

[811] Lynn White, Jr. Book review: Studies in Ancient Technology by Robert James Forbes. Isis, 48(1):77, March 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226906.

Willey:1957:BRB

[812] Gordon R. Willey. Book review: Highway of the Sun by Victor von Hagen. Isis, 48(1):77–79, March 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226907.

Saunders:1957:BRBb

[813] J. B. DeC. M. Saunders. Book review: Clio en Epidaure ou la Médecine et l’Humanisme chez les Anciens by G. Barraud. Isis, 48(1):79–80, March 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226908.

Clagett:1957:BRBa

[814] Marshall Clagett. Book review: The Classical Heritage & Its Beneficiaries by R. R. Bolgar. Isis, 48(1):80–81, March 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/226909.

Aaboe:1957:BRB

[815] Asger Aaboe. Book review: The Bequest of the Greeks by Tobias Dantzig. Isis, 48(1):81–83, March 1957. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226910.

Moody:1957:BRBa

[816] Ernest A. Moody. Book review: Opera omnia, Tomus XII: Liber de natura et origine animae by Albertus Magnus; Liber de principiis motus processivi by Bernhard Geyer; Quaestiones super de animalibus by Bernhard Geyer; Ephrem Filthaut. Isis, 48(1):83–84, March 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226911.

Maier:1957:BRB

[817] Anneliese Maier. Book review: Thomas of Bradwardine, His Tractatus de Proportionibus, Its Significance for the Development of Mathematical Physics by H. Lamar Crosby. Isis, 48(1):84–87, March 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226912.

Winter:1957:BRB

[818] H. J. J. Winter. Book review: Averroës, Tahafut al-Tahafut (The Incoherence of the Incoherence) by Averroës; Simon van den Bergh. Isis, 48(1):87–88, March 1957. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226913.

Jammer:1957:BRB

[819] Max Jammer. Book review: Cosmografía de un Judío Romano del siglo XVII by José M. Millás Vallicrosa; David Romano. Isis, 48(1):88–90, March 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226914.

Smith:1957:BRB

[820] Cyril Stanley Smith. Book review: Elizabethan Copper. The History of the Company of Mines Royal, 1568–1605 by M. B. Donald. Isis, 48(1): 90–91, March 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226915.

Kuhn:1957:BRB

[821] Thomas S. Kuhn. Book review: A Documentary History of the Problem of Fall from Kepler to Newton by Alexandre Koyré. Isis, 48(1):91–93, March 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226916.

Urdang:1957:BRB

[822] George Urdang. Book review: Geschichte der Medizin. Die historische Entwicklung der Heilkunde und des ärztlichen Lebens. II. Band: 2. Hälfte: Die Medizin vom Beginn der zellularpathologie bis zu den Anfängen der modernen Konstitutionslehre (etwa 1858–1900) by Paul Diepgen. Isis, 48(1):93, March 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226917.

Heathcote:1957:BRB

[823] N. H. de V. Heathcote. Book review: Thomas Young, Natural Philosopher, 1773–1829 by Alexander Wood; Frank Oldham; Charles E. Raven. Isis, 48(1):93–96, March 1957. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226918.

Schoenwald:1957:BRB

[824] Richard L. Schoenwald. Book review: The Life and Work of Sigmund Freud. Volume 2: Years of Maturity, 1901–1919 by Ernest Jones. Isis, 48 (1):97, March 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226919.

Zirkle:1957:BRBb

[825] Conway Zirkle. Book review: One Man in His Time by N. M. Borodin. Isis, 48(1):97–99, March 1957. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226920.

Anonymous:1957:BMa

[826] Anonymous. Back matter. Isis, 48(1):100, March 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/226921.

Anonymous:1957:FMb

[827] Anonymous. Front matter. Isis, 48(2):101, June 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227024.

Mazzeo:1957:NJD

[828] Joseph A. Mazzeo. Notes on John Donne’s alchemical imagery. Isis, 48(2):103–123, June 1957. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227025.

Guerlac:1957:JBFa

[829] Henry Guerlac. Joseph Black and fixed air a bicentenary retrospective, with some new or little known material. Isis, 48(2):124–151, June 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227026.

Gillispie:1957:DLP

[830] Charles C. Gillispie. The discovery of the Leblanc process. Isis, 48(2): 152–170, June 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227027.

Rosen:1957:WME

[831] Edward Rosen. Was Maurolico’s essay on the nova of 1572 printed? Isis, 48(2):171–175, June 1957. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227028.

Cohen:1957:EIC

[832] I. Bernard Cohen. The Eighth International Congress of the History of Science Florence–Milan, 3–9 September 1956. Isis, 48(2):176–181, June 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227029. See erratum [876].

Clagett:1957:NC

[833] Marshall Clagett, Gerald Gruman, Maurice Daumas, and Harry Woolf. Notes & correspondence. Isis, 48(2):182–188, June 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227030.

Cohen:1957:ESC

[834] I. Bernard Cohen and Katharine Strelsky. Eighty-second critical bibliography of the history of science and its cultural influences (to 1 January 1957) (endorsed by the history of science division of the international union of the history and philosophy of science). Isis, 48(2):189–280, June 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227031.

Anonymous:1957:BMb

[835] Anonymous. Back matter. Isis, 48(2):??, June 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227032.

Anonymous:1957:FMc

[836] Anonymous. Front matter. Isis, 48(3):281–282, September 1957. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226463.

Stimson:1957:DSH

[837] Dorothy Stimson. Dr. Sarton and the History of Science Society. Isis, 48 (3):283–284, September 1957. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226464.

Sarton:1957:C

[838] May Sarton. A celebration. Isis, 48(3):285, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226465.

Cohen:1957:GS

[839] I. Bernard Cohen. George Sarton. Isis, 48(3):286–300, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226466.

Conant:1957:GSH

[840] James B. Conant. George Sarton and Harvard University. Isis, 48(3): 301–305, September 1957. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226467.

Singer:1957:GSH

[841] Dorothea Singer and Charles Singer. George Sarton and the history of science. Isis, 48(3):306–310, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226468.

Fulton:1957:GSH

[842] John F. Fulton. George Sarton and the history of medicine. Isis, 48 (3):311–314, September 1957. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226469.

Vallicrosa:1957:GSH

[843] José M. Millás Vallicrosa. George Sarton y la historia de la ciencia oriental. Isis, 48(3):315–319, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226470.

Clagett:1957:GSH

[844] Marshall Clagett. George Sarton: Historian of medieval science. Isis, 48 (3):320–322, September 1957. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226471.

Thorndike:1957:SLG

[845] Lynn Thorndike. Some letters of George Sarton (1922–1955). Isis, 48 (3):323–334, September 1957. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226472.

Malik:1957:DSS

[846] Charles Habib Mâlik. Dr. Sarton’s study of Arabic. Isis, 48(3):335, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226473.

Strelsky:1957:BPG

[847] Katharine Strelsky. Bibliography of the publications of George Sarton. Isis, 48(3):336–350, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226474.

Anonymous:1957:DPH

[848] Anonymous. Dissertations in progress in the history of science. Isis, 48(3):351, September 1957. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226475.

Anonymous:1957:DHS

[849] Anonymous. Doctorates in the history of science. Isis, 48(3):351–352, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226476.

Hahn:1957:BRB

[850] Roger Hahn. Book review: Causalités et accidents de la découverte scientifique. Illustratration de quelques étapes caractéristiques de l’évolution des sciences by René Taton. Isis, 48(3):352–353, September 1957. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226477.

Cohen:1957:BRBa

[851] I. Bernard Cohen. Book review: Bibliotheca Walleriana by Erik Waller; Hans Sallander. Isis, 48(3):353–355, September 1957. CODEN ISISA4.

ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226478.

Lenzen:1957:BRB

[852] V. F. Lenzen. Book review: De Thalès à Bergson. Introduction historique à la philosophie européenne by Philippe Devaux. Isis, 48(3):355–356, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226479.

Boyer:1957:BRBa

[853] Carl B. Boyer. Book review: Die geistige Arbeit by Andreas Speiser. Isis, 48(3):356–357, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226480.

Cohen:1957:BRBb

[854] I. Bernard Cohen. Book review: The History of the Telescope by Henry C. King. Isis, 48(3):357–358, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226481.

Schneer:1957:BRBa

[855] Cecil J. Schneer. Book review: Geschichte der Geologie und des geologischen Weltbildes by Carl Christoph Beringer. Isis, 48(3):358–359, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226482.

vanderWaerden:1957:BRB

[856] B. L. van der Waerden. Book review: Greek Science in Antiquity by Marshall Clagett. Isis, 48(3):359–360, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226483.

DeSantillana:1957:BRB

[857] Giorgio De Santillana. Book review: From Platonism to Neoplatonism by Philip Merlan. Isis, 48(3):360–362, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226484.

Saunders:1957:BRBc

[858] J. B. deC. M. Saunders. Book review: On Anatomical Procedures. De Anatomicis Administrationibus by Galen; Charles Singer. Isis, 48(3):

362–364, September 1957. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226485.

Vallicrosa:1957:BRB

[859] J. M. Millás Vallicrosa and Brunilda L. Debus. Book review: Suwaru’l-Kawakib by Abu’l-Husayn ‘Abdu’r-Rahman As.-S.ufi. Isis, 48(3):364–365, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226486.

Ware:1957:BRB

[860] James R. Ware. Book review: Science and Civilisation in China. Volume II, History of Scientific Thought by Joseph Needham. Isis, 48(3):365–367, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226487.

Sisco:1957:BRBa

[861] Anneliese Sisco. Book review: Lazarus Ercker by Paul R. Beierlein. Isis, 48(3):367–368, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226488.

Saunders:1957:BRBd

[862] J. B. DeC. M. Saunders. Book review: Opuscula selecta neerlandicorum de arte medica by Volcher Coiter. Isis, 48(3):368–370, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226489.

Sisco:1957:BRBb

[863] Anneliese Sisco. Book review: Ulrich Rülein von Calw und sein Bergbüchlein by Wilhelm Pieper. Isis, 48(3):370–371, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226490.

Brown:1957:BRB

[864] Lloyd Arnold Brown. Book review: The Roanoke Voyages, 1584–1590. Documents to Illustrate the English Voyages to North America under the Patent Granted to Walter Raleigh in 1584 by David Beers Quinn. Isis, 48 (3):371–373, September 1957. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226491.

Johnson:1957:BRBa

[865] Francis R. Johnson. Book review: The Appreciation of Ancient and Medieval Science during the Renaissance (1450–1600) by George Sarton.

Isis, 48(3):373–375, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226492.

Hellman:1957:BRB

[866] C. Doris Hellman. Book review: Six Wings: Men of Science in the Renaissance by George Sarton. Isis, 48(3):375–377, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226493.

Johnson:1957:BRBb

[867] Francis R. Johnson. Book review: The Mathematical Practitioners of Tudor & Stuart England by E. G. R. Taylor. Isis, 48(3):377–378, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226494.

deSantillana:1957:BRBa

[868] Giorgio de Santillana. Book review: Discoveries and Opinions of Galileo by Stillman Drake. Isis, 48(3):378–379, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/226495.

Hiebert:1957:BRBa

[869] Erwin N. Hiebert. Book review: Johann Rudolph Glauber, 1604–1670, Leben und Werk by Kurt F. Gugel. Isis, 48(3):379–380, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226496.

Stimson:1957:BRB

[870] Dorothy Stimson. Book review: Sir Kenelm Digby: The Ornament of England (1603–1665) by R. T. Petersson. Isis, 48(3):381–382, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226497.

Woolf:1957:BRB

[871] Harry Woolf. Book review: The Journals of Captain Cook on His Voyages of Discovery. Volume I, The Voyage of the Endeavour by J. C. Beaglehole; Charts and Views Drawn by Cook and His Officers and Reproduced from the Original Manuscripts by R. A. Skelton. Isis, 48(3): 382–384, September 1957. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226498.

Brozek:1957:BRBa

[872] Josef Brožek. Book review: Georgius Prochaska (1749–1820); De Functionibus Systematis Nervosi Commentatio by Georgius Prochaska; Mojmir Petran; Ernest Gutmann; Med. Dr. Jirí Procháska, 1749–1820: Zivot, Dílo, Doba by Vladislav Kruta. Isis, 48(3):384–385, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226499.

Holton:1957:BRB

[873] Gerald Holton. Book review: Ludwig Boltzmann, Mensch, Physiker, Philosoph by Engelbert Broda. Isis, 48(3):385–386, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226500.

Eyles:1957:BRB

[874] V. A. Eyles. Book review: Illustrations of the Huttonian Theory of the Earth by John Playfair. Isis, 48(3):386–388, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226501.

Brozek:1957:BRBb

[875] Josef Brožek. Book review: Opera Omnia. Tomus VI. Dissertationes Discipulorum by Jan Evangelista Purkyne; Vladislav Kruta; Zdenek Hornof. Isis, 48(3):388–389, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226502.

Anonymous:1957:EEI

[876] Anonymous. Erratum: The Eighth International Congress of the History of Science Florence–Milan, 3–9 September 1956. Isis, 48(3):389, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226503. See [832].

Anonymous:1957:BMc

[877] Anonymous. Back matter. Isis, 48(3):390, September 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226504.

Anonymous:1957:FMd

[878] Anonymous. Front matter. Isis, 48(4):391–392, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227510.

Drake:1957:GGS

[879] Stillman Drake. Galileo gleanings i: Some unpublished anecdotes of Galileo. Isis, 48(4):393–397, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227511.

Gillispie:1957:NHI

[880] Charles C. Gillispie. The natural history of industry. Isis, 48(4):398–407, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227512.

Schofield:1957:IOS

[881] Robert E. Schofield. The industrial orientation of science in the Lunar Society of Birmingham. Isis, 48(4):408–415, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227513.

Duveen:1957:LSB

[882] Denis I. Duveen and Roger Hahn. Laplace’s succession to Bézout’s post of Examinateur des Elèves de l’Artillerie. Isis, 48(4):416–427, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227514.

Roberts:1957:SLT

[883] Victor Roberts. The solar and lunar theory of ibn ash–Shāt.ir: A precopernican Copernican model. Isis, 48(4):428–432, December 1957. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227515.

Guerlac:1957:JBFb

[884] Henry Guerlac. Joseph Black and fixed air: Part II. Isis, 48(4):433–456, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227516.

Brans:1957:NC

[885] P. H. Brans, Henry Guerlac, Lynn Thorndike, Rufus Suter, Bernard Dulsey, E. R. N. Grigg, V. F. Payne, and Marshall Clagett. Notes & correspondence. Isis, 48(4):457–470, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227517.

Hindle:1957:BRB

[886] Brooke Hindle. Book review: Science in the Federal Government. A History of Policies and Activities to 1940 by A. Hunter Dupree. Isis, 48 (4):470–471, December 1957. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227518.

Williams:1957:BRB

[887] L. Pearce Williams. Book review: Science in History by J. D. Bernal. Isis, 48(4):471–473, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227519.

Hodgen:1957:BRB

[888] Margaret T. Hodgen. Book review: The Acceptance of Histories: Toward a Perspective for Social Science by Kenneth E. Bock. Isis, 48(4):473–475, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227520.

Zirkle:1957:BRBc

[889] Conway Zirkle. Book review: The Biological Basis of Human Freedom by Theodosius Dobzhansky. Isis, 48(4):475–476, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227521.

Boyer:1957:BRBb

[890] Carl B. Boyer. Book review: The World of Mathematics by James R. Newman. Isis, 48(4):476–479, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227522.

Hiebert:1957:BRBb

[891] Erwin N. Hiebert. Book review: The Language of Modern Physics by Ernest H. Hutton. Isis, 48(4):479, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227523.

Schneer:1957:BRBb

[892] Cecil J. Schneer. Book review: The Earth We Live On. The Story of Geological Discovery by Ruth Moore. Isis, 48(4):479–481, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227524.

Gruman:1957:BRB

[893] Gerald Gruman. Book review: The Biology of Senescence by Alex Comfort. Isis, 48(4):481–482, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227525.

Cottrell:1957:BRB

[894] G. W. Cottrell, Jr. Book review: They Tell of Birds: Chaucer, Spenser, Milton, Drayton by Thomas P. Harrison. Isis, 48(4):482–483, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227526.

Leikind:1957:BRB

[895] Morris C. Leikind. Book review: Man’s Mastery of Malaria by Paul F. Russell. Isis, 48(4):483–484, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227527.

Condit:1957:BRB

[896] Carl W. Condit. Book review: Engineering in History by Richard Shelton Kirby; Sidney Withington; Arthur Burr Darling; Frederick Gridley Kilgour; History of American Technology by John W. Oliver. Isis, 48 (4):484–487, December 1957. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227528.

Stahl:1957:BRB

[897] William H. Stahl. Book review: Theophrastus on Stones by Earle R. Caley; John C. Richards. Isis, 48(4):487–488, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227529.

Moody:1957:BRBb

[898] Ernest A. Moody. Book review: William Heytesbury: Medieval Logic and the Rise of Mathematical Physics by Curtis Wilson. Isis, 48(4):488–489, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227530.

Vernet:1957:BRB

[899] Juan Vernet. Book review: Al-Qanunu’ l-Mas ‘udi by Abu Rayh.an Muh.ammad al-Biruni. Isis, 48(4):489–490, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227531.

Clagett:1957:BRBb

[900] Marshall Clagett. Book review: Commentarivm medivm in Aristotelis de generatione et corrvptione libros by Averrois Cordvbensis. Isis, 48 (4):490–491, December 1957. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227532.

Price:1957:BRB

[901] Derek J. Price. Book review: Islamic Astrolabes and Their Works by L. A. Mayer. Isis, 48(4):491–492, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227533.

Schneer:1957:BRBc

[902] Cecil J. Schneer. Book review: Robert Hooke by Margaret ’Espinasse. Isis, 48(4):493–494, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227534.

Bush:1957:BRB

[903] Douglas Bush. Book review: Milton and Science by Kester Svendsen. Isis, 48(4):494–495, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227535.

Hall:1957:BRB

[904] A. R. Hall. Book review: Franklin and Newton. An Inquiry into Speculative Newtonian Experimental Science and Franklin’s Work in Electricity as an Example Thereof by I. Bernard Cohen. Isis, 48(4):495–498, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227536.

Schofield:1957:BRB

[905] Robert E. Schofield. Book review: Early Electrical Machines by Bern Dibner. Isis, 48(4):498–499, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227537.

Gillispie:1957:BRB

[906] Charles C. Gillispie. Book review: Réaumur’s Memoirs on Steel and Iron by René Antoine Ferchault Réaumur; Anneliese Grunhaldt Sisco. Isis, 48 (4):499–500, December 1957. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227538.

Leake:1957:BRB

[907] Chauncey D. Leake. Book review: Human Generation by Arthur William Meyer. Isis, 48(4):500–501, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227539.

Anonymous:1957:VI

[908] Anonymous. Volume information. Isis, 48(4):502–506, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227509.

Anonymous:1957:BMd

[909] Anonymous. Back matter. Isis, 48(4):??, December 1957. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227540.

Anonymous:1958:FMa

[910] Anonymous. Front matter. Isis, 49(1):1–2, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/226598.

Romer:1958:TTR

[911] Alfred Romer. The transformation theory of radioactivity. Isis, 49(1): 3–12, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226599.

Greene:1958:SPA

[912] John C. Greene. Science and the public in the age of Jefferson. Isis, 49(1): 13–25, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226600.

Drake:1958:GGIa

[913] Stillman Drake. Galileo gleanings II a kind word for salusbury. Isis, 49 (1):26–33, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/226601.

Thorndike:1958:NUS

[914] Lynn Thorndike. Notes upon some Medieval Latin astronomical, astrological and mathematical manuscripts at the Vatican: Part II. Isis, 49(1): 34–49, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226602.

Multhauf:1958:BMC

[915] Robert Multhauf. The beginning of mineralogical chemistry. Isis, 49(1): 50–53, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226603.

Truesdell:1958:NBE

[916] C. Truesdell. The new Bernoulli edition. Isis, 49(1):54–62, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226604.

George:1958:CWL

[917] Katherine George. The civilized West looks at primitive Africa: 1400– 1800 a study in ethnocentrism. Isis, 49(1):62–72, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226605.

Duveen:1958:NC

[918] Denis I. Duveen, Herbert S. Klickstein, P. H. Brans, G. Polvani, and Ivolino de Vasconcellos. Notes & correspondence. Isis, 49(1):73–76, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226606.

Waller:1958:THS

[919] G. M. Waller, John C. Greene, Robert E. Schofield, A. G. Unklesbay, and Harry Woolf. Teaching the history of science. Isis, 49(1):77–78, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226607.

Figourovski:1958:AHS

[920] N. A. Figourovski, V. P. Zoubov, and Otakar Matousek. Academies and the history of science. Isis, 49(1):78–80, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226608.

Anonymous:1958:WPH

[921] Anonymous. Work in progress in the history of science. Isis, 49(1):80– 81, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226609.

Tuttle:1958:SHS

[922] Preston Tuttle and Carl B. Boyer. Sections of the History of Science Society. Isis, 49(1):81–84, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226610.

Pagel:1958:BRBa

[923] Walter Pagel. Book review: Beiträge zur Geschichte der Technologie und der Alchemie by W. Ganzenmüller. Isis, 49(1):84–86, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226611.

Ihde:1958:BRBa

[924] Aaron J. Ihde. Book review: Discovery of the Elements by Mary Elvira Weeks. Isis, 49(1):86–87, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226612.

Price:1958:BRB

[925] Derek J. Price. Book review: Deutsche und niederländische astronomische Instrumente des 11.-18. Jahrhunderts by Ernst Zinner. Isis, 49(1): 87–88, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226613.

Hiebert:1958:BRB

[926] Erwin N. Hiebert. Book review: The Historical Background of Chemistry by Henry M. Leicester. Isis, 49(1):88–89, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226614.

Smith:1958:BRB

[927] Cyril Stanley Smith. Book review: A History of Technology. Volume II, The Mediterranean Civilizations and the Middle Ages, c. 700 B.C. to c. A.D. 1500 by Charles Singer; E. J. Holmyard; A. R. Hall; Trevor I. Williams; E. Jaffé; Nan Clow; R. H. G. Thomson. Isis, 49(1):89– 90, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226615.

MacKinney:1958:BRB

[928] Loren C. MacKinney. Book review: I codici di medicina del periodo presalernitano (secoli IX, X e XI) by Augusto Beccaria. Isis, 49(1):90– 91, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226616.

Clagett:1958:BRBa

[929] Marshall Clagett. Book review: Archimedes by E. J. Dijksterhuis. Isis, 49(1):91–92, March 1958. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226617.

Ronchi:1958:BRB

[930] Vasco Ronchi. Book review: L’optique de Claude Ptolémée dans la version latine d’après l’arabe de l’émire Eugène de Sicile by Albert Lejeune. Isis, 49(1):92–93, March 1958. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226618.

Levy:1958:BRB

[931] Martin Levy. Book review: Grundsätze altindischer Medizin. (Acta Historica Scientiarum Naturalium et Medicinalium) by Reinhold F. G. Müller. Isis, 49(1):93–94, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226619.

Ware:1958:BRB

[932] James R. Ware. Book review: Fifty Years of Chinese Philosophy, 1898– 1950 by O. Brière; Laurence G. Thompson. Isis, 49(1):94, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226620.

Brown:1958:BRB

[933] Harcourt Brown. Book review: Etudes Rabelaisiennes. Isis, 49(1):94– 96, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226621.

Leicester:1958:BRB

[934] Henry M. Leicester. Book review: Lavoisier by Ya. G. Dorfman. Isis, 49 (1):97, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226622.

Rosen:1958:BRBa

[935] George Rosen. Book review: Arbeitsmedizin im 18. Jahrhundert: Werksarzt und Arbeiter im Quecksilberbergwerk Idria by Erna Lesky. Isis, 49 (1):98, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226623.

Viets:1958:BRB

[936] Henry R. Viets. Book review: New England Hospitals, 1790–1833 by Leonard K. Eaton. Isis, 49(1):98–99, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226624.

Boas:1958:ADa

[937] Marie Boas. Administrative documents. Isis, 49(1):99–104, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226625.

Anonymous:1958:BMa

[938] Anonymous. Back matter. Isis, 49(1):??, March 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226626.

Anonymous:1958:FMb

[939] Anonymous. Front matter. Isis, 49(2):105, June 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226922.

Anonymous:1958:AGS

[940] Anonymous. Award of the George Sarton Medal to Lynn Thorndike. Isis, 49(2):107–108, June 1958. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226923.

Wyckoff:1958:AMO

[941] Dorothy Wyckoff. Albertus Magnus on Ore deposits. Isis, 49(2):109– 122, June 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226924.

Beer:1958:CTD

[942] John J. Beer. Coal tar dye manufacture and the origins of the modern industrial research laboratory. Isis, 49(2):123–131, June 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226925.

Kuhn:1958:CTA

[943] Thomas S. Kuhn. The caloric theory of adiabatic compression. Isis, 49(2):132–140, June 1958. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226926.

Boyer:1958:TRH

[944] Carl B. Boyer. The tertiary rainbow: An historical account. Isis, 49(2): 141–154, June 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226927.

Drake:1958:GGIb

[945] Stillman Drake. Galileo gleanings III: A kind word for Sizzi. Isis, 49(2): 155–165, June 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226928.

Levey:1958:AAM

[946] Martin Levey. Alum in Ancient Mesopotamian technology. Isis, 49(2): 166–169, June 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226929.

Rosen:1958:HS

[947] George Rosen. Henry E. Sigerist (1891–1957). Isis, 49(2):170–171, June 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226930.

Drake:1958:QA

[948] Stillman Drake and I. Bernard Cohen. Queries & answers. Isis, 49(2): 172, June 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226931.

Dupree:1958:THS

[949] A. Hunter Dupree and Thomas S. Kuhn. Teaching the history of science. Isis, 49(2):172–173, June 1958. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226932.

deSantillana:1958:NC

[950] Giorgio de Santillana, J. M. Keynes, Herman H. Henkle, E. S. Kennedy, and C. Doris Hellman. Notes & correspondence. Isis, 49(2):173–178, June 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226933.

Cohen:1958:ETC

[951] I. Bernard Cohen and Katharine Strelsky. Eighty-third critical bibliography of the history of science and its cultural influences (to 1 January 1958). Isis, 49(2):179–296, June 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226934.

Boas:1958:ADb

[952] Marie Boas and William D. Stahlman. Administrative documents. Isis, 49(2):297–300, June 1958. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226935.

Anonymous:1958:BMb

[953] Anonymous. Back matter. Isis, 49(2):301–304, June 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226936.

Anonymous:1958:FMc

[954] Anonymous. Front matter. Isis, 49(3):305–306, September 1958. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226937.

Reingold:1958:SCW

[955] Nathan Reingold. Science in the Civil War. The permanent commission of the Navy Department. Isis, 49(3):307–318, September 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226938.

Rosen:1958:GMA

[956] Edward Rosen. Galileo’s misstatements about Copernicus. Isis, 49(3): 319–330, September 1958. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226939.

Sambursky:1958:SRE

[957] S. Sambursky. On some references to experience in stoic physics. Isis, 49 (3):331–335, September 1958. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226940.

Boyer:1958:JG

[958] Carl B. Boyer. Jekuthiel Ginsburg (1889–1957). Isis, 49(3):335–336, September 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226941.

Levey:1958:GSS

[959] Martin Levey. Gypsum, salt and soda in Ancient Mesopotamian chemical technology. Isis, 49(3):336–342, September 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226942.

Hall:1958:NC

[960] A. R. Hall, I. B. C., Stillman Drake, Denis I. Duveen, and Herbert S. Klickstein. Notes & correspondence. Isis, 49(3):342–349, September 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226943.

Boyer:1958:BRBa

[961] Carl B. Boyer. Book review: The Study of the History of Mathematics, and, The Study of the History of Science by George Sarton. Isis, 49(3): 350, September 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/226944.

Boyer:1958:BRBb

[962] Carl B. Boyer. Book review: Geschichte der Mathematik; Vol. II, Von Fermat und Descartes bis zum Ausbau der neuen Methoden by Joseph E. Hofmann; Geschichte der Mathematik; Vol.III, Von den Auseinandersetzungen um den Calculus bis zur franzosischen Revolution. Isis, 49 (3):350–352, September 1958. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226945.

Ihde:1958:BRBb

[963] Aaron Ihde. Book review: A History of Industrial Chemistry by F. Sherwood Taylor. Isis, 49(3):352, September 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226946.

Morison:1958:BRB

[964] S. E. Morison. Book review: The Haven-Finding Art. A History of Navigation from Odysseus to Captain Cook by E. G. R. Taylor. Isis, 49(3): 352–353, September 1958. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226947.

Saunders:1958:BRB

[965] J. B. deC. M. Saunders. Book review: Essai historique sur l’Hôpital Saint-Eloi de Montpellier (1183–1950) by Louis Dulieu. Isis, 49(3):353– 354, September 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/226948.

Poynter:1958:BRB

[966] F. N. L. Poynter. Book review: Medicine and the Navy, 1200–1900 by J. J. Keevil. Isis, 49(3):354–355, September 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226949.

Abetti:1958:BRBa

[967] Giorgio Abetti. Book review: Astronomical Cuneiform Texts. Babylonian Ephemerides of the Seleucid Period for the Motion of the Sun, Moon, and the Planets by O. Neugebauer. Isis, 49(3):355–356, September 1958.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226950.

Gillispie:1958:BRB

[968] Charles C. Gillispie. Book review: The Physical World of the Greeks by S. Sambursky. Isis, 49(3):356–358, September 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226951.

Clagett:1958:BRBb

[969] Marshall Clagett. Book review: The Philosophy of the Church Fathers. Volume I, Faith, Trinity, Incarnation by Harry Austryn Wolfson. Isis, 49 (3):358–359, September 1958. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226952.

Clagett:1958:BRBc

[970] Marshall Clagett. Book review: Le Système du Monde. Vol.VI by Pierre Duhem; Le Systeme du Monde. Vol.VII. Isis, 49(3):359–362, September 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226953.

Clagett:1958:BRBd

[971] Marshall Clagett. Book review: Bradwardine and the Pelagians. A Study of His De causa Dei and Its Opponents by Gordon Leff. Isis, 49(3):362, September 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226954.

Filliozat:1958:BRB

[972] J. Filliozat. Book review: History of Chemistry in Ancient and Medieval India, Incorporating the History of Hindu Chemistry by Prafulla Chandra Ray. Isis, 49(3):362–363, September 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226955.

Boas:1958:BRB

[973] Marie Boas. Book review: From the Closed World to the Infinite Universe by Alexandre Koyré. Isis, 49(3):363–366, September 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226956.

Woolf:1958:BRB

[974] Harry Woolf. Book review: The Copernican Revolution. Planetary Astronomy in the Development of Western Thought by Thomas S. Kuhn.

Isis, 49(3):366–367, September 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226957.

Multhauf:1958:BRB

[975] Robert Multhauf. Book review: The Admirable Discourses by Bernard Palissy; Aurele la Rocque. Isis, 49(3):367–368, September 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226958.

Sisco:1958:BRBa

[976] Anneliese Sisco. Book review: Georg Agricola und seine Zeit by Helmut Wilsdorf; Bermannus oder über den Bergbau. Ein Dialog by Georg Agricola; Schriften zur Geologie und Mineralogie I: Epistula ad Meurerum; De ortu et causis subterraneorum libri V; De natura eorum quae effluunt ex terra libri IV by Georgius Agricola. Isis, 49(3):368–369, September 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226959.

Sisco:1958:BRBb

[977] Anneliese Sisco. Book review: Georgius Agricola, 1494–1555, zu seinem 400. Todestag, 21. November 1955 by Rolf Wendler. Isis, 49(3):369–370, September 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226960.

Abetti:1958:BRBb

[978] Giorgio Abetti. Book review: Galileo Galilei by L. Geymonat. Isis, 49 (3):370–371, September 1958. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226961.

Anonymous:1958:BMc

[979] Anonymous. Back matter. Isis, 49(3):372, September 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226962.

Anonymous:1958:FMd

[980] Anonymous. Front matter. Isis, 49(4):373–374, December 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227280.

Woodbury:1958:FEG

[981] Robert S. Woodbury. The first epicycloidal gear teeth. Isis, 49(4):375– 377, December 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227281.

Boyer:1958:TRM

[982] Carl B. Boyer. The theory of the rainbow: Medieval triumph and failure. Isis, 49(4):378–390, December 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227282.

Patrides:1958:NAC

[983] C. A. Patrides. The numerological approach to cosmic order during the English Renaissance. Isis, 49(4):391–397, December 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227283.

Thorndike:1958:C

[984] Lynn Thorndike. De complexionibus. Isis, 49(4):398–408, December 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227284.

Drake:1958:GGIc

[985] Stillman Drake. Galileo gleanings IV bibliographical notes. Isis, 49(4): 409–413, December 1958. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227285.

Chipman:1958:MLS

[986] Robert A. Chipman. The manuscript letters of Stephen Gray, F.R.S. (1666/7–1736). Isis, 49(4):414–433, December 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227286.

Cohen:1958:WPH

[987] I. Bernard Cohen and Marie Thérèse d’Alverny. Work in progress in the history of science. Isis, 49(4):434–438, December 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227287.

Neville:1958:QA

[988] Roy G. Neville and William A. Cole. Queries & answers. Isis, 49(4): 438–439, December 1958. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227288.

Agnew:1958:NC

[989] L. R. C. Agnew, Alexandre Koyré, V. Ronchi, J. M. Millás Vallicrosa, and A. Koyré. Notes & correspondence. Isis, 49(4):439–445, December 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227289.

Sayili:1958:BRB

[990] Aydin Sayili. Book review: Histoire Générale des Sciences, publiée sous la direction de René Taton, Tome 1. La Science Antique et Médiévale (des Origines à 1450) by Réne Taton; R. Arnaldez; J. Beaujeu; G. Beaujouan; R. Bloch; L. Bourgey; P. Dupont-Sommer; J. Filliozat; R. Furon; A. Haudricourt; J. Itard; R. Labat; G. Lefebvre; L. Massignon; P.-H. Michel; J. Needham; I. Simon; G. Stresser–Péan; J. Théodoridès; J. Vercoutter; Ch. Virolleaud. Isis, 49(4):445–446, December 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227290.

Cohen:1958:BRBa

[991] I. Bernard Cohen. Book review: Dictionary of American Biography by Robert Livingston Schuyler; Edward T. James. Isis, 49(4):446–447, December 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/227291.

Stahlman:1958:BRB

[992] William D. Stahlman. Book review: Vistas in Astronomy by A. Beer. Isis, 49(4):447–449, December 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227292.

Cohen:1958:BRBb

[993] I. Bernard Cohen. Book review: The History of Photography from the Earliest Use of the Camera Obscura in the Eleventh Century up to 1914 by Helmut Gernsheim; Alison Gernsheim; The World’s First Photographer by Alison Gernsheim; L. J. M. Daguerre; Helmut Gernsheim. Isis, 49 (4):449–451, December 1958. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227293.

Nasr:1958:BRB

[994] Seyyed Hossein Nasr. Book review: Alchemy by E. J. Holmyard; Forgerons et Alchimistes by M. Eliade. Isis, 49(4):451–453, December 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227294.

Cohen:1958:BRBc

[995] I. Bernard Cohen. Book review: A History of Magic and Experimental Science. Volumes VII and VIII: The Seventeenth Century by Lynn Thorndike. Isis, 49(4):453–455, December 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227295.

Dijksterhuis:1958:BRB

[996] E. J. Dijksterhuis. Book review: The Exact Sciences in Antiquity by O. Neugebauer. Isis, 49(4):455–456, December 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227296.

Pagel:1958:BRBb

[997] W. Pagel. Book review: Asclepiades, His Life and Writings by Robert Montraville Green; Soranus’ Gynecology by Owsei Temkin; N. J. Eastman; L. Edelstein; A. F. Guttmacher. Isis, 49(4):456–457, December 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227297.

Destombes:1958:BRB

[998] M. Destombes. Book review: Arabic Astronomical and Astrological Sciences in Latin Translation: A Critical Bibliography by Francis J. Carmody. Isis, 49(4):457–458, December 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227298.

Rosen:1958:BRBb

[999] Edward Rosen. Book review: Bibliografia kopernikowska, 1509–1955 by Henryk Baranowski. Isis, 49(4):458–459, December 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227299.

Abetti:1958:BRBc

[1000] Giorgio Abetti. Book review: Discorsi e Dimostrazioni Matematiche intorno a Due Nuove Scienze a cura di Adriano Carugo e Ludovico Geymonat by Galileo Galilei. Isis, 49(4):459–460, December 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227300.

Rosen:1958:BRBc

[1001] George Rosen. Book review: Un Prince, Deux Préfets. Le mouvement scientifique et médico-social au Pays de Liège sous le règne du despotisme éclairé (1771–1830) by Marcel Florkin. Isis, 49(4):460–461, December 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227301.

Dupree:1958:BRB

[1002] A. Hunter Dupree. Book review: The Exploration of the Colorado River by John Wesley Powell. Isis, 49(4):461–462, December 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227302.

Scott:1958:BRB

[1003] J. F. Scott. Book review: Niels Henrik Abel, Mathematician Extraordinary by Øystein Øre. Isis, 49(4):462, December 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227303.

Anonymous:1958:VI

[1004] Anonymous. Volume information. Isis, 49(4):463–465, December 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227279.

Anonymous:1958:BMd

[1005] Anonymous. Back matter. Isis, 49(4):466, December 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227304.

Anonymous:1959:FMa

[1006] Anonymous. Front matter. Isis, 50(1):1–2, March 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/226532.

Guerlac:1959:OR

[1007] Henry Guerlac. To our readers. Isis, 50(1):3–4, March 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226533.

Hoff:1959:GRB

[1008] H. E. Hoff and L. A. Geddes. Graphic registration before ludwig; the antecedents of the kymograph. Isis, 50(1):5–21, March 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226534.

Neugebauer:1959:PGB

[1009] O. Neugebauer. Ptolemy’s geography, book vii, chapters 6 and 7. Isis, 50(1):22–29, March 1959. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226535.

Cope:1959:EBD

[1010] Jackson I. Cope. Evelyn, Boyle, and Dr. Wilkinson’s “Mathematico-Chymico-Mechanical School”. Isis, 50(1):30–32, March 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226536.

Thorndike:1959:NUS

[1011] Lynn Thorndike. Notes upon some medieval astronomical, astrological and mathematical manuscripts at Florence, Milan, Bologna and Venice. Isis, 50(1):33–50, March 1959. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226537.

Marsak:1959:CFF

[1012] Leonard M. Marsak. Cartesianism in Fontenelle and French science, 1686–1752. Isis, 50(1):51–60, March 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226538.

deCamp:1959:SCH

[1013] L. Sprague de Camp. Sailing close-hauled. Isis, 50(1):61–63, March 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226539.

Suter:1959:NC

[1014] Rufus Suter, Saul Benison, and Max Savelle. Notes & correspondence. Isis, 50(1):64–68, March 1959. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226540.

Zirkle:1959:BRBa

[1015] Conway Zirkle. Book review: A History of Luminescence from the Earliest Times until 1900 by E. Newton Harvey. Isis, 50(1):68–69, March 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226541.

Truesdell:1959:BRB

[1016] C. Truesdell. Book review: History of Hydraulics by Hunter Rouse; Simon Ince. Isis, 50(1):69–71, March 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226542.

Saunders:1959:BRBa

[1017] J. B. de C. M. Saunders. Book review: Les Médecins de l’Egypte Pharaonique. Essai de Prosopographie by Frans Jonckheere. Isis, 50(1): 71–72, March 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226543.

Kennedy:1959:BRB

[1018] E. S. Kennedy. Book review: Klyuch Arifmetiki; Traktat ob Okruzhnosti. The Key to Arithmetic and the Treatise on the Circumference by Dzhemshid Giyaseddin Al-Kashi; B. A. Rosenfelda; V. S. Segal; A. P. Yushkevich. Isis, 50(1):72–75, March 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226544.

Giesey:1959:BRB

[1019] Ralph E. Giesey. Book review: The Civilization of the Renaissance in Italy by Jacob Burckhardt; S. G. C. Middlemore. Isis, 50(1):75–76, March 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226545.

Boas:1959:BRB

[1020] Marie Boas. Book review: Natural Magick by John Baptista Porta; Derek J. Price. Isis, 50(1):76, March 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226546.

Gillispie:1959:BRBa

[1021] Charles C. Gillispie. Book review: Oeuvres Philosophiques de Cabanis by Cabanis. Isis, 50(1):76–78, March 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226547.

Hiebert:1959:BRBa

[1022] Erwin N. Hiebert. Book review: Gesammelte Vorträge und Aufsätze by Max Hartmann. Isis, 50(1):78–79, March 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226548.

Hiebert:1959:BRBb

[1023] Erwin N. Hiebert. Book review: Deutsche Chemiker in der ersten Hälfte des neunzehnten Jahrhunderts by Wilhelm Prandtl. Isis, 50(1):79–81, March 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226549.

Farber:1959:BRB

[1024] Eduard Farber. Book review: Proceedings of the Perkin Centennial 1856–1956 Commemorating the Discovery of Aniline Dyes by Howard J. White. Isis, 50(1):81, March 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226550.

Ihde:1959:BRBa

[1025] Aaron Ihde. Book review: Edward Williams Morley (1838–1923), His Influence on Science in America by Howard R. Williams. Isis, 50(1): 81–82, March 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226551.

Hinshaw:1959:BRB

[1026] Virgil Hinshaw, Jr. Book review: Toward a Critical Naturalism. Reflections on Contemporary American Philosophy by Patrick Romanell. Isis, 50(1):82–83, March 1959. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226552.

Anonymous:1959:BMa

[1027] Anonymous. Back matter. Isis, 50(1):84, March 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226553.

Anonymous:1959:FMb

[1028] Anonymous. Front matter. Isis, 50(2):89–151, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226730.

Anonymous:1959:GSM

[1029] Anonymous. The George Sarton Medal for 1958. Isis, 50(2):91–94, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226731.

Stahl:1959:DTE

[1030] William H. Stahl. Dominant traditions in Early Medieval Latin science. Isis, 50(2):95–124, June 1959. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226732.

Guerlac:1959:LML

[1031] Henry Guerlac. A lost memoir of Lavoisier. Isis, 50(2):125–129, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226733.

Zoubov:1959:QOA

[1032] V. P. Zoubov. Quelques observations sur l’auteur du traité anonyme “Utrum dyameter alicuius quadrati sit commensurabilis costae ejusdem”. (French) [Some observations on the author of the anonymous treatise Whether the diameter of the rib of a square is commensurable]. Isis, 50(2):130–134, June 1959. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226734.

Hindle:1959:DRI

[1033] Brooke Hindle and Helen M. Hindle. David Rittenhouse and the illusion of reversible relief. Isis, 50(2):135–140, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226735.

Watanabe:1959:CRF

[1034] Masao Watanabe. Count Rumford’s first exposition of the dynamic aspect of heat. Isis, 50(2):141–144, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226736.

Moore:1959:LPM

[1035] Marian A. Moore and Philip Melanchthon. A letter of Philip Melanchthon to the reader. Isis, 50(2):145–150, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226737.

Aiton:1959:NC

[1036] E. J. Aiton, Stillman Drake, Rufus Suter, Jacob Zeitlin, Roy G. Neville, I. Bernard Cohen, P. H. Brans, and D. A. Wittop Koning. Notes & correspondence. Isis, 50(2):152–157, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226738.

Dupree:1959:NP

[1037] Hunter Dupree, Erwin Hiebert, Chauncey Leake, Harry Woolf, Raymond P. Stearns, Morris Goran, Howard J. B. Ziegler, and Harold I. Sharlin. News of the profession. Isis, 50(2):157–158, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226739.

Holton:1959:BRBa

[1038] Gerald Holton. Book review: Space, Time and Creation. Philosophical Aspects of Scientific Cosmology by Milton K. Munitz. Isis, 50(2):159– 160, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226740.

Holton:1959:BRBb

[1039] Gerald Holton. Book review: Theories of the Universe from Babylonian Myth to Modern Science by Milton K. Munitz. Isis, 50(2):160–161, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226741.

Neville:1959:BRB

[1040] Roy G. Neville. Book review: The Pasteur Fermentation Centennial 1857–1957. Isis, 50(2):161–162, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226742.

Agnew:1959:BRBa

[1041] L. R. C. Agnew. Book review: The Bubonic Plague and England by Charles F. Mullett. Isis, 50(2):162–163, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226743.

Gillispie:1959:BRBb

[1042] Charles C. Gillispie. Book review: A History of Technology, Volume III, from the Renaissance to the Industrial Revolution, c. 1500–c. 1750 by Charles Singer; E. J. Holmyard; A. R. Hall; Trevor I. Williams; Y. Peel; J. R. Petty. Isis, 50(2):163–165, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226744.

Jasny:1959:BRB

[1043] Naum Jasny. Book review: Grain-Mills and Flour in Classical Antiquity by L. A. Moritz. Isis, 50(2):165–169, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226745.

Millas-Vallicrosa:1959:BRB

[1044] J. M. Millás-Vallicrosa. Book review: Tractat d’Astrologia by Bartomeu de Tresbéns. Isis, 50(2):169, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226746.

George:1959:BRB

[1045] Katherine George. Book review: Ethiopian Itineraries circa 1400–1524, including Those Collected by Alessandro Zorzi at Venice in the Years 1519–1524 by O. G. S. Crawford. Isis, 50(2):170–171, June 1959. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226747.

Hellman:1959:BRB

[1046] C. Doris Hellman. Book review: Agricola on Metals. The Age of Technology Waited for Better and More Abundant Metals; It Arrived So Much Sooner because Agricola Published De Re Metallica, a Mining and Metallurgical Classic by Bern Dibner. Isis, 50(2):171–172, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226748.

Roller:1959:BRB

[1047] Duane H. D. Roller. Book review: On the Magnet by William Gilbert; Silvanus P. Thompson; Gilbert Club; De Magnete by William Gilbert; P. F. Mottelay. Isis, 50(2):172–174, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226749.

Pagel:1959:BRBa

[1048] Walter Pagel. Book review: Die Kaerntner Schriften by Theophrast von Hohenheim. Isis, 50(2):174–175, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226750.

Birkmeyer:1959:BRB

[1049] Karl M. Birkmeyer. Book review: Leonardo da Vinci’s Trattato della Pittura by Kate Trauman Steinitz. Isis, 50(2):175–177, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226751.

Rosen:1959:BRB

[1050] Edward Rosen. Book review: Études coperniciennes, Vol. I by Stanislaus Wedkiewicz. Isis, 50(2):177–178, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226752.

Hall:1959:BRB

[1051] A. R. Hall. Book review: Isaac Newton’s Papers and Letters on Natural Philosophy and Related Documents by I. Bernard Cohen; Robert E. Schofield; Isaac Newton. Isis, 50(2):178–179, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226753.

Jackman:1959:BRB

[1052] Sydney W. Jackman. Book review: Alexander Gordon, M.D. of Aberdeen, 1752–1799 by Ian A. Porter. Isis, 50(2):179–180, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226754.

Brinton:1959:BRB

[1053] Crane Brinton. Book review: Carl Becker’s Heavenly City Revisited by Raymond O. Rockwood. Isis, 50(2):181–182, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226755.

Taton:1959:BRBa

[1054] René Taton. Book review: Jean-Sylvain Bailly, Astronomer, Mystic, Revolutionary, 1736–1793 by Edwin Burrows Smith. Isis, 50(2):182– 183, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226756.

Ihde:1959:BRBb

[1055] Aaron Ihde. Book review: The Chemical Industry during the Nineteenth Century. A Study of the Economic Aspects of Applied Chemistry in Europe and North America by L. F. Haber. Isis, 50(2):183–184, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226757.

Zirkle:1959:BRBb

[1056] Conway Zirkle. Book review: Gregor Mendel und das Schicksal seiner Vererbungsgesetze by Ingo Krumbiegel. Isis, 50(2):184–185, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226758.

Hiebert:1959:BRBc

[1057] Erwin N. Hiebert. Book review: Johann Conrad Fischer, 1773–1854 by Karl Schib; Rudolf Gnade. Isis, 50(2):186–187, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226759.

Anonymous:1959:PRD

[1058] Anonymous. Publications received December 1958–February 1959. Isis, 50(2):188–190, June 1959. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226760.

Hindle:1959:AD

[1059] Brooke Hindle. Administrative documents. Isis, 50(2):191–192, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226761.

Anonymous:1959:BMb

[1060] Anonymous. Back matter. Isis, 50(2):??, June 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226762.

Anonymous:1959:FMc

[1061] Anonymous. Front matter. Isis, 50(3):197–198, September 1959. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227305.

Scott:1959:SHB

[1062] Wilson L. Scott. The significance of “hard bodies” in the history of scientific thought. Isis, 50(3):199–210, September 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227306.

Partington:1959:SEA

[1063] J. R. Partington. Some early appraisals of the work of John Mayow. Isis, 50(3):211–226, September 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227307.

Kennedy:1959:PTI

[1064] E. S. Kennedy and Victor Roberts. The planetary theory of ibn al-Shāt.ir. Isis, 50(3):227–235, September 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227308.

Steadman:1959:CMN

[1065] John M. Steadman. Chauntecleer and medieval natural history. Isis, 50 (3):236–244, September 1959. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227309.

Drake:1959:GGV

[1066] Stillman Drake. Galileo gleanings vi: Galileo’s first telescopes at Padua and Venice. Isis, 50(3):245–254, September 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227310.

deSantillana:1959:BRB

[1067] Giorgio de Santillana and Stillman Drake. Book review: The Sleepwalkers. Isis, 50(3):255–260, September 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227311.

Hall:1959:NC

[1068] A. Rupert Hall and J. R. Ravetz. Notes & correspondence. Isis, 50(3): 261–265, September 1959. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227312.

Anonymous:1959:NP

[1069] Anonymous. News of the profession. Isis, 50(3):265–267, September 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227313.

Black:1959:BRB

[1070] Max Black. Book review: Patterns of Discovery by N. R. Hanson. Isis, 50 (3):267–268, September 1959. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/227314.

Boyer:1959:BRBa

[1071] Carl B. Boyer. Book review: History of Mathematics by David Eugene Smith. Isis, 50(3):268–269, September 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227315.

White:1959:BRB

[1072] George W. White. Book review: Glossary of Geology and Related Sciences by J. V. Howell. Isis, 50(3):269–270, September 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227316.

Galdston:1959:BRB

[1073] Iago Galdston. Book review: Kurze Geschichte der Psychiatrie by Erwin H. Ackerknecht. Isis, 50(3):270–271, September 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227317.

Schmidl:1959:BRB

[1074] Fritz Schmidl. Book review: Existence, a New Dimension in Psychiatry and Psychology by Rollo May; Ernest Angel; Henri F. Ellenberger. Isis, 50(3):271–272, September 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227318.

Destombes:1959:BRB

[1075] M. Destombes. Book review: A Survey of Islamic Astronomical Tables by E. S. Kennedy. Isis, 50(3):272–273, September 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227319.

Nasr:1959:BRB

[1076] Seyyed Hossein Nasr. Book review: Psychologie d’Ibn Sina (Avicenne) d’après son oeuvre As Sifa’ by Ibn Sina; Jan Bakos. Isis, 50(3):273–274, September 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227320.

Saunders:1959:BRBb

[1077] J. B. deC. M. Saunders. Book review: Paracelsus. An Introduction to Philosophical Medicine in the Era of the Renaissance by Walter Pagel. Isis, 50(3):274–276, September 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227321.

Pagel:1959:BRBb

[1078] Walter Pagel. Book review: Pagan Mysteries in the Renaissance by Edgar Wind. Isis, 50(3):276–278, September 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/227322.

Cohen:1959:BRB

[1079] I. Bernard Cohen. Book review: Collected Letters by Antoni van Leeuwenhoek; Committee of Dutch Scientists. Isis, 50(3):278–279, September 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227323.

Burger:1959:BRB

[1080] D. Burger. Book review: Het Land van Stevin en Huygens by D. J. Struik. Isis, 50(3):280–281, September 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

227324.

Boyer:1959:BRBb

[1081] Carl B. Boyer. Book review: The Principal Works of Simon Stevin. Volume II (A & B), Mathematics by D. J. Struik. Isis, 50(3):281–282, September 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/227325.

Zirkle:1959:BRBc

[1082] Conway Zirkle. Book review: Evolution by Natural Selection by Charles Darwin; Alfred Russell Wallace; Darwin’s Century by Loren Eiseley; A Century of Darwin by S. A. Barnett. Isis, 50(3):282–285, September 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227326.

Scott:1959:BRB

[1083] J. F. Scott. Book review: Carl Friedrich Gauss: Titan of Science. A Study of His Life and Work by C. Waldo Dunnington. Isis, 50(3):285– 286, September 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/227327.

Roberts:1959:BRB

[1084] David Roberts. Book review: Scottish Social Welfare, 1864–1914 by Thomas Ferguson. Isis, 50(3):286–287, September 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/227328.

Dupree:1959:BRB

[1085] A. Hunter Dupree. Book review: The Scientific Thought of Henry Adams by Henry Wasser. Isis, 50(3):288, September 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227329.

Cohen:1959:EFC

[1086] I. Bernard Cohen, Harry Woolf, and Phyllis Brooks Bosson. Eightyfourth critical bibliography of the history of science and its cultural influences (to 1 January 1959). Isis, 50(3):289–407, September 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/227330.

Anonymous:1959:BMc

[1087] Anonymous. Back matter. Isis, 50(3):408, September 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/227331.

Anonymous:1959:FMd

[1088] Anonymous. Front matter. Isis, 50(4):417–418, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226425.

Clagett:1959:IAM

[1089] Marshall Clagett. The impact of Archimedes on medieval science. Isis, 50 (4):419–429, December 1959. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226426.

Jarcho:1959:OAI

[1090] Saul Jarcho. Origin of the American Indian as suggested by Fray Joseph de Acosta (1589). Isis, 50(4):430–438, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226427.

Weisheipl:1959:PJD

[1091] James A. Weisheipl. The place of John Dumbleton in the Merton School. Isis, 50(4):439–454, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226428.

Zeitlin:1959:TSD

[1092] Jacob Zeitlin. Thomas Salusbury discovered. Isis, 50(4):455–458, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226429.

Wilczynski:1959:PDA

[1093] Jan Z. Wilczynski. On the presumed Darwinism of Alberuni eight hundred years before Darwin. Isis, 50(4):459–466, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226430.

Turbayne:1959:GAO

[1094] Colin M. Turbayne. Grosseteste and an ancient optical principle. Isis, 50 (4):467–472, December 1959. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226431.

Hall:1959:NES

[1095] Marie Boas Hall and A. Rupert Hall. Newton’s electric spirit: Four oddities. Isis, 50(4):473–476, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226432.

Neugebauer:1959:RPA

[1096] O. Neugebauer. Regula Philippi Arrhidaei. Isis, 50(4):477–478, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226433.

Carlson:1959:NC

[1097] Eric T. Carlson, Erwin Reifler, C. Truesdell, Roy G. Neville, and L. Pearce Williams. Notes & correspondence. Isis, 50(4):479–481, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226434.

Paape:1959:NP

[1098] Charles W. Paape. News of the profession. Isis, 50(4):482–484, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226435.

Schofield:1959:BRB

[1099] Robert E. Schofield. Book review: Moments of Discovery by P. H. Swartz; Philip W. Bishop. Isis, 50(4):484–486, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226436.

Greene:1959:BRB

[1100] John C. Greene. Book review: Men of Science in America by Bernard Jaffe. Isis, 50(4):486–487, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226437.

Barber:1959:BRB

[1101] Bernard Barber. Book review: Scientific Manpower in Europe. A Comparative Study of Scientific Manpower in the Public Service of Great Britain and Selected European Countries by Edward McCrensky. Isis, 50 (4):487–488, December 1959. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226438.

Smullyan:1959:BRB

[1102] Arthur Smullyan. Book review: Observation and Interpretation by S. Körner. Isis, 50(4):488–489, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226439.

Taton:1959:BRBb

[1103] René Taton. Book review: History of Analytic Geometry by Carl B. Boyer. Isis, 50(4):489–492, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226440.

Zirkle:1959:BRBd

[1104] Conway Zirkle. Book review: The Sloane Herbarium by J. E. Dandy; James Britten. Isis, 50(4):492–493, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226441.

Poynter:1959:BRB

[1105] F. N. L. Poynter. Book review: Medicine and the Navy 1200–1900. Volume II, 1649–1714 by J. J. Keevil. Isis, 50(4):493–494, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226442.

Sonnedecker:1959:BRB

[1106] Glenn Sonnedecker. Book review: Herbals of Five Centuries. A Contribution to Medical History and Bibliography by Claus Nissen. Isis, 50 (4):494–495, December 1959. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226443.

Starr:1959:BRB

[1107] Chester G. Starr. Book review: The Ancient Mariners; Seafarers and Sea Fighters of the Mediterranean in Ancient Times by Lionel Casson. Isis, 50(4):495–496, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

226444.

Ward:1959:BRB

[1108] F. A. B. Ward. Book review: Time and Its Measurement; From the Stone Age to the Nuclear Age by Harrison J. Cowan. Isis, 50(4):496–498, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226445.

Drachmann:1959:BRB

[1109] A. G. Drachmann. Book review: Studies in Early Petroleum History by R. J. Forbes. Isis, 50(4):498–499, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226446.

Oliver:1959:BRB

[1110] John W. Oliver. Book review: The Technical Development of the Royal Dutch/Shell, 1890–1940 by R. J. Forbes; R. O’Beirne. Isis, 50(4):500– 501, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226447.

Jacobs:1959:BRB

[1111] Melville Jacobs. Book review: From Ape to Angel by H. R. Hays. Isis, 50 (4):501–502, December 1959. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226448.

Pascal:1959:BRB

[1112] Paul Pascal. Book review: Libellus de Natura Animalium. Isis, 50(4): 502–503, December 1959. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226449.

Wolfnd:1959:BRB

[1113] Edwin Wolf, 2nd. Book review: The Americans; The Colonial Experience by Daniel J. Boorstin. Isis, 50(4):503–504, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226450.

Loemker:1959:BRB

[1114] L. E. Loemker. Book review: Gottfried Wilhelm Leibniz; Barock und Universalismus by Joachim Otto Fleckenstein. Isis, 50(4):504–506, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/226451.

Johnson:1959:BRB

[1115] Francis R. Johnson. Book review: The Anatomy of the World. Relations between Natural and Moral Law from Donne to Pope by Michael Macklem. Isis, 50(4):506–507, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226452.

Clark:1959:BRB

[1116] Joseph T. Clark. Book review: Pierre Gassendi: sa Vie et Son OEuvre, 1592–1655 by Bernard Rochot; Alexandre Koyré; Georges Mongrédien; Antoine Adam. Isis, 50(4):507–510, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226453.

Stimson:1959:BRB

[1117] Dorothy Stimson. Book review: History of the Royal Society by Thomas Sprat; Jackson T. Cole; Harold Whitmore Jones. Isis, 50(4):510–511, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226454.

Agnew:1959:BRBb

[1118] L. R. C. Agnew. Book review: William Harvey and the Circulation of the Blood. Isis, 50(4):511–512, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226455.

Blake:1959:BRB

[1119] John B. Blake. Book review: The Johns Hopkins Hospital and the Johns Hopkins University School of Medicine: A Chronicle. Vol. II, 1893–1905 by Alan M. Chesney. Isis, 50(4):512–514, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226456.

Hoff:1959:BRB

[1120] H. E. Hoff. Book review: Experimental Psychology and Other Essays by I. Pavlov. Isis, 50(4):514–516, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/226457.

Hewlett:1959:BRB

[1121] Richard G. Hewlett. Book review: Adventures in the World of Science by Charles G. Abbot. Isis, 50(4):516–517, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226458.

Vucinich:1959:BRB

[1122] Alexander Vucinich. Book review: Archaeology in the U.S.S.R. by Mikhail Miller. Isis, 50(4):517–518, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226459.

Anonymous:1959:PRM

[1123] Anonymous. Publications received 1 March 1959–1 September 1959. Isis, 50(4):518–522, December 1959. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226460.

Anonymous:1959:EEN

[1124] Anonymous. Errata: Evolution by natural selection. Isis, 50(4):522, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226461.

Anonymous:1959:BMd

[1125] Anonymous. Back matter. Isis, 50(4):523–530, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226462.

Anonymous:1959:VI

[1126] Anonymous. Volume information. Isis, 50(4):524–529, December 1959. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226424.